### v1

# Improved High-probability Convergence Guarantees of Decentralized SGD

Abstract
Convergence in high-probability (HP) has been receiving increasing interest, due to its attractive properties, such as exponentially decaying tail bounds and strong guarantees for each individual run of an algorithm. While HP guarantees are extensively studied in centralized settings, much less is understood in the decentralized, networked setup. Existing HP studies in decentralized settings impose strong assumptions, like uniformly bounded gradients, or asymptotically vanishing noise, resulting in a significant gap between assumptions used to establish convergence in the HP and the mean-squared error (MSE) sense, even for vanilla Decentralized Stochastic Gradient Descent (DSGD) algorithm. This is contrary to centralized settings, where it is known that SGD converges in HP under the same conditions on the cost function as needed to guarantee MSE convergence. Motivated by this observation, we revisit HP guarantees for DSGD in the presence of light-tailed noise. We show that DSGD converges in HP under the same conditions on the cost as in the MSE sense, removing uniformly bounded gradients and other restrictive assumptions, while simultaneously achieving order-optimal rates for both non-convex and strongly convex costs. Moreover, our improved analysis yields linear speed-up in the number of users, demonstrating that DSGD maintains strong performance in the HP sense and matches existing MSE guarantees. Our improved results stem from a careful analysis of the MGF of quantities of interest (norm-squared of gradient or optimality gap) and the MGF of the consensus gap between users’ models. To achieve linear speed-up, we provide a novel result on the variance-reduction effect of decentralized methods in the HP sense and more fine-grained bounds on the MGF for strongly convex costs, which are both of independent interest.
1 Introduction
Modern large-scale machine learning applications and the abundance of data necessitate alternatives to the centralized computation framework, giving rise to distributed learning, a paradigm where multiple users collaborate to jointly train a model, e.g., [1, 2, 3]. The features of distributed learning, such as storing the data locally and only exchanging smaller updates, like (quantized) model parameters or gradients, and the lack of a single point of failure, further make it an attractive paradigm from a privacy and security perspective [4, 5]. Many applications, such as federated training of models on mobile devices [6], controlling and coordinating robot swarms [7], or distributed control and power grids [8], all rely on distributed computation. From a communication/model exchange perspective, distributed frameworks can be client-server (i.e., federated) or decentralized (i.e., networked), with the main difference being that in the client-server setup users communicate with a server, while in decentralized settings users communicate directly with each other. [^1] While the client-server setup is often characterized by performing local updates and periodic communication [9], as well as partial user participation [10], we note that these features can also be incorporated in the decentralized setup, by modifying the methods to perform multiple local updates before communicating [11] and considering a dynamic network with users that are occasionally idle [12]. However, this is not the focus of our current work. Noting that, from the model update perspective, the client-server setup is equivalent to the decentralized setup with a fully connected communication network, we focus on the more general, decentralized setup.
The study of convergence guarantees of decentralized optimization algorithms has a long history, e.g., [13, 14, 15, 16, 17, 18, 19], with most works focusing on MSE convergence, e.g., [20, 21, 22, 23, 11, 24, 25, 26], see also [27] for an extensive treatment of the topic. Another type of convergence guarantees, namely convergence in HP, has garnered increasing attention recently. In particular, for a non-negative stochastic process $\{X^{t}\}_{t\in\mathbb{N}}$, the goal of HP convergence guarantees is to establish, for all $t\in\mathbb{N}$ and any $\epsilon>0$, that
$\mathbb{P}\big(X^{t}>\epsilon\big)\leq\exp\big(-Ct^{\gamma_{1}}\epsilon^{\gamma_{2}}\big),$ (1)
where $\gamma_{1},\gamma_{2},C>0$ are some positive constants. If $\{X^{t}\}_{t\in\mathbb{N}}$ is a measure of performance of an algorithm, e.g., $X^{t}=\frac{1}{t}\sum_{k\in[t]}\|\nabla f(x^{k})\|^{2}$, where $\{x^{t}\}_{t\in\mathbb{N}}$ is a sequence generated by the algorithm and $f$ is a non-convex cost, the relation (1) provides strong guarantees with respect to a single run of the algorithm. This is particularly important in modern applications like LLMs, where it is often intractable to perform multiple runs. Numerous works study HP guarantees of SGD-type methods in centralized settings, under both light-tailed [28, 29, 30, 31, 32, 33] and heavy-tailed noise [34, 35, 36, 37, 38]. Comparatively, there have been few studies on HP convergence of decentralized methods.
1.1 Literature Review
We now review the literature, focusing on works studying centralized HP convergence guarantees and decentralized MSE and HP convergence guarantees.
Centralized HP Convergence.
The authors in [28] show optimal convergence rates of SGD for convex costs under light-tailed (i.e., sub-Gaussian) stochastic gradients (see assumption (A4) ahead for a formal definition of sub-Gaussianity), while the work [29] shows the same for non-convex costs. The authors in [30, 39] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively. The works [31] and [32] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while in [40] the authors establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs. The work [33] generalizes the previous studies for non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs. The authors in [41] study HP convergence of SGD under sub-Weibull noise. Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [42, 34, 38] for various conditions used when studying heavy-tailed noise. e.g., [43, 34, 35, 44, 36, 37, 45, 46, 38], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1). Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
Decentralized MSE Convergence.
MSE guarantees are typically studied under a variety of bounds on the second noise moment. [^3] See, e.g., [47] for a good overview of the various conditions used in the literature. Following one of these settings, the work [20] shows that DSGD converges at an optimal rate for strongly convex costs, while in [21] the authors show that DSGD with a fixed step-size and the gradient tracking (GT) mechanism, e.g., [16, 48, 49], converges to a neighbourhood of the optimal solution. The works [22, 23, 50] study MSE guarantees of DSGD for non-convex costs and show that it escapes saddle points with high probability. In [26] the authors propose a general framework dubbed cooperative SGD, showing optimal rates for non-convex costs. The work [51] provides unified guarantees for DSGD with local updates and changing network topology, with optimal rates and linear speed-up in the number of users for non-convex and (strongly) convex costs. In [24] the authors show DSGD with GT and variance reduction converges at a linear rate for strongly convex costs, while in [25] optimal rates of DSGD with GT and linear speed-up in the number of users are established, for non-convex costs and costs satisfying the Polyak-Łojasiewicz (PL) condition, e.g., [52]. Finally, in [53] the authors show that DSGD with GT, local momentum and gradient normalization achieves optimal rates under heavy-tailed noise. It is also worth mentioning a rich line of works studying MSE guarantees in decentralized settings for problems such as estimation and detection, multi-objective and multitask optimization, see [54, 55, 56] and references therein.
Decentralized HP Convergence.
Compared to MSE guarantees, there is a significantly smaller body of work on HP convergence guarantees of decentralized algorithms. In particular, the work [57] studies HP convergence of DSGD for both non-convex and PL costs under light-tailed noise, requiring uniformly bounded gradients and asymptotically vanishing noise for PL costs (see discussion after Theorem 2 ahead), showing optimal rates in both cases. In [58, 59] the authors study HP convergence of a decentralized mirror descent algorithm under sub-Gaussian noise and bounded gradients, for online noncooperative games and dynamic regrets, respectively. Works [60, 61, 62] study HP convergence of decentralized algorithms with clipping under heavy-tailed noise. It is worth mentioning [63, 64, 65, 66, 67], who study asymptotic large deviation guarantees for decentralized problems like detection and inference.
A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradient stays bounded. This observation raises the question of whether decentralized algorithms can converge in HP without ensuring that the gradient is uniformly bounded, creating a gap between convergence in the HP and the MSE sense, where bounded gradients are not required. This is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence. Moreover, none of the existing HP convergence bounds in decentralized settings show linear speed-up in the number of users, established in the MSE sense in, e.g., [11, 25].
1.2 Contributions
Motivated by the observed gap between the existing literature on HP and MSE convergence in decentralized settings, we revisit the HP convergence guarantees in decentralized stochastic optimization, by studying a variant of vanilla DSGD in the presence of light-tailed noise. We establish that DSGD converges in HP under the same conditions on the cost as required in the MSE sense, for both non-convex and strongly convex costs, removing the uniformly bounded gradient condition. Next, we show that DSGD achieves optimal rates and linear speed-up in the number of users, for both non-convex and strongly convex costs, further improving on existing works and closing the gap between HP and MSE guarantees in decentralized settings. Our results are established by carefully bounding the MGF of the quantity of interest (i.e., average norm-squared of the gradient for non-convex and optimality gap of the final iterate for strongly convex costs) and the MGF of the consensus gap. Compared to [57], the work closest to ours, we provide several improvements. In particular, we remove the uniformly bounded gradient requirement, as well as the vanishing noise requirement imposed for PL costs (see the discussion after Theorem 2 ahead), while achieving linear speed-up in the number of users. Compared to works studying MSE guarantees for DSGD, e.g., [11, 20], our results require directly working with the MGF and carefully balancing between the MGF of the quantity of interest and that of the consensus gap. This is particularly challenging for strongly convex costs, where we want to show an “almost decreasing” property of the MGF, in order to get improved rates on the last iterate (see Lemma 5 and the related discussion ahead). Compared to centralized SGD, e.g., [31, 32, 33, 40], the main challenge lies in the additional consensus gap and simultaneously dealing with its MGF and that of the quantity of interest. All of these challenges are resolved by introducing several novelties, outlined next.
Novelty.
Toward establishing our improved HP convergence guarantees, we face several technical challenges, requiring novel results. In particular, to remove uniformly bounded gradients, we provide Lemma 3 and use the “offset trick” for non-convex costs (see the proof sketch of Theorem 1 in Section 4 for details), while for strongly convex costs we establish a novel bound on the MGF of the consensus gap, in the form of Lemma 4. To establish linear speed-up in the number of users, we show that the variance reduction benefit of decentralized learning is maintained in the HP sense, in the form of Lemma 2, which is of independent interest when studying HP guarantees of decentralized algorithms. Further, to guarantee linear speed-up for strongly convex costs, we require a more fine-grained bound on the MGF, provided in the form of Lemma 5, which allows for incorporating higher-order terms, generalizing the bounds from [31, 32, 40] and is of independent interest, even in centralized settings.
Paper Organization.
The rest of the paper is organized as follows. Section 2 formally introduces the problem and the DSGD method, Section 3 outlines the main results, Section 4 provides proof sketches and discussion, while Section 5 concludes the paper. Appendix contains results omitted from the main body. The remainder of this section introduces the notation.
Notation.
We use $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{R}^{d}$ to denote positive integers, real numbers and $d$-dimensional vectors. For $m\in\mathbb{N}$, we use $m]=\{1,\ldots,m\}$ to denote positive integers up to and including $m$. The notation $\langle\cdot,\cdot\rangle$ stands for the Euclidean inner product, while $\|\cdot\|$ is used for both the induced vector and matrix $\ell_{2}$ norms. We use subscripts to denote users and superscripts to denote the iteration counter, e.g., $x^{t}_{i}$ refers to the model of user $i$ in iteration $t$. We use the “big O” notation $\mathcal{O}(\cdot)$ to only hide global constants, unless stated otherwise.
2 Problem Setup and Proposed Method
In this section we introduce the problem of decentralized stochastic optimization and the DSGD algorithm. Consider a network of $n\geq 2$ users which can communicate with each other and want to jointly train a model. Formally, the problem can be cast as
$\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}\bigg\{f(x)=\frac{1}{n}\sum_{i\in[n]}f_{i}(x)\bigg\},$ (2)
where $x\in\mathbb{R}^{d}$ represents model parameters, $f_{i}:\mathbb{R}^{d}\mapsto\mathbb{R}$ is the cost function associated to user $i\in[n]$, given by $f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[\ell(x;\xi_{i})]$, with $\xi_{i}\in\Xi$ being a random variable governed by some unknown distribution $\mathcal{D}_{i}$, while $\ell:\mathbb{R}^{d}\times\Xi\mapsto\mathbb{R}$ is a loss function. Each user has access to a Stochastic First-order Oracle ($\mathcal{SFO}$), which, when queried by user $i\in[n]$ and input $x\in\mathbb{R}^{d}$, returns the gradient of $\ell$ evaluated at a random sample $\xi_{i}\sim\mathcal{D}_{i}$, i.e., $\nabla\ell(x;\xi_{i})$. The $\mathcal{SFO}$ model subsumes:
  1. 1.
    Batch (i.e., offline) learning, where users have access to a local dataset $\{\xi_{i,l}\}_{l\in[m_{i}]}$, so that $f_{i}(x)=\frac{1}{m_{i}}\sum_{l\in[m_{i}]}\ell(x;\xi_{i,l})$ and in each round users choose a sample $\xi_{i,l}$ uniformly at random, [^4] Note that it is possible to have non-uniform weights, i.e., $f_{i}(x)=\sum_{l\in[m_{i}]}p_{i,l}\ell(x;\xi_{i,l})$, where $p_{i,l}>0$ for all $l\in[m_{i}]$ and $\sum_{l\in[m_{i}]}p_{i,l}=1$, resulting in non-uniform sampling, with sample $\xi_{i,l}$ selected with probability $p_{i,l}$. which is used to compute $\nabla\ell(x;\xi_{i,l})$ and update model parameters;
  2. 2.
    Streaming (i.e., online) learning, where users do not store a local dataset, but in each round observe a random sample $\xi_{i}\sim\mathcal{D}_{i}$ from a data stream, which is used to compute $\nabla\ell(x;\xi_{i})$ and update model parameters.
The communication pattern among users is modeled as a static graph $G=(V,E)$, where $V=[n]$ is the set of vertices, representing users, while $E\subset V\times V$ is the set of edges, representing communication links between users. To solve (2) in decentralized fashion, we consider a version of Decentralized Stochastic Gradient Descent (DSGD), based on the Adapt-Then-Combine (i.e., diffusion) approach, e.g., [68, 69, 70]. The method consists of the following steps. At the start, users choose a shared step-size schedule $\{\alpha_{t}\}_{t\in\mathbb{N}}$ and each user $i\in[n]$ chooses an arbitrary, but deterministically selected initial model $x^{1}_{i}\in\mathbb{R}^{d}$. [^5] While the initial models can be any real vectors, possibly different across users, they needs to be deterministic quantities, for the sake of theoretical analysis. In iteration $t\geq 1$, users query the $\mathcal{SFO}$ with their current model $x^{t}_{i}$ and receive $\nabla\ell(x^{t}_{i};\xi_{i}^{t})$. Users first update their local model via the rule $x_{i}^{t+\nicefrac{{1}}{{2}}}=x^{t}_{i}-\alpha_{t}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$, after which the new model is produced by performing a consensus step, i.e., $x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}x^{t+\nicefrac{{1}}{{2}}}_{j}$, where $\mathcal{N}_{i}\coloneqq\{j\in V:\{i,j\}\in E\}\cup\{i\}$ is the set of users (i.e., neighbours) with whom user $i$ can communicate (including user $i$ itself), while $w_{ij}>0$ is the weight user $i$ assigns to user $j$’s model. The update rule at user $i\in[n]$ can be compactly represented as
$x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\big(x_{j}^{t}-\alpha_{t}\nabla\ell(x_{j}^{t};\xi_{j}^{t})\big).$ (3)
The method is summarized in Algorithm 1. The version of DSGD considered in our work is related to the version based on Combine-Then-Adapt (i.e., consensus ++ innovation) approach, e.g., [13, 54, 71], whose HP convergence is studied in [57].
Algorithm 1 DSGD
0: Model initialization $x^{1}_{i}\in\mathbb{R}^{d}$, $i\in[n]$, step-size schedule $\{\alpha_{t}\}_{t\in\mathbb{N}}$;
1:for $t=1,2,\ldots$, each user $i\in[n]$ in parallel do
2:  Query the oracle to obtain $\nabla\ell(x^{t}_{i};\xi_{i}^{t})$;
3:  Perform the model update: $x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\big(x_{j}^{t}-\alpha_{t}\nabla\ell(x_{j}^{t};\xi_{j}^{t})\big)$;
4:end for
3 Main Results
In this section we present the main results. Subsection 3.1 states the preliminaries, Subsection 3.2 provides results for non-convex costs, while Subsection 3.3 presents results for strongly convex costs.
3.1 Preliminaries
In this subsection we outline the assumptions used in our work. For any $T\geq 1$, let $\{\xi^{t}_{i}\}_{t\in[T]}$ be the random samples observed by user $i\in[n]$ up to time $T$ and denote by $\mathcal{F}_{T}$ the natural filtration with respect to the sequence of user models up to time $T$, i.e., $\mathcal{F}_{T}\coloneqq\sigma\left(\left\{\{x^{1}_{i}\}_{i\in[n]},\ldots,\{x_{i}^{T}\}_{i\in[n]}\right\}\right)$. For ease of notation, let $z^{t}_{i}\coloneqq\nabla\ell(x^{t}_{i};\xi_{i}^{t})-\nabla f_{i}(x^{t}_{i})$ and $W\in\mathbb{R}^{n\times n}$, where $W]_{i,j}\coloneqq w_{ij}$, denote the stochastic noise and the network communication matrix, respectively.
(A1) The network communication matrix $W\in\mathbb{R}^{n\times n}$ is primitive and doubly stochastic.
Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [24]. Moreover, it can be shown that (A1) implies that the second largest singular value of $W$, denoted by $\lambda$, satisfies $\lambda\in[0,1)$, see [72].
(A2) The global cost $f$ is bounded from below, i.e., $f^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}f(x)>$-$\infty$.
(A3) Each local cost $f_{i}$ has $L$-Lipschitz gradients, i.e., $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, for any $x,y\in\mathbb{R}^{d}$.
Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [29, 11, 25].
(A4) The stochastic quantities satisfy the following:
  1. 1.
    The random samples $\{\xi^{t}_{i}\}_{i\in[n],t\in[T]}$, are independent across users and iterations.
  2. 2.
    The stochastic gradients are unbiased, i.e., for any $i\in[n]$, $t\geq 1$ and $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
    $\mathbb{E}[\nabla\ell(x;\xi_{i}^{t})\>|\>\mathcal{F}_{t}]=\nabla f_{i}(x).$
  3. 3.
    The noise at each user $i\in[n]$ is $\sigma_{i}$-sub-Gaussian, i.e., for all $t\geq 1$ and any $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
    $\mathbb{E}\left[\exp\bigg(\frac{\|z_{i}^{t}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1).$
The first condition in (A4) is standard in decentralized stochastic optimization [25, 53], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian). Light tails are necessary to achieve exponentially decaying tail bounds of the form (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [28, 29, 30, 39, 31, 33, 32]. [^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [41] and references therein, the focus of this work is on the more widely studied sub-Gaussian noise setting.
(A5) The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, [^7] With at least one of $A$ or $B$ being strictly positive. we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
A heterogeneity bound of the type (A5) is in fact required to ensure that DSGD converges for non-convex costs [73, 8, 11, 26]. Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [57], as it allows users’ gradients to grow with the global gradient. Compared to [11], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition. While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [11], to keep the proofs simple and instructive, we will use assumption (A5).
(A6) Each $f_{i}$ is twice continuously differentiable and $\mu$-strongly convex, i.e., for every $x,y\in\mathbb{R}^{d}$, we have $f_{i}(x)\geq f_{i}(y)+\langle\nabla f_{i}(y),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}$.
Assumption (A6) is used in the analysis of strongly convex costs. It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [20, 11, 26] which is in general not required for centralized SGD. The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [25], however, this is beyond the scope of the current work.
Next, let $\overline{x}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}x^{t}_{i}$ denote the network-wise average model. Note that $\overline{x}^{t}$ is an ideal model, used for analysis purposes. Using (3) and the fact that $W$ is doubly stochastic, it can be readily seen that
$\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t},$
where $\overline{g}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$ is the average stochastic gradient. Let $\overline{z}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}z_{i}^{t}$ denote the average noise. We then have the following result.
Lemma 1.
Let (A3) hold. If $\alpha_{t}\leq\frac{1}{2L}$, we have
$f(\overline{x}^{t+1})\leq f(\overline{x}^{t})-\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}^{2}L\|\overline{z}^{t}\|^{2}+\frac{\alpha_{t}L^{2}}{2n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$
Lemma 1 provides an important deterministic descent-type inequality for DSGD and is the starting point for our analysis. The right-hand side of the above inequality consists of terms that arise in centralized SGD plus a consensus gap term $\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}$, which stems from the decentralized nature of the algorithm and bounding this term is crucial to ensure convergence. To finish this subsection, we provide a technical result on the behaviour of the stochastic noise in the HP sense. Let $\sigma^{2}\coloneqq\frac{1}{n}\sum_{i\in[n]}\sigma^{2}_{i}$ be the average noise parameter. We then have the following result.
Lemma 2.
If (A4) holds, then the following are true for any $t\geq 1$, $i\in[n]$ and $\mathcal{F}_{t}$-measurable $v\in\mathbb{R}^{d}$.
  1. 1.
    $\mathbb{E}\left[\exp\left(\langle v,z^{t}_{i}\rangle\right)\>|\>\mathcal{F}_{t}\right]\leq\exp\left(\frac{3\sigma_{i}^{2}\|v\|^{2}}{4}\right)$.
  2. 2.
    $\mathbb{E}\left[\exp\left(\langle v,\overline{z}^{t}\rangle\right)\>|\>\mathcal{F}_{t}\right]\leq\exp\left(\frac{3\sigma^{2}\|v\|^{2}}{4n}\right)$.
  3. 3.
    The average noise $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian, i.e., $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{120d\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1)$.
Lemma 2 establishes some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, highlighting the variance reduction benefit of decentralized optimization in the HP sense. This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we establish in the following subsections. We note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and the Appendix.
3.2 Non-convex Costs
In this subsection we present results for non-convex costs. Let $\Delta_{x}\coloneqq\frac{1}{n}\sum_{i\in[n]}\|x^{1}_{i}-\overline{x}^{1}\|^{2}$ denote the initial consensus gap. We then have the following result.
Lemma 3.
Let (A1) and (A5) hold. We then have, for any $t\geq 1$
$\frac{1}{n}\sum_{i\in[n]}$ $\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\leq 2\lambda^{2t}\Delta_{x}+\frac{4\lambda^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\|z^{k}_{i}\|^{2}$
$+\frac{4\lambda^{2}A^{2}}{1-\lambda}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}+\frac{4\lambda^{2}B^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\|\nabla f(x^{k}_{i})\|^{2}.$
Lemma 3 provides a deterministic bound on the consensus gap and is used to bound the resulting MGF. Define $\Delta_{f}\coloneqq f(\overline{x}^{1})-f^{\star}$ and $\overline{\sigma}\coloneqq\max_{i\in[n]}\sigma_{i}$. Building on Lemmas 1 and 3, we get the following result.
Theorem 1.
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$, where $C>0$ is a problem related constant satisfying
$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
we then have, for any $\delta\in(0,1)$, with probability at least $1-\delta$
$\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $=\mathcal{O}\Bigg(\frac{\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+L\big)}{\sqrt{nT}}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}$
$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{n\lambda^{2}(A^{2}+\sigma^{2})]}{\sigma^{2}(1-\lambda^{2})T}\Bigg).$
Theorem 1 establishes HP convergence guarantees for DSGD with general smooth non-convex costs and a fixed step-size. The bound in Theorem 1 consists of three terms, with the first two matching centralized SGD, while the third captures the effect of the network, through $\frac{\lambda^{2}}{1-\lambda^{2}}$. Next, note that the leading term $\mathcal{O}\big(\frac{\sigma\sqrt{d}}{\sqrt{nT}}\big)$ achieves a linear speed-up in the number of users, showing the benefits of decentralized learning in the HP sense. Finally, we can see that, in the case $A=0$, the above bound recovers the rates of noiseless centralized gradient descent (GD), i.e., when $\sigma=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$. Compared to [57], where the authors provide a $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$ HP convergence bound, while requiring uniformly bounded gradients, we provide several improvements, by relaxing the uniformly bounded gradients assumption, achieving a linear speed-up and recovering the rates of GD when $\sigma=0$ and the gradients of users grow at most as the global gradient.
3.3 Strongly Convex Costs
In this subsection we provide improved results for the last iterate of strongly convex costs. It is well known that strong convexity implies a unique global minimizer of $f$, see, e.g., [74]. Denote the global minimizer by $x^{\star}\in\mathbb{R}^{d}$, noting that $f^{\star}\coloneqq f(x^{\star})$ and let $\kappa\coloneqq\frac{L}{\mu}\geq 1$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively. Next, recall that DSGD does not require a heterogeneity bound to converge in the MSE sense for strongly convex costs, e.g., [20, 11]. To show similar benefits in the HP sense, a different approach to Lemma 3 is required. This is achieved by carefully bounding the MGF of the consensus gap and leveraging properties of strong convexity. The formal result is stated next.
Lemma 4.
Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$
$\mathbb{E}\Big[\exp\Big(\nu K_{t+1}\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\Big)\Big]\leq\exp\bigg(\nu K_{t+1}\Big(\sum_{k\in[t]}\lambda^{t-k}S_{k}+\sum_{k\in[t]}\lambda^{t-k}D_{k}\Big)\bigg),$
where $S_{k}\coloneqq\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}\coloneqq\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{5}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{6}}\Big)$.
Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring bounded heterogeneity. Using Lemma 6 from the Appendix, it can be shown that the RHS is of the order $e^{\mathcal{O}(\nu K_{t+1}\alpha_{t}^{2})}$. We note that the requirement of same model initialization across users can be removed, at the cost of an additional term in the exponent on the RHS, of order $\mathcal{O}(\nu K_{t+1}\lambda^{t}\Delta_{x})$, which decays geometrically (recall that $\lambda\in[0,1)$). However, we make the assumption for ease of exposition. Prior to stating the main theorem, we provide an important technical result, which allows us to establish sharper bounds and ensure linear speed-up in the number of users is achieved.
Lemma 5.
Let $\{X^{t}\}_{t\geq 2}$ be a sequence of random variables initialized by a deterministic $X^{1}>0$, such that, for some $M\in\mathbb{N}$, $a,t_{0},C_{i}>0$, $i\in[M]$ and every $t\geq 1$
$\mathbb{E}[\exp(X^{t+1})]\leq\mathbb{E}\bigg[\hskip-1.99997pt\exp\hskip-1.99997pt\bigg(\hskip-3.00003pt\Big(1-\frac{a}{t+t_{0}}\Big)X^{t}+\hskip-3.99994pt\sum_{i\in[M]}\hskip-1.99997pt\frac{C_{i}}{(t+t_{0})^{i}}\hskip-1.99997pt\bigg)\hskip-1.99997pt\bigg].$
If $a\in(1,2]$ and $t_{0}\geq a$, we then have
$\mathbb{E}$ $[\exp(X^{t+1})]\leq\exp\bigg(\frac{(t_{0}+1)^{a}X_{1}}{(t+1+t_{0})^{a}}+\frac{2^{a}C_{1}}{a}+\frac{2^{a}C_{2}/(a-1)}{t+1+t_{0}}\bigg)$
$\times\exp\bigg(\frac{2^{a}C_{3}\log(t+1+t_{0})}{(t+1+t_{0})^{a}}+\sum_{j=4}^{M}\frac{2^{a}t_{0}^{3-j}C_{j}}{(j-3)(t+1+t_{0})^{a}}\bigg).$
Lemma 5 provides a bound on the MGF of an “almost decreasing” process. Compared to bounds used in the centralized setting, e.g., [31, 32, 40], Lemma 5 gives a more fine-grained result, allowing us to incorporate higher-order terms in the final rate. We are now ready to state the main result for strongly convex costs.
Theorem 2.
Let (A1)-(A4) and (A6) hold, let the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+d\sigma^{2}\kappa/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(T+t_{0})^{2}}\bigg),$
where $\mathcal{O}(\cdot)$ hides some higher-order terms.
Theorem 2 establishes HP convergence guarantees for smooth strongly convex costs and time-varying step-size, without requiring the bounded heterogeneity condition (A5). In the Appendix we provide the full bound, containing additional terms, of order $\mathcal{O}\big(\frac{\log(T+t_{0}))}{(T+t_{0})^{3}}\big)$. We can see that the leading term in Theorem 2 decays at a rate $\mathcal{O}\big(\frac{d\sigma^{2}\kappa}{n(T+t_{0})}\big)$, achieving linear speed-up, while the network, problem and heterogeneity dependence are captured through the higher-order term $\mathcal{O}\big(\frac{\lambda^{2}\kappa^{2}\|\nabla\mathbf{f}^{\star}\|^{2}}{(1-\lambda)^{2}n(T+t_{0})^{2}}\big)$. Crucially, the term without linear speed-up, i.e., $\mathcal{O}\big(\frac{\sigma^{2}}{(T+t_{0})^{2}}\big)$, achieves a faster decay, stemming from the improved result in Lemma 5. Compared to [57], where the authors establish the $\mathcal{O}\big(\frac{1}{T}\big)$ convergence rate in HP for PL costs, Theorem 2 provides several improvements. First, in [57] the authors require uniformly path-wise bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq C_{i}$, for every $t\geq 1$ almost surely, which is difficult to guarantee uniformly across all sample paths. Secondly, they impose the following condition on the MGF of the noise
$\mathbb{E}\left[\exp\bigg(\frac{\|z_{i}^{t}\|^{2}}{\alpha_{t}^{2}\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1),$
where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size. Such condition implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that the noise asymptotically vanishes, at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$. [^8] It can be easily verified via Markov’s inequality that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\big)\geq 1-\delta$, for any $\delta\in(0,1)$. Similarly, it follows from Jensen’s inequality that $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$. We remove both requirements, showing that HP convergence of the last iterate of strongly convex functions is guaranteed without bounded gradients and under standard sub-Gaussian noise, while also achieving linear speed-up in the number of users. Finally, we believe that the techniques used in our proof can be leveraged to show HP guarantees for the broader class of PL costs, which is left for future work.
4 Proof Outlines and Discussion
In this section we provide some proof sketches, outline the main challenges, and discuss our results.
Proof sketch of Theorem 1. Using Lemma 1, rearranging and summing up the first $T$ terms, we get
$\sum_{t\in[T]}\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{L^{2}}{2}\sum_{t\in[T]}\frac{\alpha_{t}}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (4)
We can use Lemma 3 to control the last term on the RHS. To deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$, while removing the need for uniformly bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [39, 33, 46], by subtracting $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of $\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ are absorbed by the left-hand side, allowing us to show that the resulting MGF is bounded. The rest of the proof then relies on Lemma 2, some technical results, careful selection of the step-size and Hölder’s inequality, see the Appendix for details.
Proof sketch of Theorem 2. Starting from Lemma 1, subtracting $f^{\star}$ from both sides, using properties of strongly convex functions, defining $F^{t}\coloneqq n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$ and $A_{t}\coloneqq(t+t_{0}+1)\alpha_{t}$, we get
$F^{t+1}$ $\leq(1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}F^{t}-A_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}A_{t}nL\|\overline{z}^{t}\|^{2}+\frac{A_{t}L^{2}}{2}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (5)
We utilize the above inequality, Lemma 4, and a careful analysis to show that the MGF of $F^{t+1}$ satisfies the condition outlined in Lemma 5, after which we can use Lemma 5 to bound the MGF of the optimality gap. To complete the proof, we use the inequality [^9] See the Appendix for details.
$f(x^{t}_{i})-f^{\star}\leq 2(f(\overline{x}^{t})-f^{\star})+\frac{L}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2},$
and apply our results to show that both optimality and consensus gaps are small, with high probability.
Discussion. Results of Theorems 1 and 2 are strong, demonstrating that DSGD converges in the HP sense under the same conditions on the cost function as in the MSE sense, while retaining order-optimal rates and linear speed-up in the number of users. Compared to works studying HP convergence of centralized SGD, e.g., [31, 32, 33, 40], we make several contributions. First, we face the challenge of controlling the MGF of the consensus gap, stemming from the decentralized nature of the algorithm. To that end, we provide Lemma 3 for non-convex costs, relying on the bounded heterogeneity condition, and the improved Lemma 4 for strongly convex costs, which removes the bounded heterogeneity condition. Next, we provide Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense, which is of independent interest for HP studies of decentralized methods. To ensure linear speed-up in the number of users is achieved for strongly convex costs, we establish Lemma 5, which gives a more fine-grained result on the MGF [31, 32, 40], facilitating higher-order terms in the final bound, and is of independent interest. Moreover, to be able to utilize Lemma 5, a careful analysis of the MGF is needed, in order to simultaneously balance the effects of optimality and consensus gaps (recall that (5), which is used to bound the MGF of $F^{t+1}$, consists of both optimality and consensus gaps), while also ensuring that the “almost decreasing” property is maintained.
On the dimension dependence. We note that our bounds show a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for either MSE or HP bounds, e.g., [11, 57]. This dependence stems from the third result in Lemma 2, where the problem dimension shows up in the sub-Gaussianity constant, which is unavoidable, due to the nature of the definition of sub-Gaussianity in (A4). In particular, to show the third property in Lemma 2, we use the fact that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-norm-sub-Gaussian, as a direct proof fails to yield the variance reduction benefit in the number of users. One way to avoid the (linear) dependence on $d$ is to assume that the noise is norm-sub-Gaussian, which would simultaneously relax the dependence on the dimension to $\log(d)$ and achieve the variance reduction benefit, at the cost of imposing a slightly stronger noise condition, see, e.g., Corollary 7 and the conclusion in [75]. [^10] See the Appendix for the full proof of Lemma 2 and further discussion on the dimension dependence.
5 Conclusion
In this paper we revisit convergence in HP of a variant of DSGD under sub-Gaussian noise. We show that DSGD is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates for both non-convex and strongly convex costs. Moreover, our results show DSGD achieves linear speed-up in the number of users in both cases. Compared to [57], wherein the authors study HP convergence of DSGD, we relax strong conditions like uniformly bounded gradients or asymptotically vanishing noise, while improving the rates by showing linear speed-up in the number of users. Compared to works studying HP convergence of centralized SGD, e.g., [31, 32, 33, 40], we provide tight control on the MGF of the consensus gap for both non-convex and strongly convex costs, in Lemmas 3 and 4, as well as a more fine-grained bound on the MGF of an “almost decreasing” process, in Lemma 5. Future work includes extending our results to costs satisfying the PL condition, incorporating the GT mechanism to remove bounded heterogeneity condition and considering noise with heavier tails, like sub-Weibull.
Appendix A Introduction
The Appendix contains results omitted from the main body. Section B collects some important facts used in our proofs, Section C provides some technical results, Section D defines some notions used in the analysis, Sections E and F provide proofs for non-convex and strongly convex costs, respectively, while Section G contains further discussion on dimension dependence in our bounds.
Appendix B Useful Inequalities
In this section we outline some well-known inequalities and results used in our proofs. We start with Jensen’s inequality for convex/concave functions.
Proposition 1 (Jensen’s inequality).
Let $X\in\mathbb{R}$ be an integrable random variable. Then, for any convex function $h:\mathbb{R}\mapsto\mathbb{R}$, we have
$h(\mathbb{E}[X])\leq\mathbb{E}[h(X)].$
Moreover, if $h$ is concave, the reverse inequality holds, i.e., we then have
$\mathbb{E}[h(X)]\leq h(\mathbb{E}[X]).$
Proposition 2 (Cauchy-Schwartz inequality).
For any $a,b\in\mathbb{R}^{d}$, we have
$|\langle a,b\rangle|\leq\|a\|\|b\|.$
As a consequence of the Cauchy-Schwartz inequality, we have the following result.
Proposition 3 (Young’s inequality).
For any $a,b\in\mathbb{R}$ and any $\epsilon>0$, we have
$ab\leq\frac{\epsilon a^{2}}{2}+\frac{b^{2}}{2\epsilon}.$
As a consequence, for any $\theta>0$, we have
$(a+b)^{2}\leq(1+\theta)a^{2}+(1+\theta^{-1})b^{2}.$
Young’s inequality is also known as the Peter-Paul inequality.
Proposition 4 (Hölder’s inequality).
For any random variables $X,Y\in\mathbb{R}$ and any $p,q\in[1,\infty]$, such that $\frac{1}{p}+\frac{1}{q}=1$, we have
$\mathbb{E}|XY|\leq\sqrt[p]{\mathbb{E}|X|^{p}}\sqrt[q]{\mathbb{E}|Y|^{q}}.$
The coefficients $p,q\in[1,\infty]$ are known as Hölder coefficients. Note that Hölder’s inequality recovers Cauchy-Schwartz inequality for $p=q=2$. We have the following useful consequence of Hölder’s inequality.
Proposition 5.
For any $n\in\mathbb{N}$ and random variables $X_{i}\in\mathbb{R}$, $i\in[n]$, we have
$\mathbb{E}\Big[\prod_{i\in[n]}|X_{i}|\Big]\leq\prod_{i\in[n]}\sqrt[n]{\mathbb{E}|X_{i}|^{n}}.$
Next, we state a useful result from [25].
Proposition 6 ([25], Lemma 21).
For any $c,t_{0}>0$ and $0\leq a\leq b$, we have
$\prod_{k=a}^{b}\Big(1-\frac{c}{t+t_{0}}\Big)\leq\frac{(a+t_{0})^{c}}{(b+1+t_{0})^{c}}.$
The following result states some important consequences of conditions (A2) and (A3), see, e.g., [76, 74].
Proposition 7.
Let (A3) hold. Then, for any $i\in[n]$ and $x,y\in\mathbb{R}^{d}$, the following statements are true.
  1. 1.
    $f_{i}(x)\leq f_{i}(y)+\langle\nabla f_{i}(y),x-y\rangle+\frac{L}{2}\|x-y\|^{2}.$
  2. 2.
    $f$ has $L$-Lipschitz continuous gradients.
  3. 3.
    $f(x)\leq f(y)+\langle\nabla f(y),x-y\rangle+\frac{L}{2}\|x-y\|^{2}.$
If in addition (A2) holds, then for any $x\in\mathbb{R}^{d}$, the following is true.
  1. 4.
    $\|\nabla f(x)\|^{2}\leq 2L(f(x)-f^{\star})$.
Finally, we complete this section with the following important consequences of (A6), see, e.g., [74].
Proposition 8.
Let (A6) hold. Then, the following are true, for all $i\in[n]$ and $x,y\in\mathbb{R}^{d}$.
  1. 1.
    $\langle y,\nabla^{2}f_{i}(x)y\rangle\|\geq\mu\|y\|^{2}$.
  2. 2.
    $f$ is $\mu$-strongly convex.
  3. 3.
    $\|\nabla f(x)\|^{2}\geq 2\mu\big(f(x)-f^{\star}\big)$.
  4. 4.
    $f(x)-f^{\star}\geq\frac{\mu}{2}\|x-x^{\star}\|^{2}$.
Appendix C Technical Results
In this section we prove Lemmas 2 and 5. We also provide another technical result used in our proofs. For the reader’s convenience, we restate Lemmas 2 and 5 below.
Lemma 2.
If (A3) holds, then the following are true for any $t\geq 1$, $i\in[n]$ and $\mathcal{F}_{t}$-measurable $v\in\mathbb{R}^{d}$.
  1. 1.
    $\mathbb{E}\left[\exp\left(\langle v,z^{t}_{i}\rangle\right)\>|\>\mathcal{F}_{t}\right]\leq\exp\left(\frac{3\sigma_{i}^{2}\|v\|^{2}}{4}\right)$.
  2. 2.
    $\mathbb{E}\left[\exp\left(\langle v,\overline{z}^{t}\rangle\right)\>|\>\mathcal{F}_{t}\right]\leq\exp\left(\frac{3\sigma^{2}\|v\|^{2}}{4n}\right)$.
  3. 3.
    The average noise $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian, i.e., $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{120d\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1)$.
Proof.
  1. 1.
    To prove the first property, we follow steps similar to those in [39, Lemma 1]. Let $y_{i}=\frac{z^{t}_{i}}{\sigma_{i}}$ and note that $\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\leq\exp(1)$, from (A3). Assume first that $v\in\mathbb{R}^{d}$ is such that $\|v\|\leq\frac{4}{3}$. Using the inequality $\exp(a)\leq a+\exp(\nicefrac{{9a^{2}}}{{16}})$, which holds for any $a\in\mathbb{R}$, we then have
    $\mathbb{E}\left[\exp\left(\langle v,y_{i}\rangle\right)\>|\>\mathcal{F}_{t}\right]$ $\leq\mathbb{E}\left[\langle v,y_{i}\rangle+\exp\left(\frac{9\langle v,y_{i}\rangle^{2}}{16}\right)\>\big|\>\mathcal{F}_{t}\right]\stackrel{{\scriptstyle(a)}}{{=}}\mathbb{E}\left[\exp\left(\frac{9\langle v,y_{i}\rangle^{2}}{16}\right)\>\big|\>\mathcal{F}_{t}\right]$
    $\stackrel{{\scriptstyle(b)}}{{\leq}}\mathbb{E}\left[\exp\left(\frac{9\|v\|^{2}\|y_{i}\|^{2}}{16}\right)\>\big|\>\mathcal{F}_{t}\right]\stackrel{{\scriptstyle(c)}}{{\leq}}\left(\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\right)^{\nicefrac{{9\|v\|^{2}}}{{16}}}$
    $\stackrel{{\scriptstyle(d)}}{{\leq}}\exp\left(\frac{9\|v\|^{2}}{16}\right)\leq\exp\left(\frac{3\|v\|^{2}}{4}\right),$
    where $(a)$ follows from (A3) and the fact that $v$ is $\mathcal{F}_{t}$-measurable, $(b)$ follows from Proposition 2, $(c)$ follows from the fact that $\frac{9\|v\|^{2}}{16}\leq 1$ and Proposition 1, while $(d)$ follows from $\mathbb{E}[\exp(\|y_{i}\|^{2})\>|\>\mathcal{F}_{t}]\leq\exp(1)$. Next, if $\|v\|>\frac{4}{3}$, we have
    $\mathbb{E}[\exp(\langle v,y_{i}\rangle)\>|\>\mathcal{F}_{t}]$ $\leq\exp\left(\frac{3\|v\|^{2}}{8}\right)\mathbb{E}\left[\exp\left(\frac{2\|y_{i}\|^{2}}{3}\right)\>\big|\>\mathcal{F}_{t}\right]$
    $\leq\exp\left(\frac{2}{3}+\frac{3\|v\|^{2}}{8}\right)\leq\exp\left(\frac{3\|v\|^{2}}{4}\right),$
    where the first inequality follows by applying Proposition 3 with $\epsilon=\frac{4}{3}$ and the fact that $v$ is $\mathcal{F}_{t}$-measurable, the second follows from Proposition 1 and $\mathbb{E}[\exp(\|y_{i}\|^{2})\>|\>\mathcal{F}_{t}]\leq\exp(1)$, while the third inequality follows from the fact that $\frac{2}{3}<\frac{3\|v\|^{2}}{8}$, since $\|v\|>\frac{4}{3}$. Combining both cases, we get $\mathbb{E}[\exp\left(\langle v,y_{i}\rangle\right)\>|\>\mathcal{F}_{t}]\leq\exp\left(\frac{3\|v\|^{2}}{4}\right)$, for any $\mathcal{F}_{t}$-measurable vector $v\in\mathbb{R}^{d}$. The proof is completed by applying this inequality to $\langle v,z^{t}_{i}\rangle=\langle\sigma_{i}v,y_{i}\rangle$.
  2. 2.
    Recall that $\overline{z}^{t}=\frac{1}{n}\sum_{i\in[n]}z^{t}_{i}$ and $\sigma^{2}=\frac{1}{n}\sum_{i\in[n]}\sigma_{i}^{2}$. We then have, for any $\mathcal{F}_{t}$-measurable $v\in\mathbb{R}^{d}$
    $\mathbb{E}[\exp(\langle v,\overline{z}^{t}\rangle]$ $=\mathbb{E}\bigg[\exp\bigg(\frac{1}{n}\sum_{i\in[n]}\langle v,z^{t}_{i}\rangle\bigg)\bigg]\stackrel{{\scriptstyle(a)}}{{=}}\prod_{i\in[n]}\mathbb{E}\bigg[\exp\bigg(\Big\langle\frac{v}{n},z^{t}_{i}\Big\rangle\bigg)\bigg]$
    $\stackrel{{\scriptstyle(b)}}{{\leq}}\prod_{i\in[n]}\exp\bigg(\frac{3\sigma_{i}^{2}\|v\|^{2}}{4n^{2}}\bigg)\stackrel{{\scriptstyle(c)}}{{=}}\exp\bigg(\frac{3\sigma^{2}\|v\|^{2}}{4n}\bigg),$
    where $(a)$ follows from the fact that, conditioned on $\mathcal{F}_{t}$, the noise across users is independent (recall (A3)), $(b)$ follows from the first part of the proof, while $(c)$ follows from the definition of $\sigma^{2}$.
  3. 3.
    Combining the previous result with Lemma 1 from [75], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
    $\mathbb{P}\big(\|\overline{z}^{t}\|>\epsilon\big)\leq 2\exp\bigg(-\frac{n\epsilon^{2}}{24\sigma^{2}d}\bigg).$
    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., [77, Proposition 2.5.2], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
We next restate Lemma 5.
Lemma 5.
Let $\{X^{t}\}_{t\geq 2}$ be a sequence of random variables initialized by a deterministic $X^{1}>0$, such that, for some $M\in\mathbb{N}$, $a,t_{0},C_{i}>0$, $i\in[M]$ and every $t\geq 1$
$\mathbb{E}[\exp(X^{t+1})]\leq\mathbb{E}\bigg[\exp\bigg(\Big(1-\frac{a}{t+t_{0}}\Big)X^{t}+\sum_{i\in[M]}\frac{C_{i}}{(t+t_{0})^{i}}\bigg)\bigg].$ (6)
If $a\in(1,2]$ and $t_{0}\geq a$, we then have
$\mathbb{E}[\exp(X^{t+1})]$ $\leq\exp\bigg(\frac{(t_{0}+1)^{a}X_{1}}{(t+1+t_{0})^{a}}+\frac{2^{a}C_{1}}{a}+\frac{2^{a}C_{3}\log(t+1+t_{0})}{(t+1+t_{0})^{a}}\bigg)$
$\times\exp\bigg(\frac{2^{a}C_{2}/(a-1)}{t+1+t_{0}}+\sum_{j=4}^{M}\frac{2^{a}t_{0}^{3-j}C_{j}}{(j-3)(t+1+t_{0})^{a}}\bigg).$
Proof.
Starting from (6), taking the logarithm, defining $Y_{t}\coloneqq\log\mathbb{E}[\exp(X^{t})]$ and $b_{t}=1-\frac{a}{t+t_{0}}$, we then have
$Y^{t+1}\leq\sum_{i\in[M]}\frac{C_{i}}{(t+t_{0})^{i}}+\log\mathbb{E}[\exp(b_{t}X^{t})]$ $\leq\sum_{i\in[M]}\frac{C_{i}}{(t+t_{0})^{i}}+\log\bigg[\Big(\mathbb{E}[\exp(X^{t})]\Big)^{b_{t}}\bigg]$
$=b_{t}Y^{t}+\sum_{i\in[M]}\frac{C_{i}}{(t+t_{0})^{i}},$ (7)
where the second inequality follows from the fact that $b_{t}\in(0,1)$ and Proposition 1. Unrolling the recursion (C) and noting that $Y_{1}=X_{1}$, since $X_{1}>0$ is deterministic, we get
$Y^{t+1}$ $\leq X^{1}\prod_{k\in[t]}b_{k}+\sum_{i\in[M]}C_{i}\sum_{k\in[t]}\frac{1}{(k+t_{0})^{i}}\prod_{s=k+1}^{t}b_{s}$
$\leq\frac{(t_{0}+1)^{a}X^{1}}{(t+1+t_{0})^{a}}+\sum_{i\in[M]}C_{i}\sum_{k\in[t]}\frac{1}{(k+t_{0})^{i}}\times\frac{(k+1+t_{0})^{a}}{(t+1+t_{0})^{a}}$
$\leq\frac{(t_{0}+1)^{a}X^{1}}{(t+1+t_{0})^{a}}+\sum_{i\in[M]}\frac{2^{a}C_{i}}{(t+1+t_{0})^{a}}\sum_{k\in[t]}(k+t_{0})^{a-i},$ (8)
where the second inequality follows from Proposition 6, while the third inequality follows from the fact that $\Big(\frac{k+1+t_{0}}{k+t_{0}}\Big)^{a}\leq 2^{a}$. We now proceed to analyze $\sum_{k\in[t]}(k+t_{0})^{a-i}$, for different values of $i\in[M]$. Using Darboux sums and the fact that $a\in(1,2]$, it can be readily verified that
$\sum_{k\in[t]}(k+t_{0})^{a-i}\leq\begin{cases}\frac{(t+1+t_{0})^{a}}{a},&i=1\\ \frac{(t+1+t_{0})^{a-1}}{a-1},&i=2\\ \ln(t+t_{0}+1),&i=3\\ \frac{t_{0}^{3-i}}{i-3},&i\geq 4\end{cases}.$ (9)
Plugging (9) into (C), we get
$Y^{t+1}$ $\leq\frac{(t_{0}+1)^{a}X^{1}}{(t+1+t_{0})^{a}}+\frac{2^{a}C_{1}}{a}+\frac{2^{a}C_{2}}{(a-1)(t+t_{0}+1)}$
$+\frac{2^{a}C_{3}\log(t+t_{0}+1)}{(t+1+t_{0})^{a}}+\sum_{j=4}^{M}\frac{2^{a}t_{0}^{3-j}C_{j}}{(j-3)(t+1+t_{0})^{a}}.$
Taking the exponent on both sides completes the proof. ∎
To complete this section, we provide another technical result used in our proofs.
Lemma 6.
Let $\lambda\in[0,1)$ and $\alpha_{t}=\frac{a}{(t+t_{0})^{c}}$, where $a,t_{0}>0$ and $c\geq\nicefrac{{1}}{{2}}$. If $t_{0}\geq\frac{2c-1+\lambda}{1-\lambda}$, then for any $t\geq 1$, we have
$\sum_{k\in[t]}\alpha_{k}\lambda^{t-k}\leq\frac{3\alpha_{t}}{1-\lambda}.$
Proof.
Using the definition of $\alpha_{t}$, we note that
$\sum_{k\in[t]}\alpha_{k}\lambda^{t-k}=\alpha_{t}\sum_{k\in[t]}\frac{\alpha_{k}}{\alpha_{t}}\lambda^{t-k}=\alpha_{t}\sum_{k\in[t]}\lambda^{t-k}\bigg(\frac{t+t_{0}}{k+t_{0}}\bigg)^{c}=\alpha_{t}\sum_{k\in[t]}\lambda^{t-k}\bigg(1+\frac{t-k}{k+t_{0}}\bigg)^{c}.$
Next, denote by $\widetilde{\lambda}\coloneqq 1-\lambda\in(0,1]$ and use the substitution s=tks=t-k, to get
$\sum_{k\in[t]}\alpha_{k}\lambda^{t-k}$ $=\alpha_{t}\sum_{s=0}^{t-1}(1-\widetilde{\lambda})^{s}\bigg(1+\frac{s}{t-s+t_{0}}\bigg)^{c}\leq\alpha_{t}\sum_{s=0}^{t-1}\exp\bigg(-\widetilde{\lambda}s+\frac{cs}{t-s+t_{0}}\bigg)$
$\leq\alpha_{t}\sum_{s=0}^{t-1}\exp\bigg(-s\bigg(\widetilde{\lambda}-\frac{c}{1+t_{0}}\bigg)\bigg)\leq\alpha_{t}\sum_{s=0}^{t-1}\exp\bigg(-\frac{\widetilde{\lambda}s}{2}\bigg),$ (10)
where we used $1\pm x\leq\exp(\pm x)$ in the first, the fact that $t-s\geq 1$ in the second and the choice of $t_{0}$ in the third inequality. Next, we use Darboux sums, to get
$\sum_{s=0}^{t-1}\exp\bigg(-\frac{\widetilde{\lambda}s}{2}\bigg)=1+\sum_{s=1}^{t-1}\exp\bigg(-\frac{\widetilde{\lambda}s}{2}\bigg)\leq 1+\int_{0}^{t-1}\exp\bigg(-\frac{\widetilde{\lambda}s}{2}\bigg)ds\leq 1+\frac{2}{\widetilde{\lambda}}\leq\frac{3}{\widetilde{\lambda}}$. (11)
Plugging (11) into (C) completes the proof. ∎
Appendix D Analysis Setup
In this section we define some notation useful for the analysis. To begin, let $g^{t}_{i}\coloneqq\nabla\ell(x^{t}_{i};\xi_{i}^{t})$ denote the stochastic gradient of user $i$ at time $t$. We can then represent the update rule (3) as
$x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\big(x^{t}_{i}-\alpha_{t}g^{t}_{i}\big).$ (12)
It can immediately be seen that $z^{t}_{i}=g^{t}_{i}-\nabla f_{i}(x^{t}_{i})$. Next, define the network average stochastic gradient $\overline{g}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}g^{t}_{i}$ and $\overline{\nabla}f_{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}\nabla f_{i}(x^{t}_{i})$, which represents the network average of user’s gradients evaluated at their local models. It can then be seen that $\overline{g}^{t}=\overline{\nabla}f_{t}+\overline{z}^{t}$. Combining the definition of the network average model, the fact that the weight matrix is doubly stochastic and (12), it follows that $\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t}$.
We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [13, 18, 25]. Let ${\mathbf{x}}^{t}\coloneq col(x_{1}^{t},\ldots,x_{n}^{t})\in\mathbb{R}^{nd}$ denote the column vector stacking users’ local models. Using this notation, we can then represent the update rule (12) compactly as
${\mathbf{x}}^{t+1}=\mathbf{W}({\mathbf{x}}^{t}-\alpha_{t}{\mathbf{g}}^{t}),$ (13)
where $\mathbf{W}=W\otimes I_{d}\in\mathbb{R}^{nd\times nd}$, $\otimes$ denotes the Kronecker product and $I_{d}\in\mathbb{R}^{d\times d}$ denotes the $d$-dimensional identity matrix, while ${\mathbf{g}}^{t}\coloneqq col(g^{t}_{1},\ldots,g^{t}_{n})$. Next, define the matrix $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$, where $\mathbf{1}_{n}\in\mathbb{R}^{n}$ is the vector of all ones. The matrix $J$ represent the “ideal” consensus matrix, where all users can communicate with each other. [^11] This is equivalent to the client-server setup in terms of the update rule, as the server averages models from all clients in each iteration. The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [18, 25] and references therein.
Proposition 9.
Let (A1) hold. Then, the following are true.
  1. 1.
    $W\mathbf{1}_{n}=J\mathbf{1}_{n}=\mathbf{1}_{n}$.
  2. 2.
    $\|W-J\|=\lambda$, where $\lambda\in[0,1)$ is the second largest singular value of $W$.
  3. 3.
    $WJ=JW=J$.
Next, define $\overline{{\mathbf{x}}}^{t}\coloneqq\mathbf{1}_{n}\otimes\overline{x}^{t}\in\mathbb{R}^{nd}$, $\mathbf{J}\coloneqq J\otimes I_{d}\in\mathbb{R}^{nd\times nd}$ and note that $\overline{{\mathbf{x}}}^{t}=\mathbf{J}{\mathbf{x}^{t}}$. Combined with (13), it follows that $\overline{{\mathbf{x}}}^{t+1}=\overline{{\mathbf{x}}}^{t}-\alpha_{t}\overline{{\mathbf{g}}}^{t}$. Denoting by $\widetilde{\mathbf{W}}\coloneqq\mathbf{W}-\mathbf{J}$, it follows from (13) that
${\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}=\mathbf{W}({\mathbf{x}^{t}}-\alpha{\mathbf{g}^{t}})-\mathbf{J}({\mathbf{x}^{t}}-\alpha_{t}{\mathbf{g}^{t}})=\widetilde{\mathbf{W}}({\mathbf{x}^{t}}-\alpha{\mathbf{g}^{t}})=\widetilde{\mathbf{W}}({\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}-\alpha{\mathbf{g}^{t}}).$ (14)
Finally, recall that, for strongly convex costs, we denote the unique global minima by $x^{\star}\in\mathbb{R}^{d}$. We define two related concepts, namely the column stacking of $x^{\star}$, i.e., ${\mathbf{x}^{\star}}\coloneqq\mathbf{1}_{n}\otimes x^{\star}$ and the stacking of users’ gradients evaluated at the global optima, i.e., $\nabla\mathbf{f}^{\star}\coloneqq col(\nabla f_{1}(x^{\star}),\ldots,\nabla f_{n}(x^{\star}))$. Note that $\|\nabla\mathbf{f}^{\star}\|^{2}=\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ is a useful measure of heterogeneity.
Appendix E Proofs for Non-convex Costs
In this section we prove Lemmas 1 and 3, as well as Theorem 1. For the reader’s convenience, we restate the results below, starting with Lemma 1.
Lemma 1.
Let (A3) hold. If $\alpha_{t}\leq\frac{1}{2L}$, we have
$f(\overline{x}^{t+1})\leq f(\overline{x}^{t})-\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}^{2}L\|\overline{z}^{t}\|^{2}+\frac{\alpha_{t}L^{2}}{2n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$
Proof.
The proof follows similar steps as in [25, Lemma 3]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
$f(x)\leq f(y)+\langle\nabla f(y),x-y\rangle+\frac{L}{2}\|x-y\|^{2}.$ (15)
Setting $x=\overline{x}^{t+1}$ and $y=\overline{x}^{t}$ in (15), we get
$f(\overline{x}^{t+1})$ $\leq f(\overline{x}^{t})-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{g}_{t}\rangle+\frac{\alpha_{t}^{2}L}{2}\|\overline{g}_{t}\|^{2}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}f(\overline{x}^{t})-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{\nabla}f_{t}\rangle-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}^{2}L\|\overline{\nabla}f_{t}\|^{2}+\alpha_{t}^{2}L\|\overline{z}^{t}\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{=}}f(\overline{x}^{t})-\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}-(1-2\alpha_{t}L)\frac{\alpha_{t}}{2}\|\overline{\nabla}f_{t}\|^{2}$ (16)
$\quad\quad\quad\quad+\frac{\alpha_{t}}{2}\|\overline{\nabla}f_{t}-\nabla f(\overline{x}^{t})\|^{2}-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}^{2}L\|\overline{z}^{t}\|^{2}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}f(\overline{x}^{t})-\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}+\frac{\alpha_{t}}{2}\|\overline{\nabla}f_{t}-\nabla f(\overline{x}^{t})\|^{2}-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}^{2}L\|\overline{z}^{t}\|^{2},$ (17)
where $(a)$ follows by applying Proposition 3 with $\theta=1$, $(b)$ follows from the identity $\langle a,b\rangle=\frac{1}{2}\big(\|a\|^{2}+\|b\|^{2}-\|a-b\|^{2}\big)$, while $(c)$ follows from the fact that $\alpha_{t}\leq\frac{1}{2L}$. Recalling the definition of $\overline{\nabla}f_{t}$, we get
$\|\overline{\nabla}f_{t}-\nabla f(\overline{x}^{t})\|^{2}=\Big\|\frac{1}{n}\sum_{i\in[n]}\big[\nabla f_{i}(x^{t}_{i})-\nabla f_{i}(\overline{x}^{t})\big]\Big\|^{2}\leq\frac{L^{2}}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2},$ (18)
where we used Proposition 1 and (A3) in the last inequality. Plugging (18) in (E) gives the desired result. ∎
Next, we restate Lemma 3.
Lemma 3.
Let (A1) and (A5) hold. We then have, for any $t\geq 1$
$\frac{1}{n}\sum_{i\in[n]}$ $\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\leq 2\lambda^{2t}\Delta_{x}+\frac{4\lambda^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\|z^{k}_{i}\|^{2}$
$+\frac{4\lambda^{2}A^{2}}{1-\lambda}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}+\frac{4\lambda^{2}B^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\|\nabla f(x^{k}_{i})\|^{2}.$
Proof.
We start by noting that $\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}=\|{\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}\|^{2}$. Next, starting from (14) and unrolling the recursion, we get
${\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}=\widetilde{\mathbf{W}}({\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t})-\alpha_{t}\widetilde{\mathbf{W}}{\mathbf{g}^{t}}=\ldots=\widetilde{\mathbf{W}}^{t}({\mathbf{x}}^{1}-\overline{{\mathbf{x}}}^{1})-\sum_{k\in[t]}\alpha_{k}\widetilde{\mathbf{W}}^{t+1-k}{\mathbf{g}^{k}},$
Using Proposition 9, it follows that
$\|{\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}\|\leq\|\widetilde{\mathbf{W}}^{t}({\mathbf{x}}^{1}-\overline{{\mathbf{x}}}^{1})\|+\sum_{k\in[t]}\alpha_{k}\|\widetilde{\mathbf{W}}^{t+1-k}{\mathbf{g}^{k}}\|\leq\lambda^{t}\|{\mathbf{x}}^{1}-\overline{{\mathbf{x}}}^{1}\|+\lambda\sum_{k\in[t]}\alpha_{k}\lambda^{t-k}\|{\mathbf{g}^{k}}\|,$
where we used the fact that $\|\widetilde{\mathbf{W}}^{k}\|=\|\widetilde{\mathbf{W}}\|^{k}$, for any integer $k\geq 0$. This readily implies
$\|{\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}\|^{2}$ $\leq 2\lambda^{2t}\|{\mathbf{x}}^{1}-\overline{{\mathbf{x}}}^{1}\|^{2}+2\lambda^{2}\left(\sum_{k\in[t]}\alpha_{k}\lambda^{t-k}\|{\mathbf{g}}_{k}\|\right)^{2}$
$\leq 2\lambda^{2t}\|{\mathbf{x}}^{1}-\overline{{\mathbf{x}}}^{1}\|^{2}+2\lambda^{2}\sum_{s\in[t]}\lambda^{t-s}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\|{\mathbf{g}^{k}}\|^{2},$ (19)
where in the second inequality we use Proposition 2, with $a=[a_{1},\ldots,a_{t}]^{\top}$ and $b=[b_{1},\ldots,b_{t}]^{\top}$, setting $a_{k}=\lambda^{(t-k)/2}$ and $b_{k}=\alpha_{k}\lambda^{(t-k)/2}\|{\mathbf{g}^{k}}\|$. Next, consider the quantity $\|{\mathbf{g}^{k}}\|^{2}$. Using the fact that $\|{\mathbf{g}^{k}}\|^{2}=\sum_{i\in[n]}\|g^{k}_{i}\|^{2}$, we get
$\|{\mathbf{g}^{k}}\|^{2}$ $=\sum_{i\in[n]}\|g^{k}_{i}\|^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}2\sum_{i\in[n]}\|z^{k}_{i}\|^{2}+2\sum_{i\in[n]}\|\nabla f_{i}(x^{k}_{i})\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2\sum_{i\in[n]}\|z^{k}_{i}\|^{2}+2nA^{2}+2B^{2}\sum_{i\in[n]}\|\nabla f(x^{k}_{i})\|^{2},$ (20)
where $(a)$ follows from $g^{k}_{i}=\nabla f_{i}(x^{k}_{i})+z^{k}_{i}$ and using Proposition 3 with $\theta=1$, while $(b)$ follows from (A5). Plugging (20) into (19), using the fact that $\sum_{k\in[t]}\lambda^{t-k}=\sum_{k=0}^{t-1}\lambda^{k}\leq\frac{1}{1-\lambda}$, the desired result follows. ∎
We are now ready to prove Theorem 1, which we restate next, for convenience.
Theorem 1.
Let (A1)-(A5) hold. If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$, where $C>0$ is a problem related constant satisfying
$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
we then have, for any $\delta\in(0,1)$, with probability at least $1-\delta$
$\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+L\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}[\Delta_{x}+n\lambda^{2}(A^{2}+\sigma^{2})]}{(1-\lambda^{2})T}\Bigg).$
Proof.
Using Lemma 1, rearranging and summing up the first $T$ terms, we get
$\sum_{t\in[T]}\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{L^{2}}{2n}\sum_{t\in[T]}\alpha_{t}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$
Next, to offset the effect of the inner product term, we subtract $\frac{9\sigma^{2}}{4n}\sum_{t\in[T]}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}$ from both sides of the above inequality and note that, choosing $\alpha_{t}\leq\frac{n}{9\sigma^{2}}$, we have $\frac{\alpha_{t}}{2}-\frac{9\alpha_{t}^{2}\sigma^{2}}{4n}\geq\frac{\alpha_{t}}{4}$. Therefore, we obtain
$\sum_{t\in[T]}\frac{\alpha_{t}}{4}\|\nabla f(\overline{x}^{t})\|^{2}$ $\leq\Delta_{f}+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{L^{2}}{2n}\sum_{t\in[T]}\alpha_{t}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}$
$-\sum_{t\in[T]}\alpha_{t}\Big(\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\frac{9\alpha_{t}\sigma^{2}}{4n}\|\nabla f(\overline{x}^{t})\|^{2}\Big).$ (21)
Using Proposition 3 with $\theta=1$ and Lipschitz continuity of gradients of $f$, it readily follows that
$\|\nabla f(x^{t}_{i})\|^{2}\leq 2\|\nabla f(\overline{x}^{t})\|^{2}+2L^{2}\|\overline{x}^{t}-x^{t}_{i}\|^{2},$
implying
$\|\nabla f(\overline{x}^{t})\|^{2}\geq\frac{1}{2n}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}-\frac{L^{2}}{n}\sum_{i\in[n]}\|\overline{x}^{t}-x^{t}_{i}\|^{2}.$ (22)
Plugging (22) into (21) and rearranging, we get
$\frac{1}{8n}\sum_{t\in[T]}\sum_{i\in[n]}\alpha_{t}\|\nabla f(x^{t}_{i})\|^{2}$ $\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\Big(\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\frac{9\alpha_{t}\sigma^{2}}{4n}\|\nabla f(\overline{x}^{t})\|^{2}\Big)$
$+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{3L^{2}}{4n}\sum_{t\in[T]}\alpha_{t}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$
Using Lemma 3, we then have
$\frac{1}{8n}\sum_{t\in[T]}\sum_{i\in[n]}\alpha_{t}\|\nabla f(x^{t}_{i})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\Big(\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\frac{9\alpha_{t}\sigma^{2}}{4n}\|\nabla f(\overline{x}^{t})\|^{2}\Big)+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}$
$\>+\frac{3L^{2}}{4}\sum_{t\in[T]}\alpha_{t}\Big(2\lambda^{2(t-1)}\Delta_{x}+\frac{4\lambda^{2}}{n(1-\lambda)}\sum_{k\in[t-1]}\alpha_{k}^{2}\lambda^{t-1-k}\sum_{i\in[n]}\Big[\|z^{k}_{i}\|^{2}+A^{2}+B^{2}\|\nabla f(x^{k}_{i})\|^{2}\Big]\Big).$ (23)
Next, note that, for any sequence $\{a_{t}\}_{t\in[T]}$, the following identity holds
$\sum_{t\in[T]}\sum_{k\in[t-1]}\lambda^{t-1-k}a_{k}=\sum_{t\in[T]}a_{t}\sum_{k\in[T-t]}\lambda^{k-1},$
implying that $\sum_{t\in[T]}\sum_{k\in[t-1]}\lambda^{t-1-k}a_{k}\leq\frac{1}{1-\lambda}\sum_{t\in[T]}a_{t}$, if $\{a_{t}\}_{t\in[T]}$ is non-negative. Using the fact that the step-size is non-increasing, and applying the previously stated relation to $\sum_{t\in[T]}\alpha_{t}\sum_{k\in[t-1]}\alpha_{k}^{2}\lambda^{t-1-k}\big[\|z^{k}_{i}\|^{2}+A^{2}+B^{2}\|\nabla f(x^{k}_{i})\|^{2}\big]$, it follows that
$\sum_{t\in[T]}\alpha_{t}\sum_{k\in[t-1]}\alpha_{k}^{2}\lambda^{t-1-k}\Big[\|z^{k}_{i}\|^{2}+A^{2}+B^{2}\|\nabla f(x^{k}_{i})\|^{2}\Big]$
$\leq\sum_{t\in[T]}\sum_{k\in[t-1]}\alpha_{k}^{3}\lambda^{t-1-k}\Big[\|z^{k}_{i}\|^{2}+A^{2}+B^{2}\|\nabla f(x^{k}_{i})\|^{2}\Big]$
$\leq\frac{1}{1-\lambda}\sum_{t\in[T]}\alpha_{t}^{3}\Big[\|z^{t}_{i}\|^{2}+A^{2}+B^{2}\|\nabla f(x^{t}_{i})\|^{2}\Big].$ (24)
Plugging (E) into (23), we get
$\frac{1}{8n}$ $\sum_{t\in[T]}\sum_{i\in[n]}\alpha_{t}\|\nabla f(x^{t}_{i})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\Big(\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\frac{9\alpha_{t}\sigma^{2}}{4n}\|\nabla f(\overline{x}^{t})\|^{2}\Big)+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}$
$+\frac{3\Delta_{x}L^{2}}{2}\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}+\frac{3\lambda^{2}L^{2}}{n(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\sum_{i\in[n]}\Big[\|z^{t}_{i}\|^{2}+A^{2}+B^{2}\|\nabla f(x^{t}_{i})\|^{2}\Big].$
Rearranging and choosing $\alpha_{t}\leq\frac{1-\lambda}{\lambda LB\sqrt{48}}$, we get
$\frac{1}{16n}\sum_{t\in[T]}$ $\sum_{i\in[n]}\alpha_{t}\|\nabla f(x^{t}_{i})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\Big(\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\frac{9\alpha\sigma^{2}}{4n}\|\nabla f(\overline{x}^{t})\|^{2}\Big)+L\sum_{t\in[T]}\alpha^{2}_{t}\|\overline{z}^{t}\|^{2}$
$+\frac{3\Delta_{x}L^{2}}{2}\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}+\frac{3\lambda^{2}A^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}+\frac{3\lambda^{2}L^{2}}{n(1-\lambda)^{2}}\sum_{t\in[T]}\sum_{i\in[n]}\alpha_{t}^{3}\|z^{t}_{i}\|^{2}.$ (25)
Define the process $M_{T}\coloneqq\frac{1}{16n}\sum_{t\in[T]}\sum_{i\in[n]}\alpha_{t}\|\nabla f(x^{t}_{i})\|^{2}-\Delta_{f}-\frac{3\Delta_{x}L^{2}}{2}\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}-\frac{3\lambda^{2}A^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}$ and consider its’ moment generating function. Using (25), we then have
$\mathbb{E}[\exp(M_{T})]\leq\mathbb{E}\Bigg[\exp\bigg(\sum_{t\in[T]}\underbrace{-\alpha_{t}\Big(\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\frac{9\alpha_{t}\sigma^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}}_{b_{1,t}}\Big)\bigg)\exp\bigg(\sum_{t\in[T]}\underbrace{\alpha_{t}^{2}L\|\overline{z}^{t}\|^{2}}_{b_{2,t}}\bigg)$
$\times\exp\bigg(\sum_{t\in[T]}\underbrace{\frac{3\alpha_{t}^{3}\lambda^{2}L^{2}}{n(1-\lambda)^{2}}\sum_{i\in[n]}\|z^{t}_{i}\|^{2}}_{b_{3,t}}\bigg)\Bigg]\leq\sqrt[3]{\mathbb{E}[\exp(B_{1,T})]\mathbb{E}[\exp(B_{2,T})]\mathbb{E}[\exp(B_{3,T})]},$ (26)
where $B_{k,T}=3\sum_{t\in[T]}b_{1,t}$ and $B_{k,0}=0$ for all $k\in[3]$, while the last step follows by applying Proposition 5. We now analyze each quantity separately, starting with $\mathbb{E}[\exp(B_{1,T})]$. To that end, we have
$\mathbb{E}[\exp(B_{1,T})]=\mathbb{E}\Bigg[\exp\bigg(B_{1,T-1}-\frac{27\alpha_{T}^{2}\sigma^{2}\|\nabla f(\overline{x}^{T})\|^{2}}{4n}\bigg)\mathbb{E}\left[\exp\left(-3\alpha_{T}\langle\nabla f(\overline{x}^{T}),\overline{z}^{T}\rangle\right)\>|\>\mathcal{F}_{T}\right]\Bigg].$
Noting that $\nabla f(\overline{x}^{T})$ is $\mathcal{F}_{T}$-measurable and applying Lemma 2, we get
$\mathbb{E}[\exp(B_{1,T})]$ $\leq\mathbb{E}\Bigg[\exp\bigg(B_{1,T-1}-\frac{27\alpha_{T}^{2}\sigma^{2}\|\nabla f(\overline{x}^{T})\|^{2}}{4n}+\frac{27\alpha_{T}^{2}\sigma^{2}\|\nabla f(\overline{x}^{T})\|^{2}}{4n}\bigg)\Bigg]=\mathbb{E}[\exp(B_{1,T-1})].$
Unrolling the recursion, it follows that $\mathbb{E}[\exp(B_{1,T})]\leq 1$. Next, consider $\mathbb{E}[\exp(B_{2,T})]$. To that end, we have
$\mathbb{E}[\exp(B_{2,T})]$ $=\mathbb{E}\left[\exp(B_{2,T-1})\mathbb{E}\left[\exp\left(3\alpha_{T}^{2}L\|\overline{z}^{T}\|^{2}\right)\>|\>\mathcal{F}_{T}\right]\right]$
$=\mathbb{E}\left[\exp(B_{2,T-1})\mathbb{E}\left[\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\frac{n\|\overline{z}^{T}\|^{2}}{120\sigma^{2}d}\right)\>\Big|\>\mathcal{F}_{T}\right]\right]$
$\leq\mathbb{E}\left[\exp(B_{2,T-1})\left(\mathbb{E}\left[\exp\left(\frac{n\|\overline{z}^{T}\|^{2}}{120\sigma^{2}d}\right)\>\Big|\>\mathcal{F}_{T}\right]\right)^{\nicefrac{{360\alpha_{T}^{2}\sigma^{2}dL}}{{n}}}\right],$
where the last inequality follows by choosing $\alpha_{t}\leq\frac{\sqrt{n}}{6\sigma\sqrt{10dL}}$ and using Proposition 1. Using Lemma 2, we get
$\mathbb{E}[\exp(B_{2,T})]\leq\mathbb{E}\left[\exp(B_{2,T-1})\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\right)\right]=\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\right)\mathbb{E}[\exp(B_{2,T-1})].$
Unrolling the recursion, it follows that $\mathbb{E}[\exp(B_{2,T})]\leq\exp\left(\frac{3\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}\right)$. Finally, to bound $\mathbb{E}[\exp(B_{3,T})]$, we can proceed in the same way, using the conditional independence of noise across agents (recall (A4)) to note that, if $\alpha_{t}\leq\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2/3}L^{2/3}\sqrt[3]{9}}$, then $\mathbb{E}[\exp(B_{3,T})]\leq\exp\left(\frac{9\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\right)$. Combining everything, we get
$\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg).$ (27)
Using Markov’s inequality and (27), we get, for any $\epsilon>0$
$\mathbb{P}\left(M_{T}>\epsilon\right)\leq\exp(-\epsilon)\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(-\epsilon+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg),$
or equivalently, for any $\delta\in(0,1)$, with probability at least $1-\delta$
$M_{T}\leq\log(\nicefrac{{1}}{{\delta}})+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
Recalling the definition of $M_{T}$ and using the fact that the sequence of step-sizes in non-increasing we then have, with probability at least $1-\delta$
$\frac{\alpha_{T}}{16n}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $\leq\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+\frac{3\Delta_{x}L^{2}}{2}\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}$
$+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
Dividing both sides by $\frac{\alpha_{T}T}{16}$, it follows that, with probability at least $1-\delta$
$\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha_{T}T}+\frac{24\Delta_{x}L^{2}}{\alpha_{T}T}\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}$
$+\frac{1920\sigma^{2}dL}{n\alpha_{T}T}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{48\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\alpha_{T}T(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
Next, consider two regimes with respect to the time horizon $T$.
  1. 1.
    Known time horizon. In this case, we choose a fixed step-size $\alpha_{t}\equiv\alpha$, for all $t\in[T]$. Noticing that $\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}\leq\frac{\alpha}{1-\lambda^{2}}$, we then have, with probability at least $1-\delta$
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\alpha\sigma^{2}dL}{n}+\frac{48\alpha^{2}\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}.$
    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{dT}}$, then with probability at least $1-\delta$
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\sigma L\sqrt{d}}{\sqrt{nT}}+\frac{48n\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{dT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{dT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma}{\sqrt{nT}}+\frac{1}{CT}$. Therefore, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we finally have
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+120L\big)}{\sqrt{nT}}$
    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{48n\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
  2. 2.
    Unknown time horizon. In this case, we choose a time-varying step-size $\alpha_{t}=\frac{C^{\prime}}{\sqrt{t+1}}$, for all $t\geq 1$ and $C^{\prime}=\sqrt{2}C$, where $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$, again guaranteeing that all the step-size conditions are satisfied. Noting that $\alpha_{t}\leq C$, we get $\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}\leq\frac{C}{(1-\lambda^{2})}$, hence, with probability at least $1-\delta$
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $\leq\frac{16\sqrt{2}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{C\sqrt{T+1}}+\frac{24\sqrt{2}\Delta_{x}L^{2}}{C\sqrt{T+1}(1-\lambda^{2})}$
    $+\frac{3840\sqrt{2}\sigma^{2}dLC\log(T+1)}{n\sqrt{T+1}}+\frac{384\lambda^{2}L^{2}(\sigma^{2}+A^{2})C^{2}}{(1-\lambda)^{2}\sqrt{T+1}}.$
    Note that in the unknown time horizon case we lose the linear speed-up in the number of users, as, even in the case $C=\frac{\sqrt{n}}{6\sigma\sqrt{10dL}}$, we have
    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{T+1}}{{\delta}})\big)}{\sqrt{n(T+1)}}+\frac{\sigma\sqrt{d}\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
    as the last term, which does not attain a linear speed-up, is no longer of higher order. We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [11, 25].
Appendix F Proofs for Strongly Convex Costs
In this section we prove Lemma 4 and Theorem 2. To do so, we follow a similar strategy to the one in, e.g., [20, 25], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense. However, to establish guarantees in the HP sense, we instead work with the MGF of the iterates and have the following important result.
Lemma 7.
Let assumptions (A1)-(A4) and (A6) hold and let $a,t_{0},K>0$ and $\nu\in(0,1]$ be some positive constants. If the step-size satisfies $\alpha_{t}\leq\min\left\{\frac{1}{\overline{\sigma}\sqrt{2(t+t_{0}+2)K}},\frac{1}{\mu}\right\}$ for all $t\geq 1$, with $\nu\leq\min\left\{1,\frac{\mu}{24a\overline{\sigma}^{2}K}\right\}$ and $K_{t+1}=(t+t_{0}+2)K$, then
$\mathbb{E}[\exp(\nu K_{t+1}\|{\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}\|^{2})]\leq\exp\bigg(\nu K_{t+1}\bigg(\frac{4an\sigma^{2}\alpha_{t+1}}{a\mu-1}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{a\mu}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(t+1+t_{0})^{a\mu}}\bigg)\bigg).$
Proof.
Consider the update rule (13). We then have
${\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}=\mathbf{W}({\mathbf{x}^{t}}-{\mathbf{x}^{\star}}-\alpha_{t}{\mathbf{g}}_{t})=\mathbf{W}({\mathbf{x}^{t}}-{\mathbf{x}^{\star}}-\alpha_{t}\nabla\mathbf{f}^{t}-\alpha_{t}{\mathbf{z}^{t}}),$ (28)
where in the first equality we used $\mathbf{W}{\mathbf{x}^{\star}}={\mathbf{x}^{\star}}$. Using Taylor’s expansion, for each $i\in[n]$ and $x\in\mathbb{R}^{d}$, we have
$\nabla f_{i}(x)=\nabla f_{i}(x^{\star})+\int_{0}^{1}\nabla^{2}f_{i}(x^{\star}+\tau(x-x^{\star}))d\tau(x-x^{\star})=\nabla f_{i}(x^{\star})+H_{i}(x)(x-x^{\star}).$ (29)
Denote by $\mathbf{H}^{t}\coloneqq diag(H_{1}(x^{t}_{1}),\ldots,H_{n}(x^{t}_{n}))\in\mathbb{R}^{nd\times nd}$ the block diagonal matrix and recall that $\nabla\mathbf{f}^{\star}=col(\nabla f_{1}(x^{\star}),\ldots,\nabla f_{n}(x^{\star}))\in\mathbb{R}^{nd}$. Using (29), we can readily see that $\nabla\mathbf{f}^{t}=\nabla\mathbf{f}^{\star}+\mathbf{H}^{t}({\mathbf{x}^{t}}-{\mathbf{x}}^{\star})$, therefore, plugging in (28), we get
${\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}=\mathbf{W}(\mathbf{I}-\alpha_{t}\mathbf{H}^{t})({\mathbf{x}^{t}}-{\mathbf{x}^{\star}})-\alpha_{t}\mathbf{W}\nabla\mathbf{f}^{\star}-\alpha_{t}\mathbf{W}{\mathbf{z}^{t}}=C_{t}+\alpha_{t}\mathbf{W}{\mathbf{z}^{t}},$
where $C_{t}\coloneqq\mathbf{W}(\mathbf{I}-\alpha_{t}\mathbf{H}^{t})({\mathbf{x}^{t}}-{\mathbf{x}^{\star}})-\alpha_{t}\mathbf{W}\nabla\mathbf{f}^{\star}$. Therefore, we have
$\|{\mathbf{x}^{t+1}}$ $-{\mathbf{x}^{\star}}\|^{2}=\|C_{t}\|^{2}-2\alpha_{t}\langle\mathbf{W}C_{t},{\mathbf{z}^{t}}\rangle+\alpha_{t}^{2}\|\mathbf{W}{\mathbf{z}^{t}}\|^{2}\leq\|C_{t}\|^{2}-2\alpha_{t}\langle\mathbf{W}C_{t},{\mathbf{z}^{t}}\rangle+\alpha_{t}^{2}\|{\mathbf{z}^{t}}\|^{2}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}(1+\theta)\|\mathbf{W}(\mathbf{I}-\alpha_{t}\mathbf{H}^{t})({\mathbf{x}^{t}}-{\mathbf{x}^{\star}})\|^{2}+(1+\theta^{-1})\alpha_{t}^{2}\|\mathbf{W}\nabla\mathbf{f}^{\star}\|^{2}-2\alpha_{t}\langle\mathbf{W}C_{t},{\mathbf{z}^{t}}\rangle+\alpha_{t}^{2}\|{\mathbf{z}^{t}}\|^{2}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}(1+\theta)(1-\alpha_{t}\mu)^{2}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}+(1+\theta^{-1})\alpha_{t}^{2}\|\nabla\mathbf{f}^{\star}\|^{2}-2\alpha_{t}\langle\mathbf{W}C_{t},{\mathbf{z}^{t}}\rangle+\alpha_{t}^{2}\|{\mathbf{z}^{t}}\|^{2},$
where in $(i)$ we used Proposition 3, for some $\theta>0$ (to be specified later), while $(ii)$ follows from Proposition 9 and the fact that $\|\mathbf{I}-\alpha_{t}\mathbf{H}^{t}\|\leq(1-\alpha_{t}\mu)$ (as a consequence of Proposition 8). Define $D_{t}\coloneqq(1+\theta)(1-\alpha_{t}\mu)^{2}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}+(1+\theta^{-1})\alpha_{t}^{2}\|\nabla\mathbf{f}^{\star}\|^{2}$ and consider the MGF of $\nu K_{t+1}\|{\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}\|^{2}$ conditioned on $\mathcal{F}_{t}$, where we recall that $K_{t+1}=(t+t_{0}+2)K$ for some $K>0$ and $\nu\in(0,1]$. We then have
$\mathbb{E}_{t}[\exp(\nu K_{t+1}$ $\|{\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}\|^{2})]\stackrel{{\scriptstyle(a)}}{{\leq}}\exp(\nu K_{t+1}D_{t})\mathbb{E}_{t}\Big[\exp\Big(\nu K_{t+1}\big(-2\alpha_{t}\langle\mathbf{W}C_{t},{\mathbf{z}^{t}}\rangle+\alpha_{t}^{2}\|{\mathbf{z}^{t}}\|^{2}\big)\Big)\Big]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\exp(\nu K_{t+1}D_{t})\sqrt{\mathbb{E}_{t}[\exp(-4\alpha_{t}\nu K_{t+1}\langle\mathbf{W}C_{t},{\mathbf{z}^{t}}\rangle)\mathbb{E}_{t}[\exp(2\alpha_{t}^{2}\nu K_{t+1}\|{\mathbf{z}^{t}}\|^{2})]}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\exp(\nu K_{t+1}D_{t})\sqrt{\exp(12\alpha_{t}^{2}\overline{\sigma}^{2}\nu^{2}K_{t+1}^{2}\|\mathbf{W}C_{t}\|^{2}+2\alpha_{t}^{2}n\sigma^{2}\nu K_{t+1})}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\exp(\nu K_{t+1}D_{t}+6\alpha_{t}^{2}\overline{\sigma}^{2}\nu^{2}K_{t+1}^{2}\|C_{t}\|^{2}+\alpha_{t}^{2}n\sigma^{2}\nu K_{t+1})$
$\stackrel{{\scriptstyle(e)}}{{\leq}}\exp(\nu K_{t+1}(1+6\alpha_{t}^{2}\overline{\sigma}^{2}\nu K_{t+1})D_{t}+\alpha_{t}^{2}n\sigma^{2}\nu K_{t+1}),$ (30)
where $(a)$ follows from the fact that $D_{t}$ is $\mathcal{F}_{t}$-measurable, in $(b)$ we used Proposition 4, $(c)$ follows from Lemma 2 and $\alpha_{t}\leq\frac{1}{\overline{\sigma}\sqrt{2(t+t_{0}+2)K}}$, in $(d)$ we used Proposition 9, while $(e)$ follows from the definition of $D_{t}$ and the fact that $\|C_{t}\|^{2}\leq D_{t}$. We now analyze $(1+6\alpha_{t}^{2}\overline{\sigma}^{2}\nu K_{t+1})D_{t}$. To that end, if we choose $\theta=\frac{\alpha_{t}\mu}{2}$, it follows that
$(1+6\alpha_{t}^{2}\overline{\sigma}^{2}\nu K_{t+1})D_{t}$ $=(1+6\alpha_{t}^{2}\overline{\sigma}^{2}\nu K_{t+1})\big[(1+\nicefrac{{\alpha_{t}\mu}}{{2}})(1-\alpha_{t}\mu)^{2}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}+(1+\nicefrac{{2}}{{\alpha_{t}\mu}})\alpha_{t}^{2}\|\nabla\mathbf{f}^{\star}\|^{2}\big]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}(1+6\alpha_{t}^{2}\overline{\sigma}^{2}\nu K_{t+1})\big[(1-\nicefrac{{\alpha_{t}\mu}}{{2}})(1-\alpha_{t}\mu)\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}+(\alpha_{t}+\nicefrac{{2}}{{\mu}})\alpha_{t}\|\nabla\mathbf{f}^{\star}\|^{2}\big]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}(1-\alpha_{t}\mu)\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}+9\alpha_{t}\|\nabla\mathbf{f}^{\star}\|^{2}/2\mu,$ (31)
where in $(i)$ we used the fact that $(1+\frac{a}{2})(1-a)\leq(1-\frac{a}{2})$ for any $a>0$, while $(ii)$ follows by setting $\nu\leq\frac{\mu}{24a\overline{\sigma}^{2}K}$ and from the step-size choice $\alpha_{t}\leq\frac{1}{\mu}$. Plugging (F) in (F), using the shorthand $E_{t}=\alpha_{t}^{2}n\sigma^{2}+9\alpha_{t}\|\nabla\mathbf{f}^{\star}\|^{2}/2\mu$ and taking the full expectation, we get
$\mathbb{E}[\exp(\nu$ $K_{t+1}\|{\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}\|^{2})]\leq\exp(\nu K_{t+1}E_{t})\mathbb{E}[\exp((1-\alpha_{t}\mu)\nu K_{t+1}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2})]$
$\leq\exp(\nu K_{t+1}E_{t})\Big(\mathbb{E}[\exp(\nu K_{t}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2})]\Big)^{(1-\alpha_{t}\mu)\frac{t+t_{0}+2}{t+t_{0}+1}}$
$\leq\exp\big(\nu K_{t+1}\big(E_{t}+(1-\alpha_{t}\mu)E_{t-1}\big)\big)\Big(\mathbb{E}[\exp((1-\alpha_{t-1}\mu)\nu K_{t}\|{\mathbf{x}}_{t-1}-{\mathbf{x}^{\star}}\|^{2})]\Big)^{(1-\alpha_{t}\mu)\frac{t+t_{0}+2}{t+t_{0}+1}}$
$\leq\ldots\leq\exp\Big(\nu K_{t+1}\sum_{k=1}^{t}E_{k}\prod_{s=k+1}^{t}(1-\alpha_{k}\mu)+\nu K_{1}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}\prod_{k=1}^{t}(1-\alpha_{k}\mu)\frac{t+t_{0}+2}{t_{0}+2}\Big)$
$=\exp\bigg(\nu K_{t+1}\bigg(\sum_{k=1}^{t}E_{k}\prod_{s=k+1}^{t}(1-\alpha_{s}\mu)+\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}\prod_{k=1}^{t}(1-\alpha_{k}\mu)\bigg)\bigg),$
where the second inequality follows from Proposition 1 and the fact that $0<(1-\alpha_{t}\mu)\frac{t+t_{0}+2}{t+t_{0}+1}\leq 1$, for any $t\geq 1$, whenever $0<\alpha_{t}\leq\frac{1}{\mu}$. Next, we use Proposition 6, to get
$\sum_{k=1}^{t}E_{k}$ $\prod_{s=k+1}^{t}(1-\alpha_{s}\mu)\leq\sum_{k=1}^{t}\Big(\alpha_{k}^{2}n\sigma^{2}+9\alpha_{k}\|\nabla\mathbf{f}^{\star}\|^{2}/2\mu\Big)\frac{(k+1+t_{0})^{a\mu}}{(t+1+t_{0})^{a\mu}}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\frac{1}{(t+t_{0}+1)^{a\mu}}\sum_{k=1}^{t}\Big(4a^{2}n\sigma^{2}(k+1+t_{0})^{a\mu-2}+\frac{9a\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu}(k+t_{0}+1)^{a\mu-1}\Big)$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{4a^{2}n\sigma^{2}}{(a\mu-1)(t+t_{0}+1)}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}},$
where $(i)$ follows from the step-size choice $\alpha_{k}\leq\frac{1}{\mu}$, while in $(ii)$ we use the lower Darboux sum. Combining everything, we finally get
$\mathbb{E}[\exp(\nu K_{t+1}\|{\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}\|^{2})]\leq\exp\bigg(\nu K_{t+1}\bigg(\frac{4an\sigma^{2}\alpha_{t+1}}{a\mu-1}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{a\mu}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(t+t_{0}+1)^{a\mu}}\bigg)\bigg).$
Lemma 7 is an important building block for bounding the consensus gap. We next restate and prove Lemma 4.
Lemma 4.
Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$
$\mathbb{E}\Big[\exp\Big(\nu K_{t+1}\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\Big)\Big]\leq\exp\bigg(\nu K_{t+1}\Big(\sum_{k\in[t]}\lambda^{t-k}S_{k}+\sum_{k\in[t]}\lambda^{t-k}D_{k}\Big)\bigg),$
where $S_{k}=\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}=\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{5}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{6}}\Big)$.
Proof.
We start by noting that $\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}=\|{\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}\|^{2}$. Next, recall the update rule (13) and $\widetilde{\mathbf{W}}=\mathbf{W}-\mathbf{J}$. We then have
${\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}=\widetilde{\mathbf{W}}({\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}-\alpha_{t}\nabla\mathbf{f}^{t}-\alpha_{t}{\mathbf{z}^{t}}).$
Denote the consensus difference by $\widetilde{{\mathbf{x}}}^{t+1}\coloneqq{\mathbf{x}^{t+1}}-\overline{{\mathbf{x}}}^{t+1}$ and let $C_{t}\coloneqq\widetilde{\mathbf{W}}(\widetilde{{\mathbf{x}}}_{t}-\alpha_{t}\nabla\mathbf{f}^{t})$. Noting that $\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2}=\|C_{t}\|^{2}-2\alpha_{t}\langle\widetilde{\mathbf{W}}C_{t},{\mathbf{z}^{t}}\rangle+\alpha_{t}\|\widetilde{\mathbf{W}}{\mathbf{z}^{t}}\|^{2}$, we then consider the MGF of $\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2}$ conditioned on $\mathcal{F}_{t}$, where we recall that $K_{t+1}=(t+t_{0}+2)K$, for some $K>0$ and $\nu\in(0,1]$. If $\alpha_{t}\leq\frac{1}{\overline{\sigma}\lambda\sqrt{2(t+t_{0}+2)K}}$, we have
$\mathbb{E}_{t}[\exp(\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2})]$ $\leq\exp(\nu K_{t+1}\|C_{t}\|^{2})\sqrt{\mathbb{E}_{t}\big[\exp(-4\alpha_{t}\nu K_{t+1}\langle\widetilde{\mathbf{W}}C_{t},{\mathbf{z}^{t}}\rangle)\big]\mathbb{E}_{t}\big[\exp(2\alpha_{t}^{2}\lambda^{2}\nu K_{t+1}\|{\mathbf{z}^{t}}\|^{2})\big]}$
$\leq\exp\Big(\nu K_{t+1}\Big((1+6\alpha_{t}^{2}\overline{\sigma}^{2}\lambda^{2}\nu K_{t+1})\|C_{t}\|^{2}+\alpha_{t}^{2}\lambda^{2}n\sigma^{2}\Big)\Big),$ (32)
where the first inequality follows from the fact that $C_{t}$ is $\mathcal{F}_{t}$-measurable and using Proposition 4, while the second follows from Lemma 2. Next, using Proposition 9 and defining $\widetilde{\lambda}\coloneqq 1-\lambda\in(0,1]$, we get
$\|C_{t}\|^{2}$ $\leq\lambda^{2}\|\widetilde{{\mathbf{x}}}_{t}-\alpha_{t}\nabla\mathbf{f}^{t}\|^{2}\stackrel{{\scriptstyle(i)}}{{\leq}}\lambda^{2}(1+\theta)\|\widetilde{{\mathbf{x}}}_{t}\|^{2}+\alpha_{t}^{2}\lambda^{2}(1+\theta^{-1})\|\nabla\mathbf{f}^{t}\|^{2}$
$=(1-\widetilde{\lambda})(1+\theta)\lambda\|\widetilde{{\mathbf{x}}}_{t}\|^{2}+\alpha_{t}^{2}\lambda^{2}(1+\theta^{-1})\|\nabla\mathbf{f}^{t}\|^{2}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}(1-\nicefrac{{\widetilde{\lambda}}}{{2}})\lambda\|\widetilde{{\mathbf{x}}}_{t}\|^{2}+\alpha_{t}^{2}\lambda^{2}(1+\nicefrac{{2}}{{\widetilde{\lambda}}})\|\nabla\mathbf{f}^{t}\|^{2}$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}(1-\nicefrac{{\widetilde{\lambda}}}{{2}})\lambda\|\widetilde{{\mathbf{x}}}_{t}\|^{2}+\frac{6\alpha_{t}^{2}\lambda^{2}}{1-\lambda}\|\nabla\mathbf{f}^{\star}\|^{2}+\frac{6\alpha_{t}^{2}\lambda^{2}L^{2}}{1-\lambda}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2},$ (33)
where $(i)$ follows from Proposition 3, in $(ii)$ we set $\theta=\frac{\widetilde{\lambda}}{2}$, while $(iii)$ follows from the fact that $\|\nabla\mathbf{f}^{t}\|^{2}\leq 2L^{2}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|+2\|\nabla\mathbf{f}^{\star}\|^{2}$ (recall Proposition 7). Choosing $\alpha_{t}\leq\frac{\sqrt{1-\lambda}}{2\overline{\sigma}\lambda\sqrt{6(t+t_{0}+2)K}}$ and plugging (F) in (F), we get
$\mathbb{E}_{t}[\exp(\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2})]\leq\exp\bigg(\nu K_{t+1}\Big(1+\frac{\widetilde{\lambda}}{4}\Big)\Big[\lambda\Big(1-\frac{\widetilde{\lambda}}{2}\Big)\|\widetilde{{\mathbf{x}}}_{t}\|^{2}$
$\quad\quad\quad\quad+\frac{4\alpha_{t}^{2}\lambda^{2}}{1-\lambda}\|\nabla\mathbf{f}^{\star}\|^{2}+\frac{4\alpha_{t}^{2}\lambda^{2}L^{2}}{1-\lambda}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}\Big]+\alpha_{t}^{2}n\sigma^{2}\lambda^{2}\nu K_{t+1}\bigg)$
$\leq\exp\bigg(\nu K_{t+1}\Big(\lambda\Big(1-\frac{\widetilde{\lambda}}{4}\Big)\|\widetilde{{\mathbf{x}}}_{t}\|^{2}+\frac{5\alpha_{t}^{2}\lambda^{2}}{1-\lambda}\|\nabla\mathbf{f}^{\star}\|^{2}+\frac{5\alpha_{t}^{2}\lambda^{2}L^{2}}{1-\lambda}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}+\alpha_{t}^{2}n\sigma^{2}\lambda^{2}\Big)\bigg).$
Taking the full expectation and introducing the shorthand $S_{t}\coloneqq\alpha_{t}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$, we get
$\mathbb{E}[\exp(\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2})]\leq\exp(S_{t}\nu K_{t+1})\mathbb{E}\bigg[\exp\bigg(\lambda\nu K_{t+1}\Big(1-\frac{\widetilde{\lambda}}{4}\Big)\|\widetilde{{\mathbf{x}}}_{t}\|^{2}+\frac{5\alpha_{t}^{2}\lambda^{2}L^{2}\nu K_{t+1}}{1-\lambda}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}\bigg)\bigg]$
$\leq\exp(S_{t}\nu K_{t+1})\bigg(\mathbb{E}\big[\exp(\lambda\nu K_{t+1}\|\widetilde{{\mathbf{x}}}_{t}\|^{2})\big]\bigg)^{1-\nicefrac{{\widetilde{\lambda}}}{{4}}}\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{20\alpha_{t}^{2}\lambda^{2}L^{2}\nu K_{t+1}}{(1-\lambda)^{2}}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}\bigg)\bigg]\Bigg)^{\nicefrac{{\widetilde{\lambda}}}{{4}}}$
$\leq\exp(S_{t}\nu K_{t+1})\bigg(\mathbb{E}\big[\exp(\nu K_{t}\|\widetilde{{\mathbf{x}}}_{t}\|^{2})\big]\bigg)^{\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})\frac{t+t_{0}+2}{t+t_{0}+1}}\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{20\alpha_{t}^{2}\lambda^{2}L^{2}\nu K_{t+1}}{(1-\lambda)^{2}}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}\bigg)\bigg]\Bigg)^{\nicefrac{{\widetilde{\lambda}}}{{4}}},$ (34)
where the second inequality follows by applying Proposition 4 with $p=(1-\nicefrac{{\widetilde{\lambda}}}{{4}})^{-1}$ and $q=\frac{4}{\widetilde{\lambda}}$, while the third follows from the fact that $\lambda\frac{t+t_{0}+2}{t+t_{0}+1}\leq 1$ for $t_{0}\geq\frac{1}{1-\lambda}$ and applying Proposition 1. If $\alpha_{t}\leq\frac{1-\lambda}{2\lambda L\sqrt{10}}$, we get
$\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{20\alpha_{t}^{2}\lambda^{2}L^{2}\nu K_{t+1}}{(1-\lambda)^{2}}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2}\bigg)\bigg]\Bigg)^{\nicefrac{{\widetilde{\lambda}}}{{4}}}$ $\leq\bigg(\mathbb{E}\big[\exp(\nu K_{t}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2})\big]\bigg)^{\frac{5\alpha_{t}^{2}(t+t_{0}+2)\lambda^{2}L^{2}}{(1-\lambda)(t+t_{0}+1)}}$
$\leq\exp(\nu K_{t+1}D_{t}),$ (35)
where we used Proposition 1 in the first and Lemma 7 in the second inequality, with $D_{t}=\frac{5\alpha_{t}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{t}}{a\mu-1}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{a\mu}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(t+t_{0})^{a\mu}}\Big)$. Similarly, we use (F), to get
$\bigg(\mathbb{E}\big[\exp(\nu K_{t}\|\widetilde{{\mathbf{x}}}_{t}\|^{2})\big]\bigg)^{\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})\frac{t+t_{0}+2}{t+t_{0}+1}}$ $\leq\exp(\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})S_{t-1}\nu K_{t+1})\bigg(\mathbb{E}\big[\exp(\nu K_{t-1}\|\widetilde{{\mathbf{x}}}_{t-1}\|^{2})\big]\bigg)^{\lambda^{2}(1-\nicefrac{{\widetilde{\lambda}}}{{4}})^{2}\frac{t+t_{0}+2}{t+t_{0}}}$ (36)
$\times\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{20\alpha_{t}^{2}\lambda^{2}L^{2}\nu K_{t}}{(1-\lambda)^{2}}\|{\mathbf{x}}_{t-1}-{\mathbf{x}^{\star}}\|^{2}\bigg)\bigg]\Bigg)^{\nicefrac{{\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})\widetilde{\lambda}}}{{4}}}.$
Plugging (F) and (36) into (F), we get
$\mathbb{E}[\exp(\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2})]$ $\leq\exp\Big(\nu K_{t+1}\big(S_{t}+S_{t-1}\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})\big)+D_{t}+\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})D_{t-1}\big)\Big)$
$\times\Big(\mathbb{E}\big[\exp(\nu K_{t-1}\|\widetilde{{\mathbf{x}}}_{t-1}\|^{2})\big]\Big)^{\lambda^{2}(1-\nicefrac{{\widetilde{\lambda}}}{{4}})^{2}\frac{t+t_{0}+2}{t+t_{0}}}.$
Unrolling the recursion, it follows that
$\mathbb{E}[\exp(\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2})]$ $\leq\exp\bigg(\nu K_{t+1}\Big(\sum_{k=1}^{t}\lambda^{t-k}S_{k}+\sum_{k=1}^{t}\lambda^{t-k}D_{k}+\lambda^{t}\|\widetilde{{\mathbf{x}}}_{1}\|^{2}\Big)\bigg)$
$=\exp\bigg(\nu K_{t+1}\Big(\sum_{k=1}^{t}\lambda^{t-k}S_{k}+\sum_{k=1}^{t}\lambda^{t-k}D_{k}\Big)\bigg),$
where the last equality follows from the fact that $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. ∎
We are now ready to prove Theorem 2. Prior to that, we restate it, for convenience.
Theorem 2.
Let (A1)-(A4) and (A6) hold, let the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+d\sigma^{2}\kappa/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+\kappa^{2})}}{{(1-\lambda)}})}{(1-\lambda)n(T+t_{0})^{2}}$
$+\frac{\Delta_{f}}{n(T+t_{0})^{3}}+\frac{\sigma^{2}\lambda^{2}L^{3}(L\log(t+t_{0})+1)}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{\lambda^{2}L^{3}(1+L)\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(T+t_{0})^{3}}\bigg).$
Proof.
Starting from Lemma 1 and using property 3 of Proposition 8 with $x=\overline{x}^{t}$, we have
$f(\overline{x}^{t+1})\leq f(\overline{x}^{t})-\alpha_{t}\mu\big(f(\overline{x}^{t})-f^{\star}\big)-\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}_{t}\rangle+\alpha_{t}^{2}L\|\overline{z}_{t}\|^{2}+\frac{\alpha_{t}L^{2}}{2n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$
Subtracting $f^{\star}$ from both sides of the equation and defining $F_{t}=n(t+t_{0})(f(\overline{x}_{t})-f^{\star})$, for some $t_{0}>0$, it then follows that
$F_{t+1}$ $\leq(1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}F_{t}-\alpha_{t}n(t+t_{0}+1)\langle\nabla f(\overline{x}^{t}),\overline{z}_{t}\rangle$
$+\alpha_{t}^{2}n(t+t_{0}+1)L\|\overline{z}_{t}\|^{2}+\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}.$
Next, consider the MGF of $F_{t+1}$ conditioned on $\mathcal{F}_{t}$. Let $\nu\in(0,1]$ be a positive constant, we then have
$\mathbb{E}_{t}\big[\exp(\nu F_{t+1})\big]$ $\stackrel{{\scriptstyle(a)}}{{\leq}}\exp\bigg((1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}\nu F_{t}+\frac{\alpha_{t}\nu(t+t_{0}+1)L^{2}}{2}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\bigg)$
$\times\mathbb{E}_{t}\Big[\exp(-\alpha_{t}\nu n(t+t_{0}+1)\langle\nabla f(\overline{x}^{t}),\overline{z}_{t}\rangle+\alpha_{t}^{2}\nu n(t+t_{0}+1)L\|\overline{z}_{t}\|^{2})\Big]$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\exp\bigg((1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}\nu F_{t}+\frac{\alpha_{t}\nu(t+t_{0}+1)L^{2}}{2}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\bigg)$
$\times\sqrt{\mathbb{E}_{t}\Big[\exp(-2\alpha_{t}\nu n(t+t_{0}+1)\langle\nabla f(\overline{x}^{t}),\overline{z}_{t}\rangle)\Big]\mathbb{E}_{t}\Big[\exp(2\alpha_{t}^{2}\nu n(t+t_{0}+1)L\|\overline{z}_{t}\|^{2})\Big]}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\exp\bigg((1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}\nu F_{t}+\frac{\alpha_{t}\nu(t+t_{0}+1)L^{2}}{2}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\bigg)$
$\times\exp\bigg(\frac{3\alpha_{t}^{2}\nu^{2}n(t+t_{0}+1)^{2}\sigma^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{2}+120\alpha_{t}^{2}\nu\sigma^{2}d(t+t_{0}+1)L\bigg)$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\exp\bigg(\nu\Big(b_{t}F_{t}+c_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}+d_{t}\Big)\bigg),$
where in $(a)$ we used the fact that $F_{t}$ and $\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}$ are $\mathcal{F}_{t}$-measurable, $(b)$ follows from Proposition 4, in $(c)$ we use Lemma 2, Proposition 1 and impose the condition $\alpha_{t}\leq\frac{1}{4\sigma\sqrt{15(t+t_{0}+1)dL}}$, while $(d)$ follows from Proposition 7 and the definition of $F_{t}$, with $b_{t}=\Big(1-\alpha_{t}\mu+3\alpha_{t}^{2}\nu(t+t_{0}+1)L\Big)\frac{t+t_{0}+1}{t+t_{0}}$, $c_{t}=\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}$ and $d_{t}=120\alpha_{t}^{2}\sigma^{2}d(t+t_{0}+1)L$. Taking the full expectation and applying Proposition 4, we get
$\mathbb{E}\big[\exp(\nu F_{t+1})\big]$ $\leq\exp\big(\nu d_{t}\big)\mathbb{E}\Big[\exp\Big(\nu b_{t}F_{t}+\nu c_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\Big)\Big]$
$\leq\exp\big(\nu d_{t}\big)\sqrt[p]{\mathbb{E}\big[\exp\big(\nu pb_{t}F_{t}\big)\big]}\sqrt[q]{\mathbb{E}\big[\exp\big(\nu qc_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]}$ (37)
for some $p,q\in[1,\infty]$. We next analyze the expression $pb_{t}$. Recalling the definition of $b_{t}$, we get
$pb_{t}=p\bigg(1-\frac{a\mu}{t+t_{0}}+\frac{3a^{2}\nu(t+t_{0}+1)L}{(t+t_{0})^{2}}\bigg)\frac{t+t_{0}+1}{t+t_{0}}\leq p\bigg(1-\frac{a(\mu-6a\nu L)}{t+t_{0}}\bigg)\frac{t+t_{0}+1}{t+t_{0}}.$
Choosing $\nu\leq\frac{\mu}{12aL}$ and $p=1+\frac{\alpha_{t}\mu}{4}$, it follows that
$pb_{t}\leq p\bigg(1-\frac{a\mu}{2(t+t_{0})}\bigg)\frac{t+t_{0}+1}{t+t_{0}}\leq\bigg(1-\frac{a\mu}{4(t+t_{0})}\bigg)\bigg(1+\frac{1}{t+t_{0}}\bigg)\leq 1,$ (38)
where the last inequality follows since $a\mu>4$. Next, note that the choice of $p=1+\frac{\alpha_{t}\mu}{4}$ implies that $q=1+\frac{4}{\alpha_{t}\mu}$. From the definition of $c_{t}$, we then have
$qc_{t}=\bigg(1+\frac{4}{\alpha_{t}\mu}\bigg)\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}=\bigg(\frac{\alpha_{t}}{2}+\frac{2}{\mu}\bigg)(t+t_{0}+1)L^{2}\leq\frac{3L^{2}}{\mu}(t+t_{0}+1),$ (39)
where the first inequality follows from $\alpha_{t}\leq\frac{1}{\mu}$. Using (38) and (39) in (F), it follows that
$\mathbb{E}\big[\exp(\nu F_{t+1})\big]$ $\leq\exp\big(d_{t}\nu\big)\sqrt[p]{\big(\mathbb{E}\big[\exp(\nu F_{t})\big]\big)^{pb_{t}}}\sqrt[q]{\mathbb{E}\big[\exp\big(\nu(t+t_{0}+1)3\kappa L\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]}$
$=\exp\big(d_{t}\nu)\big(\mathbb{E}\big[\exp(\nu F_{t})\big]\big)^{b_{t}}\sqrt[q]{\mathbb{E}\big[\exp\big(\nu(t+t_{0}+1)3\kappa L\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]}.$ (40)
Using Lemma 4 with $K=3\kappa L$, we get
$\mathbb{E}\big[\exp\big(\nu qc_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]\leq\mathbb{E}\big[\exp\big(\nu K_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]\leq\exp\bigg(\nu K_{t}\Big(\sum_{k=1}^{t-1}\lambda^{t-1-k}S_{k}+\sum_{k=1}^{t-1}\lambda^{t-1-k}D_{k}\Big)\bigg),$
where we recall that $S_{k}=\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}=\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{a\mu-1}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{a\mu}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{a\mu}}\Big)$. To further bound the above expression, we use Lemma 6, to get
$\sum_{k=1}^{t-1}\lambda^{t-1-k}(S_{k}$ $+D_{k})\leq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}$
$+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}.$
Noting that $\frac{1}{q}=\frac{\alpha_{t}\mu}{4+\alpha_{t}\mu}\leq\frac{\alpha_{t}\mu}{4}$, we finally get
$\sqrt[q]{\mathbb{E}\big[\exp\big(\nu qc_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]}\leq\exp\bigg(\frac{\nu K_{t}\alpha_{t}\mu}{4}N_{t}\bigg)\leq\exp\bigg(\frac{3aL^{2}\nu N_{t}}{2}\bigg),$
where $N_{t}\coloneqq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}$. Define $G_{1}\coloneqq 240a^{2}\sigma^{2}dL$, $G_{2}\coloneqq\frac{6a^{3}\lambda^{2}L^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)}$, $G_{3}\coloneqq\frac{48a^{5}n\sigma^{2}\lambda^{2}L^{4}}{(1-\lambda)^{2}}$ and $G_{4}\coloneqq\frac{30a^{3}\lambda^{2}L^{4}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and plug into (F), to get
$\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]\leq\Big(\mathbb{E}\big[\exp\big(\nu F_{t}\big)\big]\Big)^{b_{t}}\exp\bigg(\sum_{i\in[3]}\frac{\nu G_{i}}{(t+t_{0})^{i}}+\frac{\nu G_{4}}{(t+t_{0})^{8}}\bigg).$
Recalling the definition of $b_{t}$ and (38), it follows that $b_{t}\leq 1-\frac{a\mu/2-1}{t+t_{0}}=1-\frac{2}{t+t_{0}}$, therefore we can bound the MGF of $\nu F_{t+1}$ using Lemma 5 with $a=2$, $M=8$, $C_{i}=G_{i}$ for $i\in[3]$, $C_{8}=G_{4}$ and $C_{j}=0$, for $j\in\{4,\ldots,7\}$, to finally get
$\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]$ $\leq\exp\bigg(\frac{(t_{0}+2)^{3}\nu\Delta_{f}}{(t+1+t_{0})^{2}}+4\nu G_{1}+\frac{4\nu G_{2}}{t+1+t_{0}}\bigg)$
$\times\exp\bigg(\frac{4\nu G_{3}\log(t+t_{0}+1)}{(t+t_{0}+1)^{2}}+\frac{4\nu G_{4}}{5(t_{0}+1)^{5}(t+1+t_{0})^{2}}\bigg).$ (41)
Applying Markov’s inequality, we then get, for any $\epsilon>0$
$\mathbb{P}\left(f(\overline{x}^{t+1})-f^{\star}>\epsilon\right)=\mathbb{P}\big(\nu F_{t+1}>\nu n(t+1+t_{0})\epsilon\big)\leq\exp\left(-\nu n(t+1+t_{0})\epsilon\right)\mathbb{E}\big[\big(\nu F_{t+1}\big)\big].$
Using (F), it can be readily verified that choosing
$\epsilon_{t}^{1}$ $=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})+4G_{1}}{n(t+t_{0})}+\frac{4G_{2}}{n(t+t_{0})^{2}}+\frac{(t_{0}+2)^{3}\Delta_{f}+4G_{3}\log(t+t_{0})+4G_{4}/5(t_{0}+1)^{5}}{n(t+t_{0})^{3}},$ (42)
for any $\delta\in(0,1)$, results in $\mathbb{P}\left(f(\overline{x}^{t})-f^{\star}>\epsilon_{t}^{1}\right)\leq\delta$. Next, using Proposition 7 with $x=x^{t}_{i}$ and $y=\overline{x}^{t}$, we get
$f(x^{t}_{i})$ $\leq f(\overline{x}^{t})+\langle\nabla f(\overline{x}^{t}),x^{t}_{i}-\overline{x}^{t}\rangle+\frac{L}{2}\|x^{t}_{i}-\overline{x}^{t}\|^{2}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}f(\overline{x}^{t})+\frac{1}{2L}\|\nabla f(\overline{x}^{t})\|^{2}+\frac{L}{2}\|x^{t}_{i}-\overline{x}^{t}\|^{2}+\frac{L}{2}\|x^{t}_{i}-\overline{x}^{t}\|^{2}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}f(\overline{x}^{t})+f(\overline{x}^{t})-f^{\star}+L\|x^{t}_{i}-\overline{x}^{t}\|^{2},$
where in $(i)$ we used Proposition 3 with $\epsilon=L$, while $(ii)$ follows from Proposition 7. Subtracting $f^{\star}$ from both sides and averaging over all users $i\in[n]$, we get
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)\leq 2\big(f(\overline{x}^{t})-f^{\star}\big)+\frac{L}{n}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}.$ (43)
We now consider two events, $A_{t,\epsilon}\coloneqq\left\{\omega:f(\overline{x}^{t})-f^{\star}>\epsilon\right\}$ and $B_{t,\epsilon}\coloneqq\left\{\omega:\frac{L}{n}\|{\mathbf{x}^{t}}-\overline{x}^{t}\|^{2}>\epsilon\right\}$. From the previous analysis, we know that, for any $\delta\in(0,1)$ and $\epsilon_{t}^{1}$ from (42), we have $\mathbb{P}\big(A_{t,\epsilon_{t}^{1}}\big)\leq\delta$. Similarly, using Markov’s inequality and Lemma 4 with $K=L$, we have, for any $\epsilon>0$
$\mathbb{P}\bigg(\frac{L}{n}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}>\epsilon\bigg)$ $=\mathbb{P}\left(\nu K_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}>\nu n(t+t_{0}+1)\epsilon\right)$
$\leq\exp(-\epsilon\nu n(t+t_{0}+1))\mathbb{E}\big[\exp\big(\nu K_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]$
$\leq\exp\Bigg(-\epsilon\nu n(t+t_{0}+1)+K_{t}\nu\bigg(\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}\bigg)\Bigg)$
$\times\exp\Bigg(K_{t}\nu\bigg(\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}\bigg)\Bigg).$
Therefore, it can be readily seen that, choosing
$\epsilon_{t}^{2}=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})}{n(t+t_{0}+1)}$ $+\frac{4a^{2}\lambda^{2}L(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{32a^{4}\sigma^{2}\lambda^{2}L^{3}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{3}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}},$ (44)
we get $\mathbb{P}\big(B_{t,\epsilon_{t}^{2}}\big)\leq\delta$, for any $\delta\in(0,1)$. Finally, let $C_{t}\coloneqq\left\{\omega:\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)>2\epsilon_{t}^{1}+\epsilon_{t}^{2}\right\}$. From (43) it readily follows that, for any $\delta\in(0,\nicefrac{{1}}{{2}})$, we have
$\mathbb{P}(C_{t})\leq\mathbb{P}\big(A_{t,\epsilon_{t}^{1}}\cap B_{t,\epsilon_{t}^{2}}\big)\leq\mathbb{P}\big(A_{t,\epsilon_{t}^{1}}\big)+\mathbb{P}\big(B_{t,\epsilon_{t}^{2}}\big)\leq 2\delta.$
Therefore, we get, for any $\delta\in(0,1)$, with probability at least $1-\delta$
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)$ $=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\sigma^{2}\lambda^{2}L^{3}(L\log(t+t_{0})+1)}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{\lambda^{2}L^{3}(1+L)\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}\bigg).$
Finally, it can be verified that the conditions on $a$, $t_{0}$ and $\nu$ in the statement of the theorem ensure that all the step-size conditions are satisfied, completing the proof. ∎
We remark that, similarly to the proof of Theorem 1, one can analyze the strongly convex case with a fixed step-size, resulting in the dependence on some terms, e.g., optimality and iterate gaps $\Delta_{f}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}}^{\star}\|$, decaying exponentially fast, i.e., $\mathcal{O}\big((\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}+\Delta_{f})e^{-CT}\big)$, for some $C>0$, as shown in, e.g., [11] for MSE guarantees in decentralized, or [40] for HP guarantees in centralized settings. For simplicity, we omit this analysis.
Appendix G On Dimension Dependence
As mentioned in Section 4 in the main body, our rates in Theorems 1 and 2 exhibit a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, with the dependence stemming from Lemma 2, where it is shown that $\overline{z}^{t}$ is $\mathcal{O}\big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\big)$-sub-Gaussian. This is a consequence of working with random vectors, where we simultaneoulsy need to show a bound on the MGF of the inner product $\langle\overline{z}^{t},v\rangle$ for any $\mathcal{F}_{t}$-measurable vector $v\in\mathbb{R}^{d}$ (shown in point 2 of Lemma 2), as well as on the MGF of the norm-squared $\|\overline{z}^{t}\|^{2}$. While the inner product maintains some desirable properties, such as being zero-mean and linear (in the sense that $\langle\overline{z}^{t},v\rangle=\frac{1}{n}\sum_{i\in[n]}\langle z^{t}_{i},v\rangle$), this is not the case with the squared norm, which is neither zero-mean, nor linear. Therefore, trying to directly establish the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\big(\frac{\sigma}{\sqrt{n}}\big)$-sub-Gaussian, in the sense of condition (A4), fails to yield the desired result. In particular, recalling that $\sigma^{2}=\frac{1}{n}\sum_{i\in[n]}\sigma_{i}^{2}$, we then have
$\mathbb{E}\bigg[\exp\bigg(\frac{\|\overline{z}^{t}\|^{2}}{\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]$ $\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{E}\Bigg[\exp\Bigg(\frac{\big(\sum_{i\in[n]}\|z^{t}_{i}\|\big)^{2}}{n\sum_{i\in[n]}\sigma_{i}^{2}}\Bigg)\>\Big|\>\mathcal{F}_{t}\Bigg]\stackrel{{\scriptstyle(ii)}}{{\leq}}\mathbb{E}\bigg[\exp\bigg(\frac{1}{n}\sum_{i\in[n]}\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}\prod_{i\in[n]}\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]\Bigg)^{1/n}\stackrel{{\scriptstyle(iv)}}{{\leq}}\prod_{i\in[n]}\exp\bigg(\frac{1}{n}\bigg)=\exp(1),$
where $(i)$ follows from Proposition 1, $(ii)$ follows from Sedrakyan’s inequality, namely that
$\frac{(\sum_{i\in[n]}a_{i})^{2}}{\sum_{i\in[n]}b_{i}}\leq\sum_{i\in[n]}\frac{a_{i}^{2}}{b_{i}},$
which holds for any $n\in\mathbb{N}$, $a_{i}\in\mathbb{R}$ and $b_{i}>0$, in $(iii)$ we used Proposition 1 and the fact that noise is conditionally independent across users, while $(iv)$ follows from (A4). Therefore, a direct approach yields that $\overline{z}^{t}$ is $\sigma$-sub-Gaussian, failing to show the variance reduction benefit. To circumvent this issue, we use a different argument, namely that a zero-mean random vector $x\in\mathbb{R}^{d}$, which is $\sigma$-sub-Gaussian in the inner product sense, i.e., for any $v\in\mathbb{R}^{d}$
$\mathbb{E}\big[\exp\big(\langle x,v\rangle\big)\big]\leq\exp\bigg(\frac{\sigma^{2}\|v\|^{2}}{2}\bigg),$
is also $\sigma\sqrt{d}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$
$\mathbb{P}\big(\|x\|>\epsilon\big)\leq 2\exp\bigg(-\frac{\epsilon^{2}}{2\sigma^{2}d}\bigg),$
see [75, Lemma 1]. This argument allows us to show the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian (in the sense of (A4)), at the cost of introducing a $\sqrt{d}$ dependence. One way to mitigate this is to directly assume that the noise $z^{t}_{i}$ at each user is zero-mean and $\sigma_{i}$-norm-sub-Gaussian. Using a similar argument to the one in Lemma 6 and Corollary 7 in [75], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$. This, however, comes at the cost of imposing a slightly stronger noise condition, as any $\sigma$-norm-sub-Gaussian random vector is also $\sigma$-sub-Gaussian (in the inner product sense), while the opposite implication inevitably introduces a $\sqrt{d}$ factor, as is the case in Lemma 2. Finally, we note that, in the regime $n\ll d$, we can simply use the above argument following Sedrakyan’s inequality, to conclude that $\overline{z}^{t}$ is $\sigma$-sub Gaussian, reducing the dependence on problem dimension $d$, at the cost of losing linear speed-up in the number of users.

### v2

@@
 1 Introduction
-Modern large-scale machine learning applications and the abundance of data necessitate alternatives to the centralized computation framework, giving rise to distributed learning, a paradigm where multiple users collaborate to jointly train a model, e.g., [1, 2, 3].
-The features of distributed learning, such as storing the data locally and only exchanging smaller updates, like (quantized) model parameters or gradients, and the lack of a single point of failure, further make it an attractive paradigm from a privacy and security perspective [4, 5].
-Many applications, such as federated training of models on mobile devices [6], controlling and coordinating robot swarms [7], or distributed control and power grids [8], all rely on distributed computation.
+Modern large-scale machine learning applications and the abundance of data necessitate alternatives to the centralized computation framework, giving rise to distributed learning, a paradigm where multiple users collaborate to jointly train a model, e.g., [sayed-networks, pmlr-v54-mcmahan17a, vlaski_et_al].
+The features of distributed learning, such as storing the data locally and only exchanging smaller updates, like (quantized) model parameters or gradients, and the lack of a single point of failure, further make it an attractive paradigm from a privacy and security perspective [bonawitz2017practical, yu-secure].
+Many applications, such as federated training of models on mobile devices [konecny2016federated], controlling and coordinating robot swarms [bullo2009distributed], or distributed control and power grids [dist-learn-noncvx], all rely on distributed computation.
 From a communication/model exchange perspective, distributed frameworks can be client-server (i.e., federated) or decentralized (i.e., networked), with the main difference being that in the client-server setup users communicate with a server, while in decentralized settings users communicate directly with each other.
-[^1] While the client-server setup is often characterized by performing local updates and periodic communication [9], as well as partial user participation [10], we note that these features can also be incorporated in the decentralized setup, by modifying the methods to perform multiple local updates before communicating [11] and considering a dynamic network with users that are occasionally idle [12].
+[^1] While the client-server setup is often characterized by performing local updates and periodic communication [stichlocal], as well as partial user participation [pmlr-v151-jee-cho22a], we note that these features can also be incorporated in the decentralized setup, by modifying the methods to perform multiple local updates before communicating [pmlr-v119-koloskova20a] and considering a dynamic network with users that are occasionally idle [Cannelli2020].
 However, this is not the focus of our current work.
 Noting that, from the model update perspective, the client-server setup is equivalent to the decentralized setup with a fully connected communication network, we focus on the more general, decentralized setup.
-The study of convergence guarantees of decentralized optimization algorithms has a long history, e.g., [13, 14, 15, 16, 17, 18, 19], with most works focusing on MSE convergence, e.g., [20, 21, 22, 23, 11, 24, 25, 26], see also [27] for an extensive treatment of the topic.
+The study of convergence guarantees of decentralized optimization algorithms has a long history, e.g., [nedic-subgrad, jakovetic-fast, shi-extra, scutari-next, kun-dgd, jakovetic-unification, swenson-dsgd], with most works focusing on MSE convergence, e.g., [conv-rates-dsgd-jakovetic, Pu2021, vlaski-dsgd-nonconv-1, vlaski-dsgd-nonconv-2, pmlr-v119-koloskova20a, ran-vr, ran-improved, wang-cooperative], see also [adaptive-learning-ali] for an extensive treatment of the topic.
 Another type of convergence guarantees, namely convergence in HP, has garnered increasing attention recently.
@@
 This is particularly important in modern applications like LLMs, where it is often intractable to perform multiple runs.
-Numerous works study HP guarantees of SGD-type methods in centralized settings, under both light-tailed [28, 29, 30, 31, 32, 33] and heavy-tailed noise [34, 35, 36, 37, 38].
+Numerous works study HP guarantees of SGD-type methods in centralized settings, under both light-tailed [nemirovski2009robust, ghadimi2013stochastic, li-orabona, pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high] and heavy-tailed noise [nguyen2023improved, liu2023breaking, hubler2024normalization, kornilov2025sign, armacki2023high].
 Comparatively, there have been few studies on HP convergence of decentralized methods.
@@
 Centralized HP Convergence.
-The authors in [28] show optimal convergence rates of SGD for convex costs under light-tailed (i.e., sub-Gaussian) stochastic gradients (see assumption (A4) ahead for a formal definition of sub-Gaussianity), while the work [29] shows the same for non-convex costs.
-The authors in [30, 39] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively.
-The works [31] and [32] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while in [40] the authors establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs.
-The work [33] generalizes the previous studies for non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs.
-The authors in [41] study HP convergence of SGD under sub-Weibull noise.
-Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [42, 34, 38] for various conditions used when studying heavy-tailed noise.
-e.g., [43, 34, 35, 44, 36, 37, 45, 46, 38], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1).
+The authors in [nemirovski2009robust] show optimal convergence rates of SGD for convex costs under light-tailed (i.e., sub-Gaussian) stochastic gradients (see assumption (A4) ahead for a formal definition of sub-Gaussianity), while the work [ghadimi2013stochastic] shows the same for non-convex costs.
+The authors in [li-orabona, li2020high] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively.
+The works [pmlr-v99-harvey19a] and [pmlr-v206-bajovic23a] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while in [liu2024revisiting] the authors establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs.
+The work [liu2023high] generalizes the previous studies for non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs.
+The authors in [madden2020high] study HP convergence of SGD under sub-Weibull noise.
+Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [heavy-tail-book, nguyen2023improved, armacki2023high] for various conditions used when studying heavy-tailed noise.
+e.g., [sadiev2023highprobability, nguyen2023improved, liu2023breaking, gorbunov2023breaking, hubler2024normalization, kornilov2025sign, armacki2024_ldp+mse, armacki2025optimal, armacki2023high], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1).
 Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
@@
 MSE guarantees are typically studied under a variety of bounds on the second noise moment.
-[^3] See, e.g., [47] for a good overview of the various conditions used in the literature.
-Following one of these settings, the work [20] shows that DSGD converges at an optimal rate for strongly convex costs, while in [21] the authors show that DSGD with a fixed step-size and the gradient tracking (GT) mechanism, e.g., [16, 48, 49], converges to a neighbourhood of the optimal solution.
-The works [22, 23, 50] study MSE guarantees of DSGD for non-convex costs and show that it escapes saddle points with high probability.
-In [26] the authors propose a general framework dubbed cooperative SGD, showing optimal rates for non-convex costs.
-The work [51] provides unified guarantees for DSGD with local updates and changing network topology, with optimal rates and linear speed-up in the number of users for non-convex and (strongly) convex costs.
-In [24] the authors show DSGD with GT and variance reduction converges at a linear rate for strongly convex costs, while in [25] optimal rates of DSGD with GT and linear speed-up in the number of users are established, for non-convex costs and costs satisfying the Polyak-Łojasiewicz (PL) condition, e.g., [52].
-Finally, in [53] the authors show that DSGD with GT, local momentum and gradient normalization achieves optimal rates under heavy-tailed noise.
-It is also worth mentioning a rich line of works studying MSE guarantees in decentralized settings for problems such as estimation and detection, multi-objective and multitask optimization, see [54, 55, 56] and references therein.
+[^3] See, e.g., [khaled2022better] for a good overview of the various conditions used in the literature.
+Following one of these settings, the work [conv-rates-dsgd-jakovetic] shows that DSGD converges at an optimal rate for strongly convex costs, while in [Pu2021] the authors show that DSGD with a fixed step-size and the gradient tracking (GT) mechanism, e.g., [scutari-next, nedic-tracking, qu-harnessing], converges to a neighbourhood of the optimal solution.
+The works [vlaski-dsgd-nonconv-1, vlaski-dsgd-nonconv-2, vlaski-dsgd-nonconv-3] study MSE guarantees of DSGD for non-convex costs and show that it escapes saddle points with high probability.
+In [wang-cooperative] the authors propose a general framework dubbed cooperative SGD, showing optimal rates for non-convex costs.
+The work [koloskova2023grad_clip] provides unified guarantees for DSGD with local updates and changing network topology, with optimal rates and linear speed-up in the number of users for non-convex and (strongly) convex costs.
+In [ran-vr] the authors show DSGD with GT and variance reduction converges at a linear rate for strongly convex costs, while in [ran-improved] optimal rates of DSGD with GT and linear speed-up in the number of users are established for non-convex costs and costs satisfying the Polyak-Łojasiewicz (PL) condition [karimi2016linear].
+Finally, in [yu2025decentralized] the authors show that DSGD with GT, local momentum and gradient normalization achieves optimal rates under heavy-tailed noise.
+It is also worth mentioning a rich line of works studying MSE guarantees in decentralized settings for problems such as estimation and detection, multi-objective and multitask optimization, see [consensus+innovation1, chen-decent-pareto, nassif-multitask] and references therein.
 Decentralized HP Convergence.
 Compared to MSE guarantees, there is a significantly smaller body of work on HP convergence guarantees of decentralized algorithms.
-In particular, the work [57] studies HP convergence of DSGD for both non-convex and PL costs under light-tailed noise, requiring uniformly bounded gradients and asymptotically vanishing noise for PL costs (see discussion after Theorem 2 ahead), showing optimal rates in both cases.
-In [58, 59] the authors study HP convergence of a decentralized mirror descent algorithm under sub-Gaussian noise and bounded gradients, for online noncooperative games and dynamic regrets, respectively.
-Works [60, 61, 62] study HP convergence of decentralized algorithms with clipping under heavy-tailed noise.
-It is worth mentioning [63, 64, 65, 66, 67], who study asymptotic large deviation guarantees for decentralized problems like detection and inference.
+In particular, the work [dsgd-high-prob] studies HP convergence of DSGD for both non-convex and PŁ costs under light-tailed noise, requiring uniformly bounded gradients and asymptotically vanishing noise for PŁ costs (see discussion after Theorem 2 ahead), showing optimal rates in both cases.
+In [dsgd-online-noncooperative-games, dsgd-online-stochastic] the authors study HP convergence of a decentralized mirror descent algorithm under sub-Gaussian noise and bounded gradients, for online noncooperative games and dynamic regrets, respectively.
+Works [clipped-dual-avg, clipped-dsgd1, clipped-dsgd2] study HP convergence of decentralized algorithms with clipping under heavy-tailed noise.
+It is also worth mentioning [bajovic-detection-ld, bajovic-detection-ld2, matta-diffusion-1, matta-difussion-2, bajovic-inference-ld], where the authors study asymptotic large deviation guarantees for decentralized problems like detection and inference.
 A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradient stays bounded.
-This observation raises the question of whether decentralized algorithms can converge in HP without ensuring that the gradient is uniformly bounded, creating a gap between convergence in the HP and the MSE sense, where bounded gradients are not required.
+This observation raises the question of whether decentralized algorithms can converge in the HP sense without ensuring that the gradient is uniformly bounded, creating a gap between convergence in the HP and MSE sense, where bounded gradients are not required.
 This is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence.
-Moreover, none of the existing HP convergence bounds in decentralized settings show linear speed-up in the number of users, established in the MSE sense in, e.g., [11, 25].
+Moreover, none of the existing HP convergence bounds in decentralized settings achieve linear speed-up in the number of users, which is a known feature of decentralized algorithms, established in the MSE sense in, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined].
 1.2 Contributions
@@
 Next, we show that DSGD achieves optimal rates and linear speed-up in the number of users, for both non-convex and strongly convex costs, further improving on existing works and closing the gap between HP and MSE guarantees in decentralized settings.
-Our results are established by carefully bounding the MGF of the quantity of interest (i.e., average norm-squared of the gradient for non-convex and optimality gap of the final iterate for strongly convex costs) and the MGF of the consensus gap.
-Compared to [57], the work closest to ours, we provide several improvements.
-In particular, we remove the uniformly bounded gradient requirement, as well as the vanishing noise requirement imposed for PL costs (see the discussion after Theorem 2 ahead), while achieving linear speed-up in the number of users.
-Compared to works studying MSE guarantees for DSGD, e.g., [11, 20], our results require directly working with the MGF and carefully balancing between the MGF of the quantity of interest and that of the consensus gap.
+Our results are established by carefully bounding the MGF of the quantity of interest (i.e., average norm-squared of the gradient for non-convex and optimality gap of the last iterate for strongly convex costs) and the MGF of the consensus gap.
+Compared to [dsgd-high-prob], the work closest to ours, we provide several improvements.
+In particular, we remove the uniformly bounded gradient requirement, as well as the vanishing noise requirement imposed for PŁ costs (see the discussion after Theorem 2 ahead), while achieving linear speed-up in the number of users.
+Compared to works studying MSE guarantees for DSGD, e.g., [pmlr-v119-koloskova20a, conv-rates-dsgd-jakovetic], our results require directly working with the MGF and carefully balancing between the MGF of the quantity of interest and that of the consensus gap.
 This is particularly challenging for strongly convex costs, where we want to show an “almost decreasing” property of the MGF, in order to get improved rates on the last iterate (see Lemma 5 and the related discussion ahead).
-Compared to centralized SGD, e.g., [31, 32, 33, 40], the main challenge lies in the additional consensus gap and simultaneously dealing with its MGF and that of the quantity of interest.
+Compared to centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], the main challenge lies in the additional consensus gap and simultaneously dealing with its MGF and that of the quantity of interest.
 All of these challenges are resolved by introducing several novelties, outlined next.
@@
 To establish linear speed-up in the number of users, we show that the variance reduction benefit of decentralized learning is maintained in the HP sense, in the form of Lemma 2, which is of independent interest when studying HP guarantees of decentralized algorithms.
-Further, to guarantee linear speed-up for strongly convex costs, we require a more fine-grained bound on the MGF, provided in the form of Lemma 5, which allows for incorporating higher-order terms, generalizing the bounds from [31, 32, 40] and is of independent interest, even in centralized settings.
+Further, to guarantee linear speed-up for strongly convex costs, we require a more fine-grained bound on the MGF, provided in the form of Lemma 5, which allows for incorporating higher-order terms, generalizing the bounds from [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting] and is of independent interest, even in centralized settings.
 Paper Organization.
@@
 We use subscripts to denote users and superscripts to denote the iteration counter, e.g., $x^{t}_{i}$ refers to the model of user $i$ in iteration $t$.
-We use the “big O” notation $\mathcal{O}(\cdot)$ to only hide global constants, unless stated otherwise.
+We use the “big O” notation $\mathcal{O}(\cdot)$ to hide global constants, unless stated otherwise.
 2 Problem Setup and Proposed Method
@@
   1. 1.
-    Batch (i.e., offline) learning, where users have access to a local dataset $\{\xi_{i,l}\}_{l\in[m_{i}]}$, so that $f_{i}(x)=\frac{1}{m_{i}}\sum_{l\in[m_{i}]}\ell(x;\xi_{i,l})$ and in each round users choose a sample $\xi_{i,l}$ uniformly at random, [^4] Note that it is possible to have non-uniform weights, i.e., $f_{i}(x)=\sum_{l\in[m_{i}]}p_{i,l}\ell(x;\xi_{i,l})$, where $p_{i,l}>0$ for all $l\in[m_{i}]$ and $\sum_{l\in[m_{i}]}p_{i,l}=1$, resulting in non-uniform sampling, with sample $\xi_{i,l}$ selected with probability $p_{i,l}$.
+    Batch (offline) learning, where users have access to a local dataset $\{\xi_{i,l}\}_{l\in[m_{i}]}$, so that $f_{i}(x)=\frac{1}{m_{i}}\sum_{l\in[m_{i}]}\ell(x;\xi_{i,l})$ and in each round users choose a sample $\xi_{i,l}$ uniformly at random, [^4] Note that it is possible to have non-uniform weights, i.e., $f_{i}(x)=\sum_{l\in[m_{i}]}p_{i,l}\ell(x;\xi_{i,l})$, where $p_{i,l}>0$ for all $l\in[m_{i}]$ and $\sum_{l\in[m_{i}]}p_{i,l}=1$, resulting in non-uniform sampling, with sample $\xi_{i,l}$ selected with probability $p_{i,l}$.
 which is used to compute $\nabla\ell(x;\xi_{i,l})$ and update model parameters;
@@
     Streaming (i.e., online) learning, where users do not store a local dataset, but in each round observe a random sample $\xi_{i}\sim\mathcal{D}_{i}$ from a data stream, which is used to compute $\nabla\ell(x;\xi_{i})$ and update model parameters.
-The communication pattern among users is modeled as a static graph $G=(V,E)$, where $V=[n]$ is the set of vertices, representing users, while $E\subset V\times V$ is the set of edges, representing communication links between users.
-To solve (2) in decentralized fashion, we consider a version of Decentralized Stochastic Gradient Descent (DSGD), based on the Adapt-Then-Combine (i.e., diffusion) approach, e.g., [68, 69, 70].
+The communication pattern between users is modeled as a static graph $G=(V,E)$, where $V=[n]$ is the set of vertices representing users, while $E\subset V\times V$ is the set of edges representing communication links between users.
+[^5] Our work can be readily extended to non-static, time-varying graphs, see the Appendix for a further discussion.
+To solve (2) in decentralized fashion, we consider a version of Decentralized Stochastic Gradient Descent (DSGD), based on the Adapt-Then-Combine (i.e., diffusion) approach, e.g., [diffusion1, diffusion2, diffusion3].
 The method consists of the following steps.
@@
 The method is summarized in Algorithm 1.
-The version of DSGD considered in our work is related to the version based on Combine-Then-Adapt (i.e., consensus ++ innovation) approach, e.g., [13, 54, 71], whose HP convergence is studied in [57].
+The version of DSGD considered in our work is related to the version based on Combine-Then-Adapt (i.e., consensus ++ innovation) approach, e.g., [nedic-subgrad, consensus+innovation1, consensus+innovation2], whose HP convergence is studied in [dsgd-high-prob].
 Algorithm 1 DSGD
@@
 (A1) The network communication matrix $W\in\mathbb{R}^{n\times n}$ is primitive and doubly stochastic.
-Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [24].
-Moreover, it can be shown that (A1) implies that the second largest singular value of $W$, denoted by $\lambda$, satisfies $\lambda\in[0,1)$, see [72].
+Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [ran-vr].
+Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix.
+It can be shown that (A1) implies $\lambda\in[0,1)$, see [Horn_Johnson_2012].
 (A2) The global cost $f$ is bounded from below, i.e., $f^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}f(x)>$-$\infty$.
 (A3) Each local cost $f_{i}$ has $L$-Lipschitz gradients, i.e., $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, for any $x,y\in\mathbb{R}^{d}$.
-Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [29, 11, 25].
+Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [ghadimi2013stochastic, pmlr-v119-koloskova20a, ran-improved].
 (A4) The stochastic quantities satisfy the following:
@@
     $\mathbb{E}\left[\exp\bigg(\frac{\|z_{i}^{t}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1).$
-The first condition in (A4) is standard in decentralized stochastic optimization [25, 53], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian).
-Light tails are necessary to achieve exponentially decaying tail bounds of the form (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [28, 29, 30, 39, 31, 33, 32].
-[^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [41] and references therein, the focus of this work is on the more widely studied sub-Gaussian noise setting.
+The first condition in (A4) is standard in decentralized stochastic optimization [ran-improved, yu2025decentralized], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian).
+Light tails are necessary to achieve exponentially decaying tail bounds of the form (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [nemirovski2009robust, ghadimi2013stochastic, li-orabona, li2020high, pmlr-v99-harvey19a, liu2023high, pmlr-v206-bajovic23a].
+[^7] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [madden2020high] and references therein, the focus of this work is on the more widely studied sub-Gaussian noise setting.
 (A5) The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, [^8] With at least one of $A$ or $B$ being strictly positive.
 we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
-A heterogeneity bound of the type (A5) is in fact required to ensure that DSGD converges for non-convex costs [73, 8, 11, 26].
-Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [57], as it allows users’ gradients to grow with the global gradient.
-Compared to [11], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition.
-While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [11], to keep the proofs simple and instructive, we will use assumption (A5).
+A heterogeneity bound of the type (A5) is in fact required to ensure that DSGD converges for non-convex costs [NIPS2017_f7552665, dist-learn-noncvx, pmlr-v119-koloskova20a, wang-cooperative].
+Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [dsgd-high-prob], as it allows users’ gradients to grow with the global gradient.
+Compared to [pmlr-v119-koloskova20a], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition.
+While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [pmlr-v119-koloskova20a], to keep the proofs simple and instructive, we will use assumption (A5).
 (A6) Each $f_{i}$ is twice continuously differentiable and $\mu$-strongly convex, i.e., for every $x,y\in\mathbb{R}^{d}$, we have $f_{i}(x)\geq f_{i}(y)+\langle\nabla f_{i}(y),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}$.
 Assumption (A6) is used in the analysis of strongly convex costs.
-It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [20, 11, 26] which is in general not required for centralized SGD.
-The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [25], however, this is beyond the scope of the current work.
+It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a, wang-cooperative] which is in general not required for centralized SGD.
+The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [ran-improved], however, this is beyond the scope of the current work.
 Next, let $\overline{x}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}x^{t}_{i}$ denote the network-wise average model.
@@
 Lemma 2 establishes some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, highlighting the variance reduction benefit of decentralized optimization in the HP sense.
-This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we establish in the following subsections.
+This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections.
 We note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and the Appendix.
@@
 Let (A1)-(A5) hold.
-If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$, where $C>0$ is a problem related constant satisfying
+If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
 $C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
 we then have, for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $=\mathcal{O}\Bigg(\frac{\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+L\big)}{\sqrt{nT}}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}$
-$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{n\lambda^{2}(A^{2}+\sigma^{2})]}{\sigma^{2}(1-\lambda^{2})T}\Bigg).$
-Theorem 1 establishes HP convergence guarantees for DSGD with general smooth non-convex costs and a fixed step-size.
-The bound in Theorem 1 consists of three terms, with the first two matching centralized SGD, while the third captures the effect of the network, through $\frac{\lambda^{2}}{1-\lambda^{2}}$.
-Next, note that the leading term $\mathcal{O}\big(\frac{\sigma\sqrt{d}}{\sqrt{nT}}\big)$ achieves a linear speed-up in the number of users, showing the benefits of decentralized learning in the HP sense.
-Finally, we can see that, in the case $A=0$, the above bound recovers the rates of noiseless centralized gradient descent (GD), i.e., when $\sigma=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$.
-Compared to [57], where the authors provide a $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$ HP convergence bound, while requiring uniformly bounded gradients, we provide several improvements, by relaxing the uniformly bounded gradients assumption, achieving a linear speed-up and recovering the rates of GD when $\sigma=0$ and the gradients of users grow at most as the global gradient.
+$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}$
+$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
+Theorem 1 establishes HP convergence guarantees for DSGD with general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few perspectives.
+Centralized SGD and linear speed-up.
+Compared to centralized SGD, which achieves the rate $\mathcal{O}\Big(\frac{1}{\sqrt{T}}+\frac{1}{T}\Big)$, e.g., [liu2023high], we can see that the rate from Theorem 1 similarly consists of two terms of different orders, namely $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}+\frac{n}{T}\Big)$.
+Crucially, the leading term achieves a linear speed-up in the number of users, which is consistent with MSE results, e.g., [pmlr-v119-koloskova20a], and shows that DSGD maintains the benefits of decentralized learning, in the HP sense.
+Network effect and transient time.
+We can see from the bound in Theorem 1 that network connectivity only affects the higher-order terms decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is again consistent with MSE results, e.g., [pmlr-v119-koloskova20a].
+Moreover, we can use the rate in Theorem 1 to establish the transient time of DSGD in the HP sense, which can be shown to be of the order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD stemming from MSE results, see, e.g., [unified-refined, Table I].
+For details on the derivation of the transient time, see the Appendix.
+Effects of noise and heterogeneity.
+The noise (through $\sigma,\overline{\sigma}$) and heterogeneity (through $A,B$) mainly affect the higher-order terms, [^9] The effect of noise on the leading term can be eliminated by choosing the step-size $\alpha=\min\Big\{C,\frac{\sqrt{n}}{\sqrt{dLT}}\Big\}$.
+which is again consistent with MSE results.
+Note that, if $A=0$, the bound in Theorem 1 recovers the convergence rate of noiseless gradient descent, in the sense that, when $\sigma=A=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$.
+Comparison with [dsgd-high-prob].
+The authors in [dsgd-high-prob] study HP guarantees of DSGD over undirected time-varying networks, establishing a rate $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$, using a time-varying step-size schedule $\alpha_{t}=\frac{C}{\sqrt{t+1}}$, for some $C>0$ and assuming uniformly bounded gradients.
+As we discuss next, our results provide several important improvements.
+First, we remove the uniformly bounded gradient assumption and replace it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details.
+Next, the authors in [dsgd-high-prob] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise.
+In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, as all terms converge at the same $\mathcal{O}\Big(\frac{1}{\sqrt{T}}\Big)$ rate.
+However, it would be impossible to obtain a linear speed-up in [dsgd-high-prob] even under a fixed step-size, as the authors show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity.
+On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used (see the Appendix for a detailed discussion on the results needed to achieve linear speed-up).
+[^10] This is also consistent with existing MSE guarantees, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined], where linear speed-up for non-convex costs is achieved using a fixed step-size.
+Finally, while we consider static networks, our results can be readily extended to the time-varying setting considered in [dsgd-high-prob], see the Appendix for details.
 3.3 Strongly Convex Costs
 In this subsection we provide improved results for the last iterate of strongly convex costs.
-It is well known that strong convexity implies a unique global minimizer of $f$, see, e.g., [74].
+It is well known that strong convexity implies a unique global minimizer of $f$, see, e.g., [nesterov-lectures_on_cvxopt].
 Denote the global minimizer by $x^{\star}\in\mathbb{R}^{d}$, noting that $f^{\star}\coloneqq f(x^{\star})$ and let $\kappa\coloneqq\frac{L}{\mu}\geq 1$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively.
-Next, recall that DSGD does not require a heterogeneity bound to converge in the MSE sense for strongly convex costs, e.g., [20, 11].
+Next, recall that DSGD does not require a heterogeneity bound to converge in the MSE sense for strongly convex costs, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a].
 To show similar benefits in the HP sense, a different approach to Lemma 3 is required.
@@
 where $S_{k}\coloneqq\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}\coloneqq\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{5}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{6}}\Big)$.
-Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring bounded heterogeneity.
+Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5).
 Using Lemma 6 from the Appendix, it can be shown that the RHS is of the order $e^{\mathcal{O}(\nu K_{t+1}\alpha_{t}^{2})}$.
 We note that the requirement of same model initialization across users can be removed, at the cost of an additional term in the exponent on the RHS, of order $\mathcal{O}(\nu K_{t+1}\lambda^{t}\Delta_{x})$, which decays geometrically (recall that $\lambda\in[0,1)$).
-However, we make the assumption for ease of exposition.
+However, we assume same model initialization, for ease of exposition.
 Prior to stating the main theorem, we provide an important technical result, which allows us to establish sharper bounds and ensure linear speed-up in the number of users is achieved.
@@
 Lemma 5 provides a bound on the MGF of an “almost decreasing” process.
-Compared to bounds used in the centralized setting, e.g., [31, 32, 40], Lemma 5 gives a more fine-grained result, allowing us to incorporate higher-order terms in the final rate.
+Compared to bounds used in the centralized setting, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], Lemma 5 gives a more fine-grained result, allowing us to incorporate higher-order terms in the final rate, which is crucial for establishing linear speed-up, see Section 4 and the Appendix for detailed discussions.
 We are now ready to state the main result for strongly convex costs.
@@
 In the Appendix we provide the full bound, containing additional terms, of order $\mathcal{O}\big(\frac{\log(T+t_{0}))}{(T+t_{0})^{3}}\big)$.
-We can see that the leading term in Theorem 2 decays at a rate $\mathcal{O}\big(\frac{d\sigma^{2}\kappa}{n(T+t_{0})}\big)$, achieving linear speed-up, while the network, problem and heterogeneity dependence are captured through the higher-order term $\mathcal{O}\big(\frac{\lambda^{2}\kappa^{2}\|\nabla\mathbf{f}^{\star}\|^{2}}{(1-\lambda)^{2}n(T+t_{0})^{2}}\big)$.
-Crucially, the term without linear speed-up, i.e., $\mathcal{O}\big(\frac{\sigma^{2}}{(T+t_{0})^{2}}\big)$, achieves a faster decay, stemming from the improved result in Lemma 5.
-Compared to [57], where the authors establish the $\mathcal{O}\big(\frac{1}{T}\big)$ convergence rate in HP for PL costs, Theorem 2 provides several improvements.
-First, in [57] the authors require uniformly path-wise bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq C_{i}$, for every $t\geq 1$ almost surely, which is difficult to guarantee uniformly across all sample paths.
-Secondly, they impose the following condition on the MGF of the noise
+Next, we discuss multiple aspects of our results.
+Centralized SGD and linear speed-up.
+Compared to centralized SGD, which achieves the HP rate $\mathcal{O}\Big(\frac{1}{T}+\frac{1}{T^{2}}\Big)$, e.g., [liu2024revisiting], the rate in Theorem 2 also consists of terms of different order, namely $\mathcal{O}\Big(\frac{1}{nT}+\frac{1}{T^{2}}\Big)$.
+Similarly to non-convex results, the leading term shows a linear speed-up in the number of users, again matching the corresponding MSE guarantees, e.g., [pmlr-v119-koloskova20a].
+Network effect and transient time.
+We can see that network connectivity again only affects higher-order terms.
+Next, we can use the rate in Theorem 2 to establish a transient time of DSGD in the HP sense.
+As shown in the Appendix, our HP results imply a transient time of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$, which is strictly sharper than the transient time of DSGD in, e.g., [unified-refined, Table II], stemming from MSE rates, which is of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$.
+The sharper transient time further highlights the tightness of our results and an improved analysis.
+Effects of noise and heterogeneity.
+While heterogeneity again affects only higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see the Appendix for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule.
+Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size.
+Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
+Comparison with [dsgd-high-prob].
+The authors in [dsgd-high-prob] establish HP convergence of DSGD for the more general case of PŁ costs, providing a rate of $\mathcal{O}\big(\frac{1}{T}\big)$, using the same time-varying step-size schedule considered in Theorem 2.
+Crucially, the authors require two very strong assumptions: $(i)$ path-wise uniformly bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq G$, for all $t\geq 1$ and some $G>0$, almost surely; $(ii)$ asymptotically vanishing noise, i.e., for all $i\in[n]$ and any $t\geq 1$, the authors require the following condition
 $\mathbb{E}\left[\exp\bigg(\frac{\|z_{i}^{t}\|^{2}}{\alpha_{t}^{2}\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1),$
 where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size.
-Such condition implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that the noise asymptotically vanishes, at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$.
+Both conditions are very difficult to satisfy in practice, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is in general not satisfied even for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that the noise asymptotically vanishes, at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$.
 [^11] It can be easily verified via Markov’s inequality that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$.
 Similarly, it follows from Jensen’s inequality that $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$.
-We remove both requirements, showing that HP convergence of the last iterate of strongly convex functions is guaranteed without bounded gradients and under standard sub-Gaussian noise, while also achieving linear speed-up in the number of users.
-Finally, we believe that the techniques used in our proof can be leveraged to show HP guarantees for the broader class of PL costs, which is left for future work.
+Crucially, the authors in [dsgd-high-prob] are unable to establish linear speed-up in the number of users, even though much stronger assumptions on the noise and cost are imposed.
+On the other hand, our work uses standard assumptions, while simultaneously showing optimal convergence rates with linear speed-up and improved transient times compared to MSE results.
+These improvements stem not only from a sharper analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity assumption (A5) and ensure linear speed-up is achieved for strongly convex costs (see the Appendix for a detailed discussion).
 4 Proof Outlines and Discussion
@@
 We can use Lemma 3 to control the last term on the RHS.
-To deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$, while removing the need for uniformly bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [39, 33, 46], by subtracting $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of $\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ are absorbed by the left-hand side, allowing us to show that the resulting MGF is bounded.
+To deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$, while removing the need for uniformly bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [li2020high, liu2023high, armacki2025optimal], by subtracting $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of $\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ are absorbed by the left-hand side, allowing us to show that the resulting MGF is bounded.
 The rest of the proof then relies on Lemma 2, some technical results, careful selection of the step-size and Hölder’s inequality, see the Appendix for details.
@@
 Results of Theorems 1 and 2 are strong, demonstrating that DSGD converges in the HP sense under the same conditions on the cost function as in the MSE sense, while retaining order-optimal rates and linear speed-up in the number of users.
-Compared to works studying HP convergence of centralized SGD, e.g., [31, 32, 33, 40], we make several contributions.
+Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we make several contributions.
 First, we face the challenge of controlling the MGF of the consensus gap, stemming from the decentralized nature of the algorithm.
-To that end, we provide Lemma 3 for non-convex costs, relying on the bounded heterogeneity condition, and the improved Lemma 4 for strongly convex costs, which removes the bounded heterogeneity condition.
-Next, we provide Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense, which is of independent interest for HP studies of decentralized methods.
-To ensure linear speed-up in the number of users is achieved for strongly convex costs, we establish Lemma 5, which gives a more fine-grained result on the MGF [31, 32, 40], facilitating higher-order terms in the final bound, and is of independent interest.
-Moreover, to be able to utilize Lemma 5, a careful analysis of the MGF is needed, in order to simultaneously balance the effects of optimality and consensus gaps (recall that (5), which is used to bound the MGF of $F^{t+1}$, consists of both optimality and consensus gaps), while also ensuring that the “almost decreasing” property is maintained.
+To that end, we provide Lemma 3 for non-convex costs, removing the uniformly bounded gradient condition and using the bounded heterogeneity condition (A5) instead, as well as the improved Lemma 4 for strongly convex costs, which removes (A5) altogether.
+Next, we provide Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense and is of independent interest for HP studies of decentralized methods.
+To ensure linear speed-up in the number of users is achieved for strongly convex costs, we establish Lemma 5, which gives a more fine-grained result on the MGF of an “almost decreasing” process compared to existing (centralized) works, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], facilitating higher-order terms in the final bound, and is of independent interest (see the Appendix for a detailed discussion on the importance of Lemma 5 for achieving linear speed-up for strongly convex costs).
+Finally, to be able to utilize Lemma 5, a careful analysis of the MGF is needed, in order to simultaneously balance the effects of optimality and consensus gaps (recall that (5), which is used to bound the MGF of $F^{t+1}$, consists of both optimality and consensus gaps), while also ensuring that the resulting process is “almost decreasing”, in the sense of Lemma 5.
 On the dimension dependence.
-We note that our bounds show a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for either MSE or HP bounds, e.g., [11, 57].
-This dependence stems from the third result in Lemma 2, where the problem dimension shows up in the sub-Gaussianity constant, which is unavoidable, due to the nature of the definition of sub-Gaussianity in (A4).
+We note that our bounds show a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for either MSE or HP bounds, e.g., [pmlr-v119-koloskova20a, dsgd-high-prob].
+This dependence stems from the third result in Lemma 2, where the dimension figures in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4).
 In particular, to show the third property in Lemma 2, we use the fact that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-norm-sub-Gaussian, as a direct proof fails to yield the variance reduction benefit in the number of users.
-One way to avoid the (linear) dependence on $d$ is to assume that the noise is norm-sub-Gaussian, which would simultaneously relax the dependence on the dimension to $\log(d)$ and achieve the variance reduction benefit, at the cost of imposing a slightly stronger noise condition, see, e.g., Corollary 7 and the conclusion in [75].
-[^10] See the Appendix for the full proof of Lemma 2 and further discussion on the dimension dependence.
+One way to avoid the (linear) dependence on $d$ is to assume that the noise is norm-sub-Gaussian, which would simultaneously relax the dependence on the dimension to $\log(d)$ and achieve the variance reduction benefit, at the cost of imposing a slightly stronger noise condition, see, e.g., Corollary 7 and the conclusion in [jin2019short]; see also the Appendix for a further discussion on the dimension dependence.
 5 Conclusion
-In this paper we revisit convergence in HP of a variant of DSGD under sub-Gaussian noise.
+In this paper, we revisit convergence in HP of a variant of DSGD under sub-Gaussian noise.
 We show that DSGD is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates for both non-convex and strongly convex costs.
 Moreover, our results show DSGD achieves linear speed-up in the number of users in both cases.
-Compared to [57], wherein the authors study HP convergence of DSGD, we relax strong conditions like uniformly bounded gradients or asymptotically vanishing noise, while improving the rates by showing linear speed-up in the number of users.
-Compared to works studying HP convergence of centralized SGD, e.g., [31, 32, 33, 40], we provide tight control on the MGF of the consensus gap for both non-convex and strongly convex costs, in Lemmas 3 and 4, as well as a more fine-grained bound on the MGF of an “almost decreasing” process, in Lemma 5.
-Future work includes extending our results to costs satisfying the PL condition, incorporating the GT mechanism to remove bounded heterogeneity condition and considering noise with heavier tails, like sub-Weibull.
+Compared to [dsgd-high-prob], wherein the authors study HP convergence of DSGD, we relax strong conditions like uniformly bounded gradients or asymptotically vanishing noise, while showing improved rates, with linear speed-up in the number of users and transient times matching, or strictly sharper than transient times stemming from corresponding MSE rates.
+Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we provide several novel results, establishing tight control on the MGF of the consensus gap for both non-convex and strongly convex costs in Lemmas 3 and 4, a more fine-grained bound on the MGF of an “almost decreasing” process in Lemma 5, as well as showing variance reduction benefits of decentralized algorithms (in the HP sense) in Lemma 2.
+Future work includes extending our results to costs satisfying the PŁ condition, incorporating the GT mechanism to remove bounded heterogeneity condition and considering noise with heavier tails, like sub-Weibull.
 Appendix A Introduction
 The Appendix contains results omitted from the main body.
-Section B collects some important facts used in our proofs, Section C provides some technical results, Section D defines some notions used in the analysis, Sections E and F provide proofs for non-convex and strongly convex costs, respectively, while Section G contains further discussion on dimension dependence in our bounds.
+Section B collects some important facts used in our proofs, Section C provides some technical results, Section D defines some notions used in the analysis, Sections E and F provide proofs for non-convex and strongly convex costs, respectively, Section G provides numerical results, Section H derives the transient times, Section I provides a discussion on the novel results needed for achieving linear speed-up, Section J discusses the extension to time-varying networks, while Section K contains further discussion on dimension dependence in our bounds.
 Appendix B Useful Inequalities
@@
 $\mathbb{E}\Big[\prod_{i\in[n]}|X_{i}|\Big]\leq\prod_{i\in[n]}\sqrt[n]{\mathbb{E}|X_{i}|^{n}}.$
-Next, we state a useful result from [25].
-Proposition 6 ([25], Lemma 21).
+Next, we state a useful result from [ran-improved].
+Proposition 6 ([ran-improved], Lemma 21).
 For any $c,t_{0}>0$ and $0\leq a\leq b$, we have
 $\prod_{k=a}^{b}\Big(1-\frac{c}{t+t_{0}}\Big)\leq\frac{(a+t_{0})^{c}}{(b+1+t_{0})^{c}}.$
-The following result states some important consequences of conditions (A2) and (A3), see, e.g., [76, 74].
+The following result states some important consequences of conditions (A2) and (A3), see, e.g., [bertsekas-gradient, nesterov-lectures_on_cvxopt].
 Proposition 7.
@@
     $\|\nabla f(x)\|^{2}\leq 2L(f(x)-f^{\star})$.
-Finally, we complete this section with the following important consequences of (A6), see, e.g., [74].
+Finally, we complete this section with the following important consequences of (A6), see, e.g., [nesterov-lectures_on_cvxopt].
 Proposition 8.
-Let (A6) hold. Then, the following are true, for all $i\in[n]$ and $x,y\in\mathbb{R}^{d}$.
+Let (A6) hold. Then, the following are true, for all $i\in[n]$ and $x\in\mathbb{R}^{d}$.
   1. 1.
-    $\langle y,\nabla^{2}f_{i}(x)y\rangle\|\geq\mu\|y\|^{2}$.
+    $\nabla^{2}f_{i}(x)\succeq\mu I_{d}$.
   2. 2.
@@
   1. 1.
-    To prove the first property, we follow steps similar to those in [39, Lemma 1].
+    To prove the first property, we follow steps similar to those in [li2020high, Lemma 1].
 Let $y_{i}=\frac{z^{t}_{i}}{\sigma_{i}}$ and note that $\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\leq\exp(1)$, from (A3).
@@
   3. 3.
-    Combining the previous result with Lemma 1 from [75], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
+    Combining the previous result with Lemma 1 from [jin2019short], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
     $\mathbb{P}\big(\|\overline{z}^{t}\|>\epsilon\big)\leq 2\exp\bigg(-\frac{n\epsilon^{2}}{24\sigma^{2}d}\bigg).$
-    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., [77, Proposition 2.5.2], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
+    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., [vershynin_2018, Proposition 2.5.2], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
 We next restate Lemma 5.
@@
 Combining the definition of the network average model, the fact that the weight matrix is doubly stochastic and (12), it follows that $\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t}$.
-We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [13, 18, 25].
+We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [nedic-subgrad, jakovetic-unification, ran-improved].
 Let ${\mathbf{x}}^{t}\coloneq col(x_{1}^{t},\ldots,x_{n}^{t})\in\mathbb{R}^{nd}$ denote the column vector stacking users’ local models.
@@
 [^13] This is equivalent to the client-server setup in terms of the update rule, as the server averages models from all clients in each iteration.
-The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [18, 25] and references therein.
+The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [jakovetic-unification, ran-improved] and references therein.
 Proposition 9.
@@
 Proof.
-The proof follows similar steps as in [25, Lemma 3]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
+The proof follows similar steps as in [ran-improved, Lemma 3]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
 $f(x)\leq f(y)+\langle\nabla f(y),x-y\rangle+\frac{L}{2}\|x-y\|^{2}.$ (15)
@@
 Let (A1)-(A5) hold.
-If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$, where $C>0$ is a problem related constant satisfying
+If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
 $C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
 we then have, for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+L\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}[\Delta_{x}+n\lambda^{2}(A^{2}+\sigma^{2})]}{(1-\lambda^{2})T}\Bigg).$
+$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}$
+$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
 Proof.
@@
     $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\alpha\sigma^{2}dL}{n}+\frac{48\alpha^{2}\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}.$
-    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{dT}}$, then with probability at least $1-\delta$
-    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\sigma L\sqrt{d}}{\sqrt{nT}}+\frac{48n\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
-    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{dT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{dT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma}{\sqrt{nT}}+\frac{1}{CT}$.
+    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{dLT}}$, then with probability at least $1-\delta$
+    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
+    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{dLT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{dLT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{1}{CT}$.
 Therefore, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we finally have
-    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{d}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+120L\big)}{\sqrt{nT}}$
-    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{48n\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
+    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+120\big)}{\sqrt{nT}}$
+    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
   2. 2.
@@
     Note that in the unknown time horizon case we lose the linear speed-up in the number of users, as, even in the case $C=\frac{\sqrt{n}}{6\sigma\sqrt{10dL}}$, we have
-    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{T+1}}{{\delta}})\big)}{\sqrt{n(T+1)}}+\frac{\sigma\sqrt{d}\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
+    $\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{T+1}}{{\delta}})\big)}{\sqrt{n(T+1)}}$
+    $+\frac{\sigma\sqrt{d}\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
     as the last term, which does not attain a linear speed-up, is no longer of higher order.
-We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [11, 25].
+We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [pmlr-v119-koloskova20a, ran-improved].
 Appendix F Proofs for Strongly Convex Costs
 In this section we prove Lemma 4 and Theorem 2.
-To do so, we follow a similar strategy to the one in, e.g., [20, 25], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense.
+To do so, we follow a similar strategy to the one in, e.g., [conv-rates-dsgd-jakovetic, ran-improved], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense.
 However, to establish guarantees in the HP sense, we instead work with the MGF of the iterates and have the following important result.
@@
 If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
-$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+d\sigma^{2}\kappa/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+\kappa^{2})}}{{(1-\lambda)}})}{(1-\lambda)n(T+t_{0})^{2}}$
-$+\frac{\Delta_{f}}{n(T+t_{0})^{3}}+\frac{\sigma^{2}\lambda^{2}L^{3}(L\log(t+t_{0})+1)}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{\lambda^{2}L^{3}(1+L)\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(T+t_{0})^{3}}\bigg).$
+$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
+$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
+$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
 Proof.
@@
 where $N_{t}\coloneqq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}$.
-Define $G_{1}\coloneqq 240a^{2}\sigma^{2}dL$, $G_{2}\coloneqq\frac{6a^{3}\lambda^{2}L^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)}$, $G_{3}\coloneqq\frac{48a^{5}n\sigma^{2}\lambda^{2}L^{4}}{(1-\lambda)^{2}}$ and $G_{4}\coloneqq\frac{30a^{3}\lambda^{2}L^{4}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and plug into (F), to get
+Define $G_{1}\coloneqq\frac{8640\kappa\sigma^{2}d}{\mu}$, $G_{2}\coloneqq\frac{1296\kappa^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)}$, $G_{4}\coloneqq\frac{6480\kappa^{3}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and $G_{3}\coloneqq\frac{373248n\kappa^{4}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}}$, and plug into (F), to get
 $\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]\leq\Big(\mathbb{E}\big[\exp\big(\nu F_{t}\big)\big]\Big)^{b_{t}}\exp\bigg(\sum_{i\in[3]}\frac{\nu G_{i}}{(t+t_{0})^{i}}+\frac{\nu G_{4}}{(t+t_{0})^{8}}\bigg).$
@@
 Therefore, it can be readily seen that, choosing
-$\epsilon_{t}^{2}=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})}{n(t+t_{0}+1)}$ $+\frac{4a^{2}\lambda^{2}L(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)n(t+t_{0})^{2}}$
-$+\frac{32a^{4}\sigma^{2}\lambda^{2}L^{3}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{3}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}},$ (44)
+$\epsilon_{t}^{2}=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})}{n(t+t_{0}+1)}$ $+\frac{144\kappa\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)n(t+t_{0})^{2}}$
+$+\frac{41472\kappa^{3}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{720\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}},$ (44)
 we get $\mathbb{P}\big(B_{t,\epsilon_{t}^{2}}\big)\leq\delta$, for any $\delta\in(0,1)$.
@@
 Therefore, we get, for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)$ $=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(t+t_{0})^{2}}$
-$+\frac{\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\sigma^{2}\lambda^{2}L^{3}(L\log(t+t_{0})+1)}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{\lambda^{2}L^{3}(1+L)\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}\bigg).$
+$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
+$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
+$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
 Finally, it can be verified that the conditions on $a$, $t_{0}$ and $\nu$ in the statement of the theorem ensure that all the step-size conditions are satisfied, completing the proof. ∎
-We remark that, similarly to the proof of Theorem 1, one can analyze the strongly convex case with a fixed step-size, resulting in the dependence on some terms, e.g., optimality and iterate gaps $\Delta_{f}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}}^{\star}\|$, decaying exponentially fast, i.e., $\mathcal{O}\big((\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}+\Delta_{f})e^{-CT}\big)$, for some $C>0$, as shown in, e.g., [11] for MSE guarantees in decentralized, or [40] for HP guarantees in centralized settings.
+We remark that, similarly to the proof of Theorem 1, one can analyze the strongly convex case with a fixed step-size, resulting in the dependence on some terms, e.g., optimality and iterate gaps $\Delta_{f}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}}^{\star}\|$, decaying exponentially fast, i.e., $\mathcal{O}\big((\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}+\Delta_{f})e^{-CT}\big)$, for some $C>0$, as shown in, e.g., [pmlr-v119-koloskova20a] for MSE guarantees in decentralized, or [liu2024revisiting] for HP guarantees in centralized settings.
 For simplicity, we omit this analysis.
-Appendix G On Dimension Dependence
+Appendix G Numerical Experiments
+In this section we provide some numerical results, which serve to verify our theory.
+Methodology.
+We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive semi-definite, making each $f_{i}$ strongly convex, with unbounded gradients.
+We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [XIAO200465].
+When queried by user $i\in[n]$ in iteration $t\geq 1$, the $\mathcal{SFO}$ returns a noisy gradient of $f_{i}$ evaluated at $x^{t}_{i}$, i.e., $g^{t}_{i}=A_{i}x^{t}_{i}+b_{i}+z^{t}_{i}$, where $z^{t}_{i}\in\mathbb{R}^{d}$ is a zero-mean Gaussian random vector, making the noise consistent with assumption (A4).
+We use the DSGD method outlined in Algorithm 1, with the time-varying step-size schedule $\alpha_{t}=\frac{1}{t+1}$ and shared initializatio $x^{1}_{i}=0$, for all $i\in[n]$.
+We are interested in testing the following two facets of our theory.
+  1. 1.
+    Exponentially decaying tails - we want to verify that the tail probability decays at an exponential scale, as defined in (1) and predicted in Theorem 2.
+  2. 2.
+    Linear speed-up - we want to verify that the tail probability decays faster as the number of users increases, as predicted in Theorem 2.
+To verify these two facets, we measure the performance in terms of the empirical tail probability $\mathbb{P}^{t}_{n,\varepsilon}$, computed as follows.
+We first run DSGD for $T$ iterations and repeat it over $R$ runs.
+Next, for each $t\in[T]$, we use Monte-Carlo sampling to create a set $S_{t}$ of indices from $[R]$, of some fixed length $|S_{t}|=S$.
+For any $\varepsilon>0$, the empirical probability is computed as
+$\mathbb{P}^{t}_{n,\varepsilon}=\frac{1}{S}\sum_{r\in S_{t}}\mathbb{I}\bigg(\frac{1}{n}\sum_{i\in[n]}\|x_{i}^{t,r}-x^{\star}\|^{2}>\varepsilon\bigg),$
+where $x_{i}^{t,r}\in\mathbb{R}^{d}$ is the model of user $i$ in iteration $t$ and run $r$, with $x^{\star}=\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}f(x)$ being the solution of the global problem and $\mathbb{I}(A)$ being the indicator of event $A$.
+The empirical tail probability is a proxy to the true tail probability and is the main metric in our experiments.
+To further illustrate our results, we also compute the empirical mean-squared error $\mathbb{E}^{t}_{n}$, which is the optimality gap at time $t$, averaged across all users and runs, i.e., $\mathbb{E}^{t}_{n}=\frac{1}{nR}\sum_{i\in[n]}\sum_{r\in[R]}\|x_{i}^{t,r}-x^{\star}\|^{2}$.
+We next present the results.
+Exponentially decaying tails.
+To verify that the (empirical) tail probability decays at an exponential rate, we consider a fixed network of $n=30$ users.
+For local costs, each matrix $A_{i}$ is generated using python’s sklearn library function make_sparse_spd_matrix, with dimension $d=50$ and value $\texttt{alpha}=0.9$, while vectors $b_{i}$ are drawn from a multivariate normal distribution $\mathcal{N}(\mathbf{0}_{d},\sigma_{i}^{2}I_{d})$, where $\sigma_{i}^{2}=1$ for $i\in\{1,\ldots,10\}$, $\sigma_{i}^{2}=2$ for $i\in\{11,\ldots,20\}$, and $\sigma^{2}_{i}=4$ for $i\in\{21,\ldots,30\}$, ensuring that the data is heterogeneous across users.
+We run DSGD for $T=10000$ iterations and repeat across $R=5000$ runs.
+To compute the empirical probability, in each iteration we draw $S=3000$ Monte-Carlo samples and consider threshold values $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$.
+The results are presented in Figure 1, where the left plot shows the MSE behaviour, while the right plot shows the tail probability behaviour.
+We can see that the empirical tail probability decays exponentially fast for all values of $\varepsilon$, as predicted by our theory.
+Note that for the threshold value $\varepsilon=10^{-4}$, the tail probability starts decaying exponentially after approximately $t=6000$ iterations, which is consistent with the MSE behaviour on the left plot, where we can see that it takes DSGD around the same number of iterations to reach the average accuracy $\mathbb{E}_{n}^{t}=10^{-4}$.
+Refer to caption Refer to caption
+Figure 1: Performance of DSGD, in the MSE sense (left) and HP sense (right).
+We can see that DSGD achieves an exponential tail decay for all values of threshold $\varepsilon$.
+For the threshold $\varepsilon=10^{-4}$, the tail probability starts decaying exponentially after approximately $t=6000$ iterations, which is consistent with the MSE behaviour, where we can see that DSGD takes around the same number of iterations to reach the average accuracy $\mathbb{E}^{t}_{n}=10^{-4}$.
+Linear speed-up.
+To verify that linear speed-up in the number of users is achieved, we consider three networks, with $n=\{25,50,100\}$ users.
+In all three cases we use Erdős–Rényi communications graphs which satisfy $\lambda=\|W-J\|\approx 0.6$.
+To ensure that the effect of heterogeneity is consistent across different networks, we fix the matrices $A_{i}$, i.e., $A_{i}\equiv A=S+I_{d}$, where $S\in\mathbb{R}^{d\times d}$ is generated using sklearn’s make_sparse_spd_matrix, with $d=50$ and $\texttt{alpha}=0.9$, while vectors $b_{i}$ are given by $b_{i}=\beta_{i}\mathbf{1}_{d}$, where $\beta_{i}\in\{-2,-1,0,1,3\}$ are selected uniformly at random and in equal proportion across users (i.e., one fifth of users has $\beta_{i}=-2$, one fifth $\beta_{i}=-1$, etc).
+Generating the networks and costs in this manner ensures that both network connectivity and user heterogeneity are constant across all three network settings, allowing us to properly capture the effect of linear speed-up.
+We run DSGD for $T=3000$ iterations and repeat across $R=1000$ runs.
+To compute the empirical probability, in each iteration we draw $S=600$ Monte-Carlo samples, with thresholds $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$.
+The results are presented in Figure 2, where the plots left to right and top to bottom show the MSE and tail probability behaviours for different values of $\varepsilon$, respectively.
+We can again see that the empirical tail probability decays exponentially fast for all values of $\varepsilon$, with the decay being consistently faster for larger number of users, demonstrating linear speed-up in the HP sense.
+Finally, note that in the lower right plot (i.e., threshold $\varepsilon=10^{-4}$) the network with $n=25$ users has a constant tail probability $\mathbb{P}^{t}_{n,\varepsilon}=1$, while for the network with $n=50$ the tail probability is slowly starting to decrease, which is again consistent with the MSE behaviour on the upper left plot, as the accuracy $\mathbb{E}^{t}_{n}=10^{-4}$ for these two networks is not reached (on average) in the allocated number of iterations.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 2: Linear speed-up of DSGD, in the MSE and HP sense.
+Left to right and top to bottom: MSE performance and tail decay with threshold $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$.
+We can see that DSGD consistently achieves faster exponential tail decay for larger networks, across all values of threshold $\varepsilon$, illustrating the effect of the linear speed-up in the HP sense.
+Appendix H Deriving the Transient Time
+In this section we provide details on the transient time resulting from our HP bounds.
+Similarly to [unified-refined], we focus on the dependence on network connectivity and the number of users, ignoring other problem related constants.
+First, consider non-convex costs.
+Recall that the bound in Theorem 1 is of the form
+$\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}+\frac{1}{CT}+\frac{\Delta_{x}}{(1-\lambda^{2})T}+\frac{n}{(1-\lambda)^{2}T}\bigg),$ (45)
+where the problem related constant $C$ depends on the network connectivity via the expression $C=\mathcal{O}\Big(\min\Big\{\frac{1-\lambda}{\lambda},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\lambda^{2}}\Big\}\Big)$.
+Since we are interested in the worst-case transient time (corresponding to poor network connectivity, i.e., $\lambda\approx 1$), it follows that $C=\mathcal{O}\Big(\frac{1-\lambda}{\lambda}\Big)$.
+For simplicity, assume that users have a shared initialization, i.e., $x^{1}_{i}=x_{j}^{1}$, for all $i,j\in[n]$, so that $\Delta_{x}=0$.
+In this case, the bound in (45) becomes
+$\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}+\frac{1}{(1-\lambda)T}+\frac{n}{(1-\lambda)^{2}T}\bigg).$ (46)
+It is now evident that, for any $T\geq\max\Big\{\frac{n}{(1-\lambda)^{2}},\frac{n^{3}}{(1-\lambda)^{4}}\Big\}=\frac{n^{3}}{(1-\lambda)^{4}}$, the bound in (46) becomes $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$.
+This matches the transient time of DSGD obtained from the MSE rate in [pmlr-v119-koloskova20a], see, e.g., [unified-refined, Table I].
+Next, consider strongly convex costs. Recalling the expressions in (42) and (F), it follows that the full bound for strongly convex costs is of the form
+$\mathcal{O}$ $\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}+\frac{t_{0}^{3}}{n(T+t_{0})^{3}}$
+$+\frac{\log(T+t_{0})}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{t_{0}}{n(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{t_{0}^{6}}{n(1-\lambda)^{2}(T+t_{0})^{8}}\bigg),$ (47)
+where $t_{0}$ depends on the network connectivity via the expression $t_{0}=\Omega\Big(\max\Big\{\frac{\lambda}{1-\lambda},\frac{\lambda^{2}}{1-\lambda},\frac{1}{1-\lambda}\Big\}\Big)$.
+Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{\lambda}{1-\lambda}\Big)$, so that the bound in (H) becomes
+$\quad\quad\mathcal{O}\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}$
+$+\frac{\log(T+t_{0})}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{3}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{8}(T+t_{0})^{8}}\bigg).$ (48)
+It can be seen that, for any $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (H) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$.
+Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [pmlr-v119-koloskova20a], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., [unified-refined, Table II].
+As such, our HP analysis provides either matching or strictly sharper transient times for DSGD compared to the transient times stemming from the corresponding MSE results, further highlighting the tightness of our analysis.
+Appendix I Achieving Linear Speed-up
+In this section we provide further discussions on the main results and techniques needed to obtain linear speed-up, for both non-convex and strongly convex costs.
+Linear speed-up for non-convex costs.
+The crucial result for obtaining linear speed for non-convex costs is Lemma 2, which establishes the variance reduction benefits of decentralized methods in the HP sense.
+To see why this is the case, note that using a fixed step-size $\alpha_{t}\equiv\alpha$ and following the same steps as in the proof of Theorem 1, while relying only on assumption (A4) to control the noise, one would get a final bound where the leading term is of the form
+$\mathcal{O}\bigg(\frac{1}{\alpha T}+\alpha\sigma^{2}\bigg).$ (49)
+It is obvious that no choice of $\alpha$ leading to a linear speed-up is possible, i.e., such that (49) becomes $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}\Big)$.
+On the other hand, using Lemma 2 to handle the noise, we are able to establish a bound where the leading term is of the form
+$\mathcal{O}\bigg(\frac{1}{\alpha T}+\frac{\alpha\sigma^{2}}{n}\bigg),$ (50)
+where it is easy to see that choosing $\alpha=\Theta\Big(\frac{\sqrt{n}}{\sqrt{T}}\Big)$ leads to the desired linear speed-up.
+Works like [dsgd-high-prob, dsgd-online-noncooperative-games, dsgd-online-stochastic] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity.
+As such, even if a fixed step-size was used, the authors would be unable to obtain a bound of the form (50) and achieve linear speed-up.
+As discussed in Appendix E, using a time-varying step-size for non-convex costs inevitably leads to a loss of linear speed-up, as all terms decay at the same rate, leading to a bound of the form
+$\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}+\frac{n}{\sqrt{T}}\bigg).$
+On the other hand, using a fixed step-size $\alpha=\Theta\Big(\frac{\sqrt{n}}{\sqrt{T}}\Big)$, we get the bound
+$\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}+\frac{n}{T}\bigg),$
+where the term not achieving linear speed-up is of higher order.
+This is consistent with MSE guarantees, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined], where a fixed step-size is also needed to achieve linear speed-up.
+Linear speed-up for strongly convex costs.
+For strongly convex costs, in addition to Lemma 2, a tighter control of the MGF of the process of interest is required, which is achieved in Lemma 5.
+To see why this is the case, consider the existing techniques in centralized settings, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a], where the authors show via induction that the MGF of the process of interest is uniformly bounded, i.e., that there exists a $B>0$, such that, for all $t\geq 1$
+$\mathbb{E}[\exp(BF^{t})]\leq\exp(1).$ (51)
+Recalling that in our work $F^{t}=n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$, it then follows from (51) and Markov’s inequality that, for any $\epsilon>0$
+$\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}>\frac{\epsilon+1}{nB(t+t_{0})}\bigg)=\mathbb{P}\big(BF^{t}>\epsilon+1\big)\leq\exp(-\epsilon),$
+or equivalently, with probability at least $1-\delta$
+$f(\overline{x}^{t})-f^{\star}\leq\frac{\log(\nicefrac{{1}}{{\delta}})+1}{nB(t+t_{0})}.$ (52)
+While the above bound seemingly achieves linear speed-up, we now show that the constant $B$ inevitably depends on $n$, effectively canceling the linear speed-up.
+To see that this is the case, we go back to (51) and note that, in order for the identity to hold for $t=1$, we have
+$\mathbb{E}[\exp(BF^{1})]=\exp\Big(nB(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)\Big)\leq\exp(1),$
+where the equality follows from the definition of $F^{t}$ and the fact that the initial models are chosen deterministically, while the inequality follows if and only if $B\leq\frac{1}{n(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)}$.
+As such, in order for (51) to hold, the constant $B$ has to decay with $n$, canceling the linear speed-up in (52) and resulting in a HP bound of the form
+$f(\overline{x}^{t})-f^{\star}\leq\frac{\log(\nicefrac{{1}}{{\delta}})+1}{B^{\prime}(t+t_{0})},$
+where $B^{\prime}>0$ is independent of $n$ and linear speed-up is lost. On the other hand, Lemma 5 implies the following bound on the MGF
+$\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]$ $\leq\exp\bigg(\frac{(t_{0}+2)^{3}\nu\Delta_{f}}{(t+1+t_{0})^{2}}+4\nu G_{1}+\frac{4\nu G_{2}}{t+1+t_{0}}\bigg)$
+$\times\exp\bigg(\frac{4\nu G_{3}\log(t+t_{0}+1)}{(t+t_{0}+1)^{2}}+\frac{4\nu G_{4}}{5(t_{0}+1)^{5}(t+1+t_{0})^{2}}\bigg),$
+for some problem related constants $G_{i}>0$, $i\in[4]$, providing a much tighter and more fine-grained control of the MGF compared to (52) and crucially allowing for a bound where the leading term achieves linear speed-up in the number of users (as $G_{1}$ is independent of $n$).
+On the other hand, using a different approach to [pmlr-v99-harvey19a, pmlr-v206-bajovic23a], the work [liu2024revisiting] establishes the following HP bound for the last iterate of SGD in centralized setting and smooth, strongly convex costs
+$f(x^{t})-f^{\star}=\mathcal{O}\bigg(\frac{\log(t)\log(\nicefrac{{1}}{{\delta}})}{t+t_{0}}+\frac{1}{(t+t_{0})^{2}}\bigg).$
+While such result could potentially be leveraged in decentralized settings to maintain linear speed-up (provided that the constant growing linearly with $n$ affects only the higher-order term), it still accrues a factor of $\log(t)$ in the leading term, whereas in our bound it appears only in the higher-order term, decaying at rate $\mathcal{O}\Big(\frac{\log(t+t_{0})}{(t+t_{0})^{3}}\Big)$.
+As such, Lemma 5 provides the most fine-grained control of the MGF and is of independent interest when studying “almost decreasing” type processes, even in centralized settings.
+Appendix J Extending to Time-varying Networks
+As discussed in the main body, to model the communication network, we use assumption (A1), which subsumes fixed undirected networks, as well as a class of fixed directed networks.
+The authors in [dsgd-high-prob] consider a more general case of directed, time-varying networks, facilitated by the following condition.
+[…]
+Our work can be readily extended to the setting of time-varying networks, with assumption (A1) replaced by (A1).
+To see why this is the case, note that in our proofs we only use the fact that under (A1), we have $\lambda=\|W-J\|\in[0,1)$.
+The authors in [dsgd-high-prob, Lemma 1] use a similar result for time-varying networks, namely that under (A1), for any time indices $t\geq s$
+$\Big|[\Phi(t,s)]_{ij}-\frac{1}{n}\Big|\leq M\rho^{t-s},$ (53)
+where $\Phi(t,s)=W^{t-1}\times\ldots\times W^{s}$ is the transition matrix from time $s$ to $t$ (with $\Phi(t,t)=I_{n}$), while $M>0$ and $\rho\in(0,1)$ are graph related constants.
+As discussed in [dsgd-high-prob], the constant $\rho$ plays the same role of the network connectivity parameter, as $\lambda$ does in our work.
+It can be readily verified that replacing (A1) with (A1) and using (53) to control the consensus gap, our results readily go through, with $\lambda$ being replaced by $\rho$.
+As such, our results can be easily extended to the (directed) time-varying network setting.
+Appendix K On Dimension Dependence
 As mentioned in Section 4 in the main body, our rates in Theorems 1 and 2 exhibit a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, with the dependence stemming from Lemma 2, where it is shown that $\overline{z}^{t}$ is $\mathcal{O}\big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\big)$-sub-Gaussian.
@@
 $\mathbb{P}\big(\|x\|>\epsilon\big)\leq 2\exp\bigg(-\frac{\epsilon^{2}}{2\sigma^{2}d}\bigg),$
-see [75, Lemma 1].
+see [jin2019short, Lemma 1].
 This argument allows us to show the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian (in the sense of (A4)), at the cost of introducing a $\sqrt{d}$ dependence.
 One way to mitigate this is to directly assume that the noise $z^{t}_{i}$ at each user is zero-mean and $\sigma_{i}$-norm-sub-Gaussian.
-Using a similar argument to the one in Lemma 6 and Corollary 7 in [75], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$.
+Using a similar argument to the one in Lemma 6 and Corollary 7 in [jin2019short], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$.
 This, however, comes at the cost of imposing a slightly stronger noise condition, as any $\sigma$-norm-sub-Gaussian random vector is also $\sigma$-sub-Gaussian (in the inner product sense), while the opposite implication inevitably introduces a $\sqrt{d}$ factor, as is the case in Lemma 2.

### v3

@@
-Improved High-probability Convergence Guarantees of Decentralized SGD
+High-probability Convergence Guarantees of Decentralized SGD
 Abstract
-Convergence in high-probability (HP) has been receiving increasing interest, due to its attractive properties, such as exponentially decaying tail bounds and strong guarantees for each individual run of an algorithm.
-While HP guarantees are extensively studied in centralized settings, much less is understood in the decentralized, networked setup.
-Existing HP studies in decentralized settings impose strong assumptions, like uniformly bounded gradients, or asymptotically vanishing noise, resulting in a significant gap between assumptions used to establish convergence in the HP and the mean-squared error (MSE) sense, even for vanilla Decentralized Stochastic Gradient Descent (DSGD) algorithm.
-This is contrary to centralized settings, where it is known that SGD converges in HP under the same conditions on the cost function as needed to guarantee MSE convergence.
-Motivated by this observation, we revisit HP guarantees for DSGD in the presence of light-tailed noise.
-We show that DSGD converges in HP under the same conditions on the cost as in the MSE sense, removing uniformly bounded gradients and other restrictive assumptions, while simultaneously achieving order-optimal rates for both non-convex and strongly convex costs.
-Moreover, our improved analysis yields linear speed-up in the number of users, demonstrating that DSGD maintains strong performance in the HP sense and matches existing MSE guarantees.
-Our improved results stem from a careful analysis of the MGF of quantities of interest (norm-squared of gradient or optimality gap) and the MGF of the consensus gap between users’ models.
-To achieve linear speed-up, we provide a novel result on the variance-reduction effect of decentralized methods in the HP sense and more fine-grained bounds on the MGF for strongly convex costs, which are both of independent interest.
+Convergence in high-probability (HP) has attracted increasing interest, due to implying exponentially decaying tail bounds and strong guarantees for individual runs of an algorithm.
+While many works study HP guarantees in centralized settings, much less is understood in the decentralized setup, where existing works require strong assumptions, like uniformly bounded gradients, or asymptotically vanishing noise.
+This results in a significant gap between the assumptions used to establish convergence in the HP and the mean-squared error (MSE) sense, and is also contrary to centralized settings, where it is known that SGD converges in HP under the same conditions on the cost function as needed for MSE convergence.
+Motivated by these observations, we study the HP convergence of Decentralized SGD (DSGD) in the presence of light-tailed noise, providing several strong results.
+First, we show that DSGD converges in HP under the same conditions on the cost as in the MSE sense, removing the restrictive assumptions used in prior works.
+Second, our sharp analysis yields order-optimal rates for both non-convex and strongly convex costs.
+Third, we establish a linear speed-up in the number of users, leading to matching, or strictly better transient times than those obtained from MSE results, further underlining the tightness of our analysis.
+To the best of our knowledge, this is the first work that shows DSGD achieves a linear speed-up in the HP sense.
+Our relaxed assumptions and sharp rates stem from several technical results of independent interest, including a result on the variance-reduction effect of decentralized methods in the HP sense, as well as a novel bound on the MGF of strongly convex costs, which is of interest even in centralized settings.
+Finally, we provide experiments that validate our theory.
 1 Introduction
-Modern large-scale machine learning applications and the abundance of data necessitate alternatives to the centralized computation framework, giving rise to distributed learning, a paradigm where multiple users collaborate to jointly train a model, e.g., [sayed-networks, pmlr-v54-mcmahan17a, vlaski_et_al].
-The features of distributed learning, such as storing the data locally and only exchanging smaller updates, like (quantized) model parameters or gradients, and the lack of a single point of failure, further make it an attractive paradigm from a privacy and security perspective [bonawitz2017practical, yu-secure].
-Many applications, such as federated training of models on mobile devices [konecny2016federated], controlling and coordinating robot swarms [bullo2009distributed], or distributed control and power grids [dist-learn-noncvx], all rely on distributed computation.
-From a communication/model exchange perspective, distributed frameworks can be client-server (i.e., federated) or decentralized (i.e., networked), with the main difference being that in the client-server setup users communicate with a server, while in decentralized settings users communicate directly with each other.
-[^1] While the client-server setup is often characterized by performing local updates and periodic communication [stichlocal], as well as partial user participation [pmlr-v151-jee-cho22a], we note that these features can also be incorporated in the decentralized setup, by modifying the methods to perform multiple local updates before communicating [pmlr-v119-koloskova20a] and considering a dynamic network with users that are occasionally idle [Cannelli2020].
+Modern large-scale machine learning applications and the abundance of data necessitate alternatives to the centralized computation framework, giving rise to distributed learning, a paradigm where multiple users collaborate to jointly train a model, e.g., [63, 49, 68].
+The features of distributed learning, such as storing the data locally and only exchanging smaller updates, like (quantized) model parameters or gradients, and the lack of a single point of failure, further make it an attractive paradigm from a privacy and security perspective [10, 80].
+Many applications, such as federated training of models on mobile devices [34], controlling and coordinating robot swarms [11], or distributed control and power grids [14], all rely on distributed computation.
+From a communication perspective, distributed frameworks can be client-server (i.e., federated) or decentralized (i.e., networked), with the main difference being that in the client-server setup users communicate with a central server, while in decentralized settings users communicate directly with each other.
+[^1] While the client-server setup is often characterized by performing local updates and periodic communication [65], as well as partial user participation [17], we note that these features can also be incorporated in the decentralized setup, by modifying the methods to perform multiple local updates before communicating [33] and considering a dynamic network with users that are occasionally idle [12].
 However, this is not the focus of our current work.
 Noting that, from the model update perspective, the client-server setup is equivalent to the decentralized setup with a fully connected communication network, we focus on the more general, decentralized setup.
-The study of convergence guarantees of decentralized optimization algorithms has a long history, e.g., [nedic-subgrad, jakovetic-fast, shi-extra, scutari-next, kun-dgd, jakovetic-unification, swenson-dsgd], with most works focusing on MSE convergence, e.g., [conv-rates-dsgd-jakovetic, Pu2021, vlaski-dsgd-nonconv-1, vlaski-dsgd-nonconv-2, pmlr-v119-koloskova20a, ran-vr, ran-improved, wang-cooperative], see also [adaptive-learning-ali] for an extensive treatment of the topic.
+The study of convergence guarantees of decentralized optimization algorithms has a long history, e.g., [53, 25, 64, 43, 81, 26, 66], with most works focusing on MSE convergence, e.g., [24, 57, 70, 69, 33, 74, 75, 72], see also [62] for an extensive treatment of the topic.
 Another type of convergence guarantees, namely convergence in HP, has garnered increasing attention recently.
-In particular, for a non-negative stochastic process $\{X^{t}\}_{t\in\mathbb{N}}$, the goal of HP convergence guarantees is to establish, for all $t\in\mathbb{N}$ and any $\epsilon>0$, that
+In particular, for a non-negative stochastic process $\{X^{t}\}_{t\in\mathbb{N}}$, the goal of HP convergence is to establish that, for all $t\in\mathbb{N}$ and any $\epsilon>0$
 $\mathbb{P}\big(X^{t}>\epsilon\big)\leq\exp\big(-Ct^{\gamma_{1}}\epsilon^{\gamma_{2}}\big),$ (1)
-where $\gamma_{1},\gamma_{2},C>0$ are some positive constants.
-If $\{X^{t}\}_{t\in\mathbb{N}}$ is a measure of performance of an algorithm, e.g., $X^{t}=\frac{1}{t}\sum_{k\in[t]}\|\nabla f(x^{k})\|^{2}$, where $\{x^{t}\}_{t\in\mathbb{N}}$ is a sequence generated by the algorithm and $f$ is a non-convex cost, the relation (1) provides strong guarantees with respect to a single run of the algorithm.
+where $\gamma_{1},\gamma_{2},C>0$ are constants.
+If $\{X^{t}\}_{t\in\mathbb{N}}$ is a measure of performance, e.g., $X^{t}=\frac{1}{t}\sum_{k\in[t]}\|\nabla f(x^{k})\|^{2}$, where $\{x^{t}\}_{t\in\mathbb{N}}$ is a sequence generated by some algorithm and $f$ is a non-convex cost, the relation (1) provides strong guarantees with respect to a single run of that algorithm.
 This is particularly important in modern applications like LLMs, where it is often intractable to perform multiple runs.
-Numerous works study HP guarantees of SGD-type methods in centralized settings, under both light-tailed [nemirovski2009robust, ghadimi2013stochastic, li-orabona, pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high] and heavy-tailed noise [nguyen2023improved, liu2023breaking, hubler2024normalization, kornilov2025sign, armacki2023high].
-Comparatively, there have been few studies on HP convergence of decentralized methods.
+Numerous works study HP guarantees of SGD-type methods in centralized settings, under both light-tailed [54, 20, 36, 21, 5, 39] and heavy-tailed noise [56, 40, 23, 35, 4].
+Comparatively, there have been very few studies of HP convergence of decentralized methods.
 1.1 Literature Review
-We now review the literature, focusing on works studying centralized HP convergence guarantees and decentralized MSE and HP convergence guarantees.
+We now review the literature, focusing on centralized HP results and decentralized MSE and HP convergence results.
 Centralized HP Convergence.
-The authors in [nemirovski2009robust] show optimal convergence rates of SGD for convex costs under light-tailed (i.e., sub-Gaussian) stochastic gradients (see assumption (A4) ahead for a formal definition of sub-Gaussianity), while the work [ghadimi2013stochastic] shows the same for non-convex costs.
-The authors in [li-orabona, li2020high] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively.
-The works [pmlr-v99-harvey19a] and [pmlr-v206-bajovic23a] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while in [liu2024revisiting] the authors establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs.
-The work [liu2023high] generalizes the previous studies for non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs.
-The authors in [madden2020high] study HP convergence of SGD under sub-Weibull noise.
-Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [heavy-tail-book, nguyen2023improved, armacki2023high] for various conditions used when studying heavy-tailed noise.
-e.g., [sadiev2023highprobability, nguyen2023improved, liu2023breaking, gorbunov2023breaking, hubler2024normalization, kornilov2025sign, armacki2024_ldp+mse, armacki2025optimal, armacki2023high], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1).
+[54] show optimal convergence rates of SGD for convex costs under light-tailed stochastic gradients (see assumption (A4) ahead for a formal definition of light-tailed noise), while [20] show the same for non-convex costs.
+[36, 37] establish high-probability guarantees of momentum SGD and a variant of AdaGrad for non-convex costs, respectively.
+[21] and [5] respectively provide optimal HP convergence rates for the last iterate of SGD for non-smooth and smooth strongly convex costs, while [41] establish unified HP convergence guarantees for smooth and non-smooth convex and strongly convex costs.
+[39] generalize the previous works on non-convex and convex costs, providing unified guarantees for several algorithms, including SGD and AdaGrad for smooth and non-smooth costs.
+[46] study HP convergence of SGD under sub-Weibull noise.
+Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [50, 56, 4] for various conditions used to study heavy-tailed noise.
+e.g., [61, 56, 40, 58, 23, 35, 3, 2, 4], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1).
 Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
@@
 MSE guarantees are typically studied under a variety of bounds on the second noise moment.
-[^3] See, e.g., [khaled2022better] for a good overview of the various conditions used in the literature.
-Following one of these settings, the work [conv-rates-dsgd-jakovetic] shows that DSGD converges at an optimal rate for strongly convex costs, while in [Pu2021] the authors show that DSGD with a fixed step-size and the gradient tracking (GT) mechanism, e.g., [scutari-next, nedic-tracking, qu-harnessing], converges to a neighbourhood of the optimal solution.
-The works [vlaski-dsgd-nonconv-1, vlaski-dsgd-nonconv-2, vlaski-dsgd-nonconv-3] study MSE guarantees of DSGD for non-convex costs and show that it escapes saddle points with high probability.
-In [wang-cooperative] the authors propose a general framework dubbed cooperative SGD, showing optimal rates for non-convex costs.
-The work [koloskova2023grad_clip] provides unified guarantees for DSGD with local updates and changing network topology, with optimal rates and linear speed-up in the number of users for non-convex and (strongly) convex costs.
-In [ran-vr] the authors show DSGD with GT and variance reduction converges at a linear rate for strongly convex costs, while in [ran-improved] optimal rates of DSGD with GT and linear speed-up in the number of users are established for non-convex costs and costs satisfying the Polyak-Łojasiewicz (PL) condition [karimi2016linear].
-Finally, in [yu2025decentralized] the authors show that DSGD with GT, local momentum and gradient normalization achieves optimal rates under heavy-tailed noise.
-It is also worth mentioning a rich line of works studying MSE guarantees in decentralized settings for problems such as estimation and detection, multi-objective and multitask optimization, see [consensus+innovation1, chen-decent-pareto, nassif-multitask] and references therein.
+[^3] See, e.g., [31] for a good overview of the various conditions used in the literature.
+Following one of these settings, [24] show that DSGD converges at an optimal rate for strongly convex costs, while [57] show that DSGD with a fixed step-size and the gradient tracking (GT) mechanism, e.g., [43, 52, 60], converges to a neighbourhood of the optimal solution.
+[70, 69, 71] study MSE guarantees of DSGD for non-convex costs and show that it escapes saddle points with high probability.
+[72] propose a general framework dubbed cooperative SGD, showing optimal rates for non-convex costs.
+[32] provide unified guarantees for DSGD with local updates and changing network topology, with optimal rates and linear speed-up in the number of users for non-convex and (strongly) convex costs.
+[74] show DSGD with GT and variance reduction converges at a linear rate for strongly convex costs, while [75] establish optimal rates of DSGD with GT and linear speed-up in the number of users for non-convex costs and costs satisfying the Polyak-Łojasiewicz (PL) condition, e.g., [30].
+[79] show that DSGD with GT, momentum and normalization achieves optimal rates under heavy-tailed noise.
+It is worth mentioning a rich line of works studying MSE guarantees for decentralized problems such as estimation, detection, multi-objective and multitask optimization, see, e.g., [28, 16, 51] and references therein.
 Decentralized HP Convergence.
 Compared to MSE guarantees, there is a significantly smaller body of work on HP convergence guarantees of decentralized algorithms.
-In particular, the work [dsgd-high-prob] studies HP convergence of DSGD for both non-convex and PŁ costs under light-tailed noise, requiring uniformly bounded gradients and asymptotically vanishing noise for PŁ costs (see discussion after Theorem 2 ahead), showing optimal rates in both cases.
-In [dsgd-online-noncooperative-games, dsgd-online-stochastic] the authors study HP convergence of a decentralized mirror descent algorithm under sub-Gaussian noise and bounded gradients, for online noncooperative games and dynamic regrets, respectively.
-Works [clipped-dual-avg, clipped-dsgd1, clipped-dsgd2] study HP convergence of decentralized algorithms with clipping under heavy-tailed noise.
-It is also worth mentioning [bajovic-detection-ld, bajovic-detection-ld2, matta-diffusion-1, matta-difussion-2, bajovic-inference-ld], where the authors study asymptotic large deviation guarantees for decentralized problems like detection and inference.
-A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradient stays bounded.
-This observation raises the question of whether decentralized algorithms can converge in the HP sense without ensuring that the gradient is uniformly bounded, creating a gap between convergence in the HP and MSE sense, where bounded gradients are not required.
-This is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence.
-Moreover, none of the existing HP convergence bounds in decentralized settings achieve linear speed-up in the number of users, which is a known feature of decentralized algorithms, established in the MSE sense in, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined].
+In particular, [44] study HP convergence of DSGD for general non-convex and PŁ costs under light-tailed noise, requiring uniformly bounded gradients, as well as asymptotically vanishing noise for PŁ costs (see the discussion after Theorem 2 for details), while showing optimal rates in both cases.
+[45, 76] study HP convergence of a decentralized mirror descent algorithm under light-tailed noise, for online noncooperative games and dynamic regrets, respectively, requiring bounded gradients and compact domains.
+[59, 77] study HP convergence of decentralized algorithms with clipping, for convex costs under heavy-tailed noise, while [78] also consider general non-convex costs.
+It is worth mentioning works like [7, 6, 47, 48, 8], where the authors study asymptotic large deviation guarantees for decentralized problems such as detection and inference.
+A common thread for all these works is the need for uniformly bounded gradients, or algorithmic modifications like gradient clipping, which ensure that the gradients stay bounded, as well as the fact that none of the existing HP studies achieve linear speed-up in the number of users, which is a known feature of decentralized algorithms in the MSE sense in, e.g., [33, 75, 1].
+These observations raise the following important questions: (i) Can decentralized algorithms converge in the HP sense under the same assumptions on the cost as used for MSE results, where bounded gradients are not required?
+(ii) Is it possible to achieve a linear speed-up in the number of users in the HP sense, which is a well-known feature of decentralized algorithms?
+This situation is further contrasted by the centralized setting, where it is known that SGD-type algorithms require the same conditions on the cost function to guarantee both HP and MSE convergence.
 1.2 Contributions
-Motivated by the observed gap between the existing literature on HP and MSE convergence in decentralized settings, we revisit the HP convergence guarantees in decentralized stochastic optimization, by studying a variant of vanilla DSGD in the presence of light-tailed noise.
-We establish that DSGD converges in HP under the same conditions on the cost as required in the MSE sense, for both non-convex and strongly convex costs, removing the uniformly bounded gradient condition.
+Motivated by the observed gaps between the existing literature on HP and MSE convergence in decentralized settings, we revisit the HP convergence guarantees in decentralized optimization, by studying a variant of vanilla DSGD in the presence of light-tailed noise, establishing several strong results in the process.
+In particular, we show that DSGD converges in HP under the same conditions on the cost as required in the MSE sense, for both non-convex and strongly convex costs, removing strong conditions like uniformly bounded gradients and (asymptotically) vanishing noise.
 Next, we show that DSGD achieves optimal rates and linear speed-up in the number of users, for both non-convex and strongly convex costs, further improving on existing works and closing the gap between HP and MSE guarantees in decentralized settings.
 Our results are established by carefully bounding the MGF of the quantity of interest (i.e., average norm-squared of the gradient for non-convex and optimality gap of the last iterate for strongly convex costs) and the MGF of the consensus gap.
-Compared to [dsgd-high-prob], the work closest to ours, we provide several improvements.
+Compared to [44], the work closest to ours, we provide several improvements.
 In particular, we remove the uniformly bounded gradient requirement, as well as the vanishing noise requirement imposed for PŁ costs (see the discussion after Theorem 2 ahead), while achieving linear speed-up in the number of users.
-Compared to works studying MSE guarantees for DSGD, e.g., [pmlr-v119-koloskova20a, conv-rates-dsgd-jakovetic], our results require directly working with the MGF and carefully balancing between the MGF of the quantity of interest and that of the consensus gap.
+Compared to works studying MSE guarantees for DSGD, e.g., [33, 24], our results require directly working with the MGF and carefully balancing between the MGF of the quantity of interest and that of the consensus gap.
 This is particularly challenging for strongly convex costs, where we want to show an “almost decreasing” property of the MGF, in order to get improved rates on the last iterate (see Lemma 5 and the related discussion ahead).
-Compared to centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], the main challenge lies in the additional consensus gap and simultaneously dealing with its MGF and that of the quantity of interest.
-All of these challenges are resolved by introducing several novelties, outlined next.
+Compared to centralized SGD, e.g., [21, 5, 39, 41], the main challenge lies in the additional consensus gap and simultaneously dealing with its MGF and that of the quantity of interest.
+These challenges are resolved by introducing several novelties, outlined next.
 Novelty.
 Toward establishing our improved HP convergence guarantees, we face several technical challenges, requiring novel results.
-In particular, to remove uniformly bounded gradients, we provide Lemma 3 and use the “offset trick” for non-convex costs (see the proof sketch of Theorem 1 in Section 4 for details), while for strongly convex costs we establish a novel bound on the MGF of the consensus gap, in the form of Lemma 4.
+In particular, to remove uniformly bounded gradients, we provide Lemma 3 and use the “offset trick” for non-convex costs (see the proof sketch of Theorem 1 in Section 4 for details), while for strongly convex costs we establish a novel bound on the MGF of the consensus gap, removing the need for bounded gradients and bounded heterogeneity, in the form of Lemma 4.
 To establish linear speed-up in the number of users, we show that the variance reduction benefit of decentralized learning is maintained in the HP sense, in the form of Lemma 2, which is of independent interest when studying HP guarantees of decentralized algorithms.
-Further, to guarantee linear speed-up for strongly convex costs, we require a more fine-grained bound on the MGF, provided in the form of Lemma 5, which allows for incorporating higher-order terms, generalizing the bounds from [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting] and is of independent interest, even in centralized settings.
+Further, we establish a novel result on the MGF of the optimality gap for strongly convex costs (more broadly, “almost decreasing” processes, see Section 3.3 ahead), in the form of Lemma 5, which is essential in ensuring linear speed-up, by providing a more fine-grained bound on the MGF compared to existing centralized results [21, 5, 41] and is of independent interest, even in centralized settings.
 Paper Organization.
 The rest of the paper is organized as follows.
-Section 2 formally introduces the problem and the DSGD method, Section 3 outlines the main results, Section 4 provides proof sketches and discussion, while Section 5 concludes the paper.
+Section 2 outlines the problem and DSGD method, Section 3 presents the main results, Section 4 provides proof sketches and discussions, while Section 5 concludes the paper.
 Appendix contains results omitted from the main body.
@@
 For $m\in\mathbb{N}$, we use $m]=\{1,\ldots,m\}$ to denote positive integers up to and including $m$.
-The notation $\langle\cdot,\cdot\rangle$ stands for the Euclidean inner product, while $\|\cdot\|$ is used for both the induced vector and matrix $\ell_{2}$ norms.
+The notation $\langle\cdot,\cdot\rangle$ stands for the Euclidean inner product, while $\|\cdot\|$ is used for both vector and matrix induced norms.
 We use subscripts to denote users and superscripts to denote the iteration counter, e.g., $x^{t}_{i}$ refers to the model of user $i$ in iteration $t$.
-We use the “big O” notation $\mathcal{O}(\cdot)$ to hide global constants, unless stated otherwise.
-2 Problem Setup and Proposed Method
-In this section we introduce the problem of decentralized stochastic optimization and the DSGD algorithm.
-Consider a network of $n\geq 2$ users which can communicate with each other and want to jointly train a model.
+The “big O” notation $\mathcal{O}(\cdot)$ hides only global constants, unless stated otherwise.
+2 Problem Setup and DSGD
+In this section we introduce the problem of interest and the DSGD algorithm.
+Consider a network of $n\geq 2$ users who can communicate with each other and want to jointly train a model.
 Formally, the problem can be cast as
-$\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}\bigg\{f(x)=\frac{1}{n}\sum_{i\in[n]}f_{i}(x)\bigg\},$ (2)
-where $x\in\mathbb{R}^{d}$ represents model parameters, $f_{i}:\mathbb{R}^{d}\mapsto\mathbb{R}$ is the cost function associated to user $i\in[n]$, given by $f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[\ell(x;\xi_{i})]$, with $\xi_{i}\in\Xi$ being a random variable governed by some unknown distribution $\mathcal{D}_{i}$, while $\ell:\mathbb{R}^{d}\times\Xi\mapsto\mathbb{R}$ is a loss function.
-Each user has access to a Stochastic First-order Oracle ($\mathcal{SFO}$), which, when queried by user $i\in[n]$ and input $x\in\mathbb{R}^{d}$, returns the gradient of $\ell$ evaluated at a random sample $\xi_{i}\sim\mathcal{D}_{i}$, i.e., $\nabla\ell(x;\xi_{i})$.
+$\operatorname*{arg\,min}_{x\in\mathbb{R}^{d}}\Big\{f(x)=\frac{1}{n}\sum_{i\in[n]}f_{i}(x)\Big\},$ (2)
+where $x\in\mathbb{R}^{d}$ represents model parameters, $f_{i}:\mathbb{R}^{d}\mapsto\mathbb{R}$ is the cost function of user $i\in[n]$, given by $f_{i}(x)=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[\ell(x;\xi_{i})]$, with $\xi_{i}\in\Xi$ being a random variable governed by an unknown distribution $\mathcal{D}_{i}$, while $\ell:\mathbb{R}^{d}\times\Xi\mapsto\mathbb{R}$ is a loss function.
+Each user has access to a Stochastic First-order Oracle ($\mathcal{SFO}$), which, when queried by user $i\in[n]$ with input $x\in\mathbb{R}^{d}$, returns the gradient of $\ell$ evaluated at a random sample, i.e., $\nabla\ell(x;\xi_{i})$.
 The $\mathcal{SFO}$ model subsumes:
-  1. 1.
-    Batch (offline) learning, where users have access to a local dataset $\{\xi_{i,l}\}_{l\in[m_{i}]}$, so that $f_{i}(x)=\frac{1}{m_{i}}\sum_{l\in[m_{i}]}\ell(x;\xi_{i,l})$ and in each round users choose a sample $\xi_{i,l}$ uniformly at random, [^4] Note that it is possible to have non-uniform weights, i.e., $f_{i}(x)=\sum_{l\in[m_{i}]}p_{i,l}\ell(x;\xi_{i,l})$, where $p_{i,l}>0$ for all $l\in[m_{i}]$ and $\sum_{l\in[m_{i}]}p_{i,l}=1$, resulting in non-uniform sampling, with sample $\xi_{i,l}$ selected with probability $p_{i,l}$.
-which is used to compute $\nabla\ell(x;\xi_{i,l})$ and update model parameters;
-  2. 2.
-    Streaming (i.e., online) learning, where users do not store a local dataset, but in each round observe a random sample $\xi_{i}\sim\mathcal{D}_{i}$ from a data stream, which is used to compute $\nabla\ell(x;\xi_{i})$ and update model parameters.
-The communication pattern between users is modeled as a static graph $G=(V,E)$, where $V=[n]$ is the set of vertices representing users, while $E\subset V\times V$ is the set of edges representing communication links between users.
-[^5] Our work can be readily extended to non-static, time-varying graphs, see the Appendix for a further discussion.
-To solve (2) in decentralized fashion, we consider a version of Decentralized Stochastic Gradient Descent (DSGD), based on the Adapt-Then-Combine (i.e., diffusion) approach, e.g., [diffusion1, diffusion2, diffusion3].
-The method consists of the following steps.
-At the start, users choose a shared step-size schedule $\{\alpha_{t}\}_{t\in\mathbb{N}}$ and each user $i\in[n]$ chooses an arbitrary, but deterministically selected initial model $x^{1}_{i}\in\mathbb{R}^{d}$.
-[^6] While the initial models can be any real vectors, possibly different across users, they needs to be deterministic quantities, for the sake of theoretical analysis.
-In iteration $t\geq 1$, users query the $\mathcal{SFO}$ with their current model $x^{t}_{i}$ and receive $\nabla\ell(x^{t}_{i};\xi_{i}^{t})$.
-Users first update their local model via the rule $x_{i}^{t+\nicefrac{{1}}{{2}}}=x^{t}_{i}-\alpha_{t}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$, after which the new model is produced by performing a consensus step, i.e., $x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}x^{t+\nicefrac{{1}}{{2}}}_{j}$, where $\mathcal{N}_{i}\coloneqq\{j\in V:\{i,j\}\in E\}\cup\{i\}$ is the set of users (i.e., neighbours) with whom user $i$ can communicate (including user $i$ itself), while $w_{ij}>0$ is the weight user $i$ assigns to user $j$’s model.
-The update rule at user $i\in[n]$ can be compactly represented as
-$x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\big(x_{j}^{t}-\alpha_{t}\nabla\ell(x_{j}^{t};\xi_{j}^{t})\big).$ (3)
-The method is summarized in Algorithm 1.
-The version of DSGD considered in our work is related to the version based on Combine-Then-Adapt (i.e., consensus ++ innovation) approach, e.g., [nedic-subgrad, consensus+innovation1, consensus+innovation2], whose HP convergence is studied in [dsgd-high-prob].
+1. Batch (offline) learning
+- users have access to a local dataset $\{\xi_{i,l}\}_{l\in[m_{i}]}$, so that $f_{i}(x)=\frac{1}{m_{i}}\sum_{l\in[m_{i}]}\ell(x;\xi_{i,l})$ and in each round users choose a random sample $\xi_{i,l}$, which is used to compute $\nabla\ell(x;\xi_{i,l})$ and update the model;
+2. Streaming (online) learning
+- users do not store a local dataset, but in each round observe a random sample $\xi_{i}$, which is used to compute $\nabla\ell(x;\xi_{i})$ and update the model.
 Algorithm 1 DSGD
@@
 4:end for
+The communication pattern between users is modeled as a static graph $G=(V,E)$, where $V=[n]$ is the set of vertices representing users, while $E\subset V\times V$ is the set of edges representing communication links between users.
+To solve (2) in decentralized fashion, we consider a version of Decentralized Stochastic Gradient Descent (DSGD), based on the Adapt-Then-Combine (i.e., diffusion) approach, e.g., [42, 13, 15].
+The method consists of the following steps.
+First, users choose a shared step-size schedule $\{\alpha_{t}\}_{t\in\mathbb{N}}$ and each user $i\in[n]$ chooses an arbitrary, but deterministically selected initial model $x^{1}_{i}\in\mathbb{R}^{d}$.
+[^4] While initial models can be any real vectors, possibly different across users, for analysis purposes they need to be deterministic.
+In iteration $t\geq 1$, users query the $\mathcal{SFO}$ with their current model $x^{t}_{i}$ and receive $\nabla\ell(x^{t}_{i};\xi_{i}^{t})$.
+Users then update their models via the rule
+$x^{t+1}_{i}=\sum_{j\in\mathcal{N}_{i}}w_{ij}\big(x_{j}^{t}-\alpha_{t}\nabla\ell(x_{j}^{t};\xi_{j}^{t})\big),$ (3)
+which consists of a local update, followed by a consensus step, where $\mathcal{N}_{i}\coloneqq\{j\in V:\{i,j\}\in E\}\cup\{i\}$ is the set of users (i.e., neighbours) with whom user $i$ can communicate (including $i$ itself), while $w_{ij}>0$ is the weight user $i$ assigns to user $j$’s model.
+The method is summarized in Algorithm 1.
+The version of DSGD considered in our work is closely related to the variant based on Combine-Then-Adapt (i.e., consensus ++ innovation) approach, e.g., [53, 28, 29], whose HP convergence is studied by [44].
 3 Main Results
-In this section we present the main results.
-Subsection 3.1 states the preliminaries, Subsection 3.2 provides results for non-convex costs, while Subsection 3.3 presents results for strongly convex costs.
+In this section we present the main results. Subsection 3.1 states the preliminaries, while Subsections 3.2 and 3.3 provide results for non-convex and strongly convex costs.
 3.1 Preliminaries
@@
 For ease of notation, let $z^{t}_{i}\coloneqq\nabla\ell(x^{t}_{i};\xi_{i}^{t})-\nabla f_{i}(x^{t}_{i})$ and $W\in\mathbb{R}^{n\times n}$, where $W]_{i,j}\coloneqq w_{ij}$, denote the stochastic noise and the network communication matrix, respectively.
-(A1) The network communication matrix $W\in\mathbb{R}^{n\times n}$ is primitive and doubly stochastic.
-Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [ran-vr].
+The network communication matrix $W\in\mathbb{R}^{n\times n}$ is primitive and doubly stochastic.
+Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [74].
+[^5] Our work can be readily extended to the setting of time-varying networks, see Appendix J for a detailed discussion.
 Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix.
-It can be shown that (A1) implies $\lambda\in[0,1)$, see [Horn_Johnson_2012].
-(A2) The global cost $f$ is bounded from below, i.e., $f^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}f(x)>$-$\infty$.
-(A3) Each local cost $f_{i}$ has $L$-Lipschitz gradients, i.e., $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, for any $x,y\in\mathbb{R}^{d}$.
-Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [ghadimi2013stochastic, pmlr-v119-koloskova20a, ran-improved].
-(A4) The stochastic quantities satisfy the following:
+It can be shown that (A1) implies $\lambda\in[0,1)$, see, e.g., [22].
+The global cost $f$ is bounded from below, i.e., $f^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}f(x)>$-$\infty$.
+Each local cost $f_{i}$ has $L$-Lipschitz gradients, i.e., $\|\nabla f_{i}(x)-\nabla f_{i}(y)\|\leq L\|x-y\|$, for any $x,y\in\mathbb{R}^{d}$.
+Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [20, 75].
+The stochastic quantities satisfy the following:
   1. 1.
@@
   2. 2.
-    The stochastic gradients are unbiased, i.e., for any $i\in[n]$, $t\geq 1$ and $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
-    $\mathbb{E}[\nabla\ell(x;\xi_{i}^{t})\>|\>\mathcal{F}_{t}]=\nabla f_{i}(x).$
+    The stochastic gradients are unbiased, i.e., for any $i\in[n]$, $t\geq 1$ and $\mathcal{F}_{t}$-measurable vector $x\in\mathbb{R}^{d}$, we have $\mathbb{E}[\nabla\ell(x;\xi_{i}^{t})\>|\>\mathcal{F}_{t}]=\nabla f_{i}(x)$.
   3. 3.
-    The noise at each user $i\in[n]$ is $\sigma_{i}$-sub-Gaussian, i.e., for all $t\geq 1$ and any $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
+    The noise at user $i\in[n]$ is $\sigma_{i}$-sub-Gaussian (light-tailed), i.e., for all $t\geq 1$ and $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
     $\mathbb{E}\left[\exp\bigg(\frac{\|z_{i}^{t}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1).$
-The first condition in (A4) is standard in decentralized stochastic optimization [ran-improved, yu2025decentralized], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian).
-Light tails are necessary to achieve exponentially decaying tail bounds of the form (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [nemirovski2009robust, ghadimi2013stochastic, li-orabona, li2020high, pmlr-v99-harvey19a, liu2023high, pmlr-v206-bajovic23a].
-[^7] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [madden2020high] and references therein, the focus of this work is on the more widely studied sub-Gaussian noise setting.
-(A5) The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, [^8] With at least one of $A$ or $B$ being strictly positive.
-we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
-A heterogeneity bound of the type (A5) is in fact required to ensure that DSGD converges for non-convex costs [NIPS2017_f7552665, dist-learn-noncvx, pmlr-v119-koloskova20a, wang-cooperative].
-Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [dsgd-high-prob], as it allows users’ gradients to grow with the global gradient.
-Compared to [pmlr-v119-koloskova20a], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition.
-While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [pmlr-v119-koloskova20a], to keep the proofs simple and instructive, we will use assumption (A5).
-(A6) Each $f_{i}$ is twice continuously differentiable and $\mu$-strongly convex, i.e., for every $x,y\in\mathbb{R}^{d}$, we have $f_{i}(x)\geq f_{i}(y)+\langle\nabla f_{i}(y),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}$.
-Assumption (A6) is used in the analysis of strongly convex costs.
-It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a, wang-cooperative] which is in general not required for centralized SGD.
-The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [ran-improved], however, this is beyond the scope of the current work.
-Next, let $\overline{x}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}x^{t}_{i}$ denote the network-wise average model.
-Note that $\overline{x}^{t}$ is an ideal model, used for analysis purposes.
-Using (3) and the fact that $W$ is doubly stochastic, it can be readily seen that
-$\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t},$
-where $\overline{g}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$ is the average stochastic gradient.
-Let $\overline{z}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}z_{i}^{t}$ denote the average noise.
+The first condition in (A4) is standard in decentralized stochastic optimization [75, 79], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian).
+Light tails are necessary to achieve exponentially decaying tail bounds of the form in (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [54, 20, 36, 37, 21, 39, 5].
+[^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [46] and references therein, the focus of this work is on sub-Gaussian noise.
+The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
+A heterogeneity bound of the type in (A5) is in fact required to ensure that DSGD converges for non-convex costs [38, 14, 33, 72].
+Note that (A5) is strictly weaker than the uniformly bounded gradient assumption used in [44], as it allows users’ gradients to grow with the global gradient.
+Compared to [33], who analyze the MSE convergence and require a bound on the average heterogeneity, i.e., $\frac{1}{n}\sum_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$, we impose a slightly stronger condition.
+While we believe that (A5) can be relaxed to average heterogeneity by incorporating a similar analysis technique to the one in [33], to keep the proofs simple and instructive, we will use assumption (A5).
+Each $f_{i}$ is twice continuously differentiable and $\mu$-strongly convex, i.e., for every $x,y\in\mathbb{R}^{d}$, we have $f_{i}(x)\geq f_{i}(y)+\langle\nabla f_{i}(y),x-y\rangle+\frac{\mu}{2}\|x-y\|^{2}$.
+Assumption (A6) is used for the strongly convex case.
+It is well known that DSGD requires (A3) and (A6) to hold for each $f_{i}$, e.g., [24, 33, 72].
+The heterogeneity bound in (A5) and strong convexity of each cost in (A6) can be removed by deploying the GT technique, e.g., [75], however, this is beyond the scope of the current work.
+Next, let $\overline{x}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}x^{t}_{i}$, $\overline{g}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}\nabla\ell(x^{t}_{i};\xi_{i}^{t})$ and $\overline{z}^{t}\coloneqq\frac{1}{n}\sum_{i\in[n]}z_{i}^{t}$ denote the network-average model, stochastic gradient and noise, respectively.
+Using (3) and the double stochasticity of $W$, it follows that
+$\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t}.$
 We then have the following result.
@@
 The right-hand side of the above inequality consists of terms that arise in centralized SGD plus a consensus gap term $\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}$, which stems from the decentralized nature of the algorithm and bounding this term is crucial to ensure convergence.
-To finish this subsection, we provide a technical result on the behaviour of the stochastic noise in the HP sense.
+To close the subsection, we provide an important result on the behaviour of the stochastic noise in the HP sense.
 Let $\sigma^{2}\coloneqq\frac{1}{n}\sum_{i\in[n]}\sigma^{2}_{i}$ be the average noise parameter.
@@
     The average noise $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian, i.e., $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{120d\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1)$.
-Lemma 2 establishes some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, highlighting the variance reduction benefit of decentralized optimization in the HP sense.
+Lemma 2 provides some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, establishing the variance reduction benefit of decentralized optimization in the HP sense.
 This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections.
-We note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and the Appendix.
+Note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and Appendix K.
 3.2 Non-Convex Costs
 In this subsection we present results for non-convex costs.
-Let $\Delta_{x}\coloneqq\frac{1}{n}\sum_{i\in[n]}\|x^{1}_{i}-\overline{x}^{1}\|^{2}$ denote the initial consensus gap.
-We then have the following result.
+Let $\Delta_{x}\coloneqq\frac{1}{n}\sum_{i\in[n]}\|x^{1}_{i}-\overline{x}^{1}\|^{2}$ and $\Delta_{f}\coloneqq f(\overline{x}^{1})-f^{\star}$ denote the initial consensus and optimality gap.
 Lemma 3.
-Let (A1) and (A5) hold. We then have, for any $t\geq 1$
-$\frac{1}{n}\sum_{i\in[n]}$ $\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\leq 2\lambda^{2t}\Delta_{x}+\frac{4\lambda^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\|z^{k}_{i}\|^{2}$
-$+\frac{4\lambda^{2}A^{2}}{1-\lambda}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}+\frac{4\lambda^{2}B^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\|\nabla f(x^{k}_{i})\|^{2}.$
-Lemma 3 provides a deterministic bound on the consensus gap and is used to bound the resulting MGF.
-Define $\Delta_{f}\coloneqq f(\overline{x}^{1})-f^{\star}$ and $\overline{\sigma}\coloneqq\max_{i\in[n]}\sigma_{i}$.
-Building on Lemmas 1 and 3, we get the following result.
+If (A1) and (A5) hold, then for any $t\geq 1$
+$\frac{1}{n}\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\leq 2\lambda^{2t}\Delta_{x}+\frac{4\lambda^{2}A^{2}}{1-\lambda}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}+\frac{4\lambda^{2}}{n(1-\lambda)}\sum_{k\in[t]}\alpha_{k}^{2}\lambda^{t-k}\sum_{i\in[n]}\Big[\|z^{k}_{i}\|^{2}+B^{2}\|\nabla f(x^{k}_{i})\|^{2}\Big].$
+Lemma 3 provides a useful deterministic bound on the consensus gap. Define $\overline{\sigma}\coloneqq\max_{i\in[n]}\sigma_{i}$. Building on Lemmas 1 and 3, we get the following result.
 Theorem 1.
@@
 If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
-$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
-we then have, for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}$
-$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
-Theorem 1 establishes HP convergence guarantees for DSGD with general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few perspectives.
+$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$
+then for any $\delta\in(0,1)$, with probability at least $1-\delta$
+$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
+Theorem 1 establishes HP convergence of DSGD for general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few different perspectives.
 Centralized SGD and linear speed-up.
-Compared to centralized SGD, which achieves the rate $\mathcal{O}\Big(\frac{1}{\sqrt{T}}+\frac{1}{T}\Big)$, e.g., [liu2023high], we can see that the rate from Theorem 1 similarly consists of two terms of different orders, namely $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}+\frac{n}{T}\Big)$.
-Crucially, the leading term achieves a linear speed-up in the number of users, which is consistent with MSE results, e.g., [pmlr-v119-koloskova20a], and shows that DSGD maintains the benefits of decentralized learning, in the HP sense.
+Compared to centralized SGD, which achieves the rate $\mathcal{O}\Big(\frac{1}{\sqrt{T}}+\frac{1}{T}\Big)$, e.g., [39], we can see that the rate in Theorem 1 also consists of two different order terms, namely $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}+\frac{n}{T}\Big)$.
+Crucially, the leading term achieves a linear speed-up in the number of users, which is consistent with MSE results, e.g., [33], showing that DSGD retains the benefits of decentralized learning in the HP sense.
 Network effect and transient time.
-We can see from the bound in Theorem 1 that network connectivity only affects the higher-order terms decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is again consistent with MSE results, e.g., [pmlr-v119-koloskova20a].
-Moreover, we can use the rate in Theorem 1 to establish the transient time of DSGD in the HP sense, which can be shown to be of the order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD stemming from MSE results, see, e.g., [unified-refined, Table I].
-For details on the derivation of the transient time, see the Appendix.
+We can see from the bound in Theorem 1 that network connectivity only affects higher-order terms, decaying at rate $\mathcal{O}\Big(\frac{1}{T}\Big)$, which is consistent with MSE results, e.g., [33].
+Moreover, the rate in Theorem 1 implies that the transient time of DSGD in the HP sense is of order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD implied by MSE rates, see, e.g., Table I in [1].
+For details on the derivation of the transient time, see Appendix H.
 Effects of noise and heterogeneity.
@@
 Note that, if $A=0$, the bound in Theorem 1 recovers the convergence rate of noiseless gradient descent, in the sense that, when $\sigma=A=0$, the above bound becomes $\mathcal{O}\big(\frac{1}{T}\big)$.
-Comparison with [dsgd-high-prob].
-The authors in [dsgd-high-prob] study HP guarantees of DSGD over undirected time-varying networks, establishing a rate $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$, using a time-varying step-size schedule $\alpha_{t}=\frac{C}{\sqrt{t+1}}$, for some $C>0$ and assuming uniformly bounded gradients.
-As we discuss next, our results provide several important improvements.
-First, we remove the uniformly bounded gradient assumption and replace it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details.
-Next, the authors in [dsgd-high-prob] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise.
-In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, as all terms converge at the same $\mathcal{O}\Big(\frac{1}{\sqrt{T}}\Big)$ rate.
-However, it would be impossible to obtain a linear speed-up in [dsgd-high-prob] even under a fixed step-size, as the authors show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity.
-On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used (see the Appendix for a detailed discussion on the results needed to achieve linear speed-up).
-[^10] This is also consistent with existing MSE guarantees, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined], where linear speed-up for non-convex costs is achieved using a fixed step-size.
-Finally, while we consider static networks, our results can be readily extended to the time-varying setting considered in [dsgd-high-prob], see the Appendix for details.
+Comparison with [44].
+The authors in [44] study HP guarantees of DSGD over undirected time-varying networks, under a time-varying step-size schedule $\alpha_{t}=\frac{a}{\sqrt{t+1}}$, assuming uniformly bounded gradients and establish the rate $\mathcal{O}\big(\frac{\log(\nicefrac{{T}}{{\delta}})}{\sqrt{T}}\big)$.
+As highlighted next, we provide several important improvements.
+First, we remove the uniformly bounded gradient assumption, replacing it with the more general heterogeneity condition (A5), which is facilitated by the use of the “offset trick”, see the discussion in Section 4 for details.
+Next, [44] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise.
+In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, see Appendix E for a detailed discussion.
+However, it would be impossible for [44] to obtain a linear speed-up even under a fixed step-size, as they show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity.
+On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used, see Appendix I for detailed discussion on the results needed to achieve linear speed-up.
+[^8] This is consistent with MSE results, e.g., [33, 75, 1], where linear speed-up in non-convex case is achieved only via a fixed step-size.
+Finally, while we consider static networks, our results can be extended to the time-varying setting in [44], see Appendix J for details.
 3.3 Strongly Convex Costs
-In this subsection we provide improved results for the last iterate of strongly convex costs.
-It is well known that strong convexity implies a unique global minimizer of $f$, see, e.g., [nesterov-lectures_on_cvxopt].
-Denote the global minimizer by $x^{\star}\in\mathbb{R}^{d}$, noting that $f^{\star}\coloneqq f(x^{\star})$ and let $\kappa\coloneqq\frac{L}{\mu}\geq 1$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively.
-Next, recall that DSGD does not require a heterogeneity bound to converge in the MSE sense for strongly convex costs, e.g., [conv-rates-dsgd-jakovetic, pmlr-v119-koloskova20a].
-To show similar benefits in the HP sense, a different approach to Lemma 3 is required.
-This is achieved by carefully bounding the MGF of the consensus gap and leveraging properties of strong convexity.
-The formal result is stated next.
+In this subsection we provide results for the last iterate of strongly convex costs.
+Recall that strong convexity implies a unique global minimizer, e.g., [55], denoted by $x^{\star}\in\mathbb{R}^{d}$, with $f^{\star}\coloneqq f(x^{\star})$, and let $\kappa\coloneqq\frac{L}{\mu}$ and $\|\nabla\mathbf{f}^{\star}\|^{2}\coloneqq\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ be the condition number and heterogeneity measure, respectively.
+It is known that DSGD does not require a heterogeneity bound of the form in (A5) for strongly convex costs in the MSE sense, e.g., [24, 33].
+To show this is also the case in the HP sense, we next provide a refined version of Lemma 3, which carefully leverages properties of strong convexity to bound the MGF of the consensus gap.
 Lemma 4.
 Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$.
-If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$
+If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$, we have
 $\mathbb{E}\Big[\exp\Big(\nu K_{t+1}\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\Big)\Big]\leq\exp\bigg(\nu K_{t+1}\Big(\sum_{k\in[t]}\lambda^{t-k}S_{k}+\sum_{k\in[t]}\lambda^{t-k}D_{k}\Big)\bigg),$
-where $S_{k}\coloneqq\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}\coloneqq\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{5}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{6}}\Big)$.
-Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5).
-Using Lemma 6 from the Appendix, it can be shown that the RHS is of the order $e^{\mathcal{O}(\nu K_{t+1}\alpha_{t}^{2})}$.
-We note that the requirement of same model initialization across users can be removed, at the cost of an additional term in the exponent on the RHS, of order $\mathcal{O}(\nu K_{t+1}\lambda^{t}\Delta_{x})$, which decays geometrically (recall that $\lambda\in[0,1)$).
-However, we assume same model initialization, for ease of exposition.
-Prior to stating the main theorem, we provide an important technical result, which allows us to establish sharper bounds and ensure linear speed-up in the number of users is achieved.
+where $S_{k},D_{k}>0$ are problem related constants.
+Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5), with $S_{k},D_{k}$ depending on the step-size and other problem parameters, see Appendix F for explicit forms of $S_{k},D_{k}$.
+We note that the requirement of same model initialization across users can be removed, at the cost of an additional, geometrically decaying term in the exponent on the RHS, however, we use same initialization, for ease of exposition.
+Prior to stating the main theorem, we provide an important technical result, which facilitates sharper bounds and ensures linear speed-up in the number of users.
 Lemma 5.
-Let $\{X^{t}\}_{t\geq 2}$ be a sequence of random variables initialized by a deterministic $X^{1}>0$, such that, for some $M\in\mathbb{N}$, $a,t_{0},C_{i}>0$, $i\in[M]$ and every $t\geq 1$
-$\mathbb{E}[\exp(X^{t+1})]\leq\mathbb{E}\bigg[\hskip-1.99997pt\exp\hskip-1.99997pt\bigg(\hskip-3.00003pt\Big(1-\frac{a}{t+t_{0}}\Big)X^{t}+\hskip-3.99994pt\sum_{i\in[M]}\hskip-1.99997pt\frac{C_{i}}{(t+t_{0})^{i}}\hskip-1.99997pt\bigg)\hskip-1.99997pt\bigg].$
-If $a\in(1,2]$ and $t_{0}\geq a$, we then have
+Let $\{X^{t}\}_{t\in\mathbb{N}}$ be a sequence of random variables initialized by a deterministic $X^{1}>0$, such that, for some $M\in\mathbb{N}$, $a,t_{0}>0$ and every $t\geq 1$
+$\mathbb{E}[\exp(X^{t+1})]\leq\mathbb{E}\bigg[\hskip-1.99997pt\exp\hskip-1.99997pt\bigg(\hskip-3.00003pt\Big(1-\frac{a}{t+t_{0}}\Big)X^{t}+\hskip-3.99994pt\sum_{i\in[M]}\hskip-1.99997pt\frac{C_{i}}{(t+t_{0})^{i}}\hskip-1.99997pt\bigg)\hskip-1.99997pt\bigg],$
+where $C_{i}>0$, $i\in[M]$. If $a\in(1,2]$ and $t_{0}\geq a$, we have
 $\mathbb{E}[\exp(X^{t+1})]\leq\exp\bigg(\frac{(t_{0}+1)^{a}X_{1}}{(t+1+t_{0})^{a}}+\frac{2^{a}C_{1}}{a}+\frac{2^{a}C_{2}/(a-1)}{t+1+t_{0}}\bigg)$
 $\times\exp\bigg(\frac{2^{a}C_{3}\log(t+1+t_{0})}{(t+1+t_{0})^{a}}+\sum_{j=4}^{M}\frac{2^{a}t_{0}^{3-j}C_{j}}{(j-3)(t+1+t_{0})^{a}}\bigg).$
-Lemma 5 provides a bound on the MGF of an “almost decreasing” process.
-Compared to bounds used in the centralized setting, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], Lemma 5 gives a more fine-grained result, allowing us to incorporate higher-order terms in the final rate, which is crucial for establishing linear speed-up, see Section 4 and the Appendix for detailed discussions.
+Lemma 5 provides a tight bound on the MGF of an “almost decreasing” process, with the optimality gap of strongly convex costs being an instance of such a process.
+Compared to bounds used in the centralized setting, e.g., [21, 5, 41], Lemma 5 is significantly sharper and incorporates higher-order terms, which are crucial for ensuring linear speed-up, see Section 4 and Appendix I for detailed discussions.
 We are now ready to state the main result for strongly convex costs.
 Theorem 2.
-Let (A1)-(A4) and (A6) hold, let the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$.
-If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
+Let (A1)-(A4) and (A6) hold, let the step-size schedule be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$.
+If constants $a$ and $t_{0}$ are chosen as $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
 $\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+d\sigma^{2}\kappa/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(T+t_{0})^{2}}\bigg),$
 where $\mathcal{O}(\cdot)$ hides some higher-order terms.
-Theorem 2 establishes HP convergence guarantees for smooth strongly convex costs and time-varying step-size, without requiring the bounded heterogeneity condition (A5).
-In the Appendix we provide the full bound, containing additional terms, of order $\mathcal{O}\big(\frac{\log(T+t_{0}))}{(T+t_{0})^{3}}\big)$.
+Theorem 2 establishes HP convergence for smooth strongly convex costs and time-varying step-size.
+In Appendix F we provide the full bound, with additional higher-order terms.
 Next, we discuss multiple aspects of our results.
 Centralized SGD and linear speed-up.
-Compared to centralized SGD, which achieves the HP rate $\mathcal{O}\Big(\frac{1}{T}+\frac{1}{T^{2}}\Big)$, e.g., [liu2024revisiting], the rate in Theorem 2 also consists of terms of different order, namely $\mathcal{O}\Big(\frac{1}{nT}+\frac{1}{T^{2}}\Big)$.
-Similarly to non-convex results, the leading term shows a linear speed-up in the number of users, again matching the corresponding MSE guarantees, e.g., [pmlr-v119-koloskova20a].
+Compared to centralized SGD, which achieves the HP rate $\mathcal{O}\Big(\frac{1}{T}+\frac{1}{T^{2}}\Big)$, e.g., [41], the rate in Theorem 2 also consists of two terms of different order, namely $\mathcal{O}\Big(\frac{1}{nT}+\frac{1}{T^{2}}\Big)$.
+Similarly to non-convex results, the leading term shows a linear speed-up in the number of users, again matching the corresponding MSE guarantees, e.g., [33].
 Network effect and transient time.
-We can see that network connectivity again only affects higher-order terms.
-Next, we can use the rate in Theorem 2 to establish a transient time of DSGD in the HP sense.
-As shown in the Appendix, our HP results imply a transient time of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$, which is strictly sharper than the transient time of DSGD in, e.g., [unified-refined, Table II], stemming from MSE rates, which is of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$.
-The sharper transient time further highlights the tightness of our results and an improved analysis.
+We can again see that network connectivity only affects higher-order terms.
+Morevoer, the rate in Theorem 2 implies a transient time of DSGD in the HP sense, of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$.
+This is strictly sharper than the transient time of DSGD implied by MSE rates, of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see Table II in [1], further highlighting the tightness of our results and improved analysis.
 Effects of noise and heterogeneity.
-While heterogeneity again affects only higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see the Appendix for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule.
+While heterogeneity only affects higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see Appendix F for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule.
 Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size.
 Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
-Comparison with [dsgd-high-prob].
-The authors in [dsgd-high-prob] establish HP convergence of DSGD for the more general case of PŁ costs, providing a rate of $\mathcal{O}\big(\frac{1}{T}\big)$, using the same time-varying step-size schedule considered in Theorem 2.
-Crucially, the authors require two very strong assumptions: $(i)$ path-wise uniformly bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq G$, for all $t\geq 1$ and some $G>0$, almost surely; $(ii)$ asymptotically vanishing noise, i.e., for all $i\in[n]$ and any $t\geq 1$, the authors require the following condition
+Comparison with [44].
+The authors in [44] establish HP convergence of DSGD for the more general case of PŁ costs, providing a rate of $\mathcal{O}\big(\frac{1}{T}\big)$, using the same time-varying step-size schedule as in Theorem 2.
+Crucially, the authors require two very strong assumptions: $(i)$ path-wise uniformly bounded gradients, i.e., $\|\nabla f_{i}(x^{t}_{i})\|\leq G$, for some $G>0$ and all $t\geq 1$, almost surely; $(ii)$ asymptotically vanishing noise, i.e., for all $i\in[n]$ and any $t\geq 1$, they require the following
 $\mathbb{E}\left[\exp\bigg(\frac{\|z_{i}^{t}\|^{2}}{\alpha_{t}^{2}\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1),$
 where $\alpha_{t}=\frac{a}{t+t_{0}}$ is the step-size.
-Both conditions are very difficult to satisfy in practice, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is in general not satisfied even for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that the noise asymptotically vanishes, at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$.
-[^11] It can be easily verified via Markov’s inequality that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$.
-Similarly, it follows from Jensen’s inequality that $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$.
-Crucially, the authors in [dsgd-high-prob] are unable to establish linear speed-up in the number of users, even though much stronger assumptions on the noise and cost are imposed.
-On the other hand, our work uses standard assumptions, while simultaneously showing optimal convergence rates with linear speed-up and improved transient times compared to MSE results.
-These improvements stem not only from a sharper analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity assumption (A5) and ensure linear speed-up is achieved for strongly convex costs (see the Appendix for a detailed discussion).
+Both conditions are difficult to satisfy, as $(i)$ is a path-wise relaxation of the uniformly bounded gradient condition (which is not satisfied for strongly convex costs), while $(ii)$ implies that the noise is $\frac{\sigma_{i}}{t+t_{0}}$-sub-Gaussian at time $t$, meaning that it vanishes at rate $\mathcal{O}\big(\frac{1}{t+t_{0}}\big)$.
+[^9] Markov’s inequality implies that a $\frac{\sigma}{t+t_{0}}$-sub-Gaussian random variable $X$ satisfies $\mathbb{P}\Big(X^{2}\leq\frac{\sigma^{2}(\log(\nicefrac{{1}}{{\delta}})+1)}{(t+t_{0})^{2}}\Big)\geq 1-\delta$, for any $\delta\in(0,1)$.
+Similarly, Jensen’s inequality implies $\mathbb{E}[X^{2}]\leq\frac{\sigma^{2}}{(t+t_{0})^{2}}$.
+Crucially, [44] are unable to establish linear speed-up, even though they impose much stronger assumptions on the noise and cost.
+On the other hand, we show optimal convergence rate with linear speed-up using standard assumptions, with sharper transient times compared to MSE results.
+These improvements stem not only from a tighter analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity condition (A5) and ensure linear speed-up is achieved, see Appendix I for a further discussion.
+Numerical experiments.
+In Appendix G we provide numerical experiments which validate the main findings of the paper, notably showing that (i) DSGD converges in HP under the same conditions on the cost as in the MSE sense and (ii) DSGD achieves linear speed-up in the HP sense.
+The reader is kindly referred to the Appendix for details.
 4 Proof Outlines and Discussion
-In this section we provide some proof sketches, outline the main challenges, and discuss our results.
-Proof sketch of Theorem 1. Using Lemma 1, rearranging and summing up the first $T$ terms, we get
-$\sum_{t\in[T]}\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{L^{2}}{2}\sum_{t\in[T]}\frac{\alpha_{t}}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (4)
-We can use Lemma 3 to control the last term on the RHS.
-To deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$, while removing the need for uniformly bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [li2020high, liu2023high, armacki2025optimal], by subtracting $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of $\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ are absorbed by the left-hand side, allowing us to show that the resulting MGF is bounded.
-The rest of the proof then relies on Lemma 2, some technical results, careful selection of the step-size and Hölder’s inequality, see the Appendix for details.
+In this section we briefly outline proof sketches of the main results and provide further discuss on our results.
 Proof sketch of Theorem 1.
-Starting from Lemma 1, subtracting $f^{\star}$ from both sides, using properties of strongly convex functions, defining $F^{t}\coloneqq n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$ and $A_{t}\coloneqq(t+t_{0}+1)\alpha_{t}$, we get
-$F^{t+1}$ $\leq(1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}F^{t}-A_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\alpha_{t}A_{t}nL\|\overline{z}^{t}\|^{2}+\frac{A_{t}L^{2}}{2}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (5)
-We utilize the above inequality, Lemma 4, and a careful analysis to show that the MGF of $F^{t+1}$ satisfies the condition outlined in Lemma 5, after which we can use Lemma 5 to bound the MGF of the optimality gap.
-To complete the proof, we use the inequality [^12] See the Appendix for details.
-$f(x^{t}_{i})-f^{\star}\leq 2(f(\overline{x}^{t})-f^{\star})+\frac{L}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2},$
-and apply our results to show that both optimality and consensus gaps are small, with high probability.
-Discussion.
-Results of Theorems 1 and 2 are strong, demonstrating that DSGD converges in the HP sense under the same conditions on the cost function as in the MSE sense, while retaining order-optimal rates and linear speed-up in the number of users.
-Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we make several contributions.
+Using Lemma 1, rearranging and summing up the first $T$ terms, we get
+$\sum_{t\in[T]}\frac{\alpha_{t}}{2}\|\nabla f(\overline{x}^{t})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$
+$+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}+\frac{L^{2}}{2}\sum_{t\in[T]}\frac{\alpha_{t}}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (4)
+We use Lemma 3 to control the last term on the RHS, while to deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ and remove the need for bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [37, 39, 2], i.e., we subtract $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of the inner product is absorbed by the left-hand side.
+The rest of the proof relies on Lemma 2, some technical results and a careful selection of the step-size, see Appendix E for details.
+Proof sketch of Theorem 2.
+Using Lemma 1, properties of strongly convex functions, and defining $F^{t}\coloneqq n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$ and $A_{t}\coloneqq(t+t_{0}+1)\alpha_{t}$, we get
+$F^{t+1}$ $\leq(1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}F^{t}-A_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$
+$+\alpha_{t}A_{t}nL\|\overline{z}^{t}\|^{2}+\frac{A_{t}L^{2}}{2}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2}.$ (5)
+We use (4), Lemma 4, and a careful analysis to show that the MGF of $F^{t}$ satisfies the conditions of Lemma 5, which we can then apply.
+The proof is completed by using these results to show that both optimality and consensus gaps are small with high probability, see Appendix F for details.
+Comparison with centralized works.
+Compared to works studying HP convergence of centralized SGD, e.g., [21, 5, 39, 41], we make several contributions.
 First, we face the challenge of controlling the MGF of the consensus gap, stemming from the decentralized nature of the algorithm.
-To that end, we provide Lemma 3 for non-convex costs, removing the uniformly bounded gradient condition and using the bounded heterogeneity condition (A5) instead, as well as the improved Lemma 4 for strongly convex costs, which removes (A5) altogether.
-Next, we provide Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense and is of independent interest for HP studies of decentralized methods.
-To ensure linear speed-up in the number of users is achieved for strongly convex costs, we establish Lemma 5, which gives a more fine-grained result on the MGF of an “almost decreasing” process compared to existing (centralized) works, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2024revisiting], facilitating higher-order terms in the final bound, and is of independent interest (see the Appendix for a detailed discussion on the importance of Lemma 5 for achieving linear speed-up for strongly convex costs).
-Finally, to be able to utilize Lemma 5, a careful analysis of the MGF is needed, in order to simultaneously balance the effects of optimality and consensus gaps (recall that (5), which is used to bound the MGF of $F^{t+1}$, consists of both optimality and consensus gaps), while also ensuring that the resulting process is “almost decreasing”, in the sense of Lemma 5.
+To that end, we provide Lemma 3 for non-convex, as well as the improved Lemma 4 for strongly convex costs, which removes the bounded heterogeneity condition (A5).
+Next, we establish Lemma 2, which shows the variance reduction benefits of decentralized methods in the HP sense and is of independent interest when studying HP guarantees in decentralized settings.
+Finally, we provide a novel result on the MGF of the optimality gap of strongly convex costs (more broadly, “almost decreasing” processes), in the form of Lemma 5.
+This result provides a tighter control on the MGF compared to state-of-the-art bounds in [21, 5, 41] and is, in addition to Lemma 2, paramount to ensuring linear speed-up is achieved for strongly convex costs.
+In particular, following the same approach as in [21, 5], one can show that the MGF of the quantity of interest is uniformly bounded, i.e., that for some $B>0$ and all $t\geq 1$ and $\nu\in(0,B]$
+$\mathbb{E}[\exp(\nu F^{t})]\leq\exp\Big(\frac{\nu}{B}\Big),$ (6)
+where $F^{t}=n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$. Using (6) and Markov’s inequality, setting $\nu=B$, one can show that
+$\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}\leq\frac{\log(\nicefrac{{1}}{{\delta}})+1}{nB(t+t_{0})}\bigg)\geq 1-\delta,$ (7)
+losing the linear speed-up, as $B=\mathcal{O}(1/n)$, which we show in Appendix I.
+On the other hand, Lemma 5 provides a much sharper bound on the MGF, ensuring that linear speed-up is achieved.
+Similar observations hold with respect to [41], see Appendix I for details.
 On the dimension dependence.
-We note that our bounds show a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for either MSE or HP bounds, e.g., [pmlr-v119-koloskova20a, dsgd-high-prob].
-This dependence stems from the third result in Lemma 2, where the dimension figures in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4).
-In particular, to show the third property in Lemma 2, we use the fact that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-norm-sub-Gaussian, as a direct proof fails to yield the variance reduction benefit in the number of users.
-One way to avoid the (linear) dependence on $d$ is to assume that the noise is norm-sub-Gaussian, which would simultaneously relax the dependence on the dimension to $\log(d)$ and achieve the variance reduction benefit, at the cost of imposing a slightly stronger noise condition, see, e.g., Corollary 7 and the conclusion in [jin2019short]; see also the Appendix for a further discussion on the dimension dependence.
+Our bounds contain a factor of $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for MSE or HP bounds in [33, 44].
+This stems from Lemma 2, where $d$ appears in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4).
+One way to relax this dependence to $\log(d)$ is to assume that the noise is norm-sub-Gaussian, which comes with the cost of a slightly stronger noise condition, see, e.g., Corollary 7 in [27]; see also Appendix K for a further discussion on the dimension dependence.
 5 Conclusion
-In this paper, we revisit convergence in HP of a variant of DSGD under sub-Gaussian noise.
-We show that DSGD is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates for both non-convex and strongly convex costs.
-Moreover, our results show DSGD achieves linear speed-up in the number of users in both cases.
-Compared to [dsgd-high-prob], wherein the authors study HP convergence of DSGD, we relax strong conditions like uniformly bounded gradients or asymptotically vanishing noise, while showing improved rates, with linear speed-up in the number of users and transient times matching, or strictly sharper than transient times stemming from corresponding MSE rates.
-Compared to works studying HP convergence of centralized SGD, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a, liu2023high, liu2024revisiting], we provide several novel results, establishing tight control on the MGF of the consensus gap for both non-convex and strongly convex costs in Lemmas 3 and 4, a more fine-grained bound on the MGF of an “almost decreasing” process in Lemma 5, as well as showing variance reduction benefits of decentralized algorithms (in the HP sense) in Lemma 2.
-Future work includes extending our results to costs satisfying the PŁ condition, incorporating the GT mechanism to remove bounded heterogeneity condition and considering noise with heavier tails, like sub-Weibull.
+We studied convergence in HP of a variant of DSGD under light-tailed noise, showing that it is guaranteed to converge in the HP sense under the same conditions on the cost as in the MSE sense, achieving order-optimal rates and linear speed-up for both non-convex and strongly convex costs.
+Compared to existing works, we relax strong assumptions like uniformly bounded gradients and asymptotically vanishing noise, while simultaneously showing improved rates and sharper transient times than the ones obtained from MSE rates.
+Future work includes extending our results to costs satisfying the PŁ condition, incorporating bias-correction mechanisms like GT and exact diffusion to remove bounded heterogeneity, and considering heavy-tailed noise.
 Appendix A Introduction
@@
 $\mathbb{E}\Big[\prod_{i\in[n]}|X_{i}|\Big]\leq\prod_{i\in[n]}\sqrt[n]{\mathbb{E}|X_{i}|^{n}}.$
-Next, we state a useful result from [ran-improved].
-Proposition 6 ([ran-improved], Lemma 21).
+Next, we state a useful result from [75].
+Proposition 6 ([75], Lemma 21).
 For any $c,t_{0}>0$ and $0\leq a\leq b$, we have
-$\prod_{k=a}^{b}\Big(1-\frac{c}{t+t_{0}}\Big)\leq\frac{(a+t_{0})^{c}}{(b+1+t_{0})^{c}}.$
-The following result states some important consequences of conditions (A2) and (A3), see, e.g., [bertsekas-gradient, nesterov-lectures_on_cvxopt].
+$\prod_{k=a}^{b}\Big(1-\frac{c}{k+t_{0}}\Big)\leq\frac{(a+t_{0})^{c}}{(b+1+t_{0})^{c}}.$
+The following result states some important consequences of conditions (A2) and (A3), see, e.g., [9, 55].
 Proposition 7.
@@
     $\|\nabla f(x)\|^{2}\leq 2L(f(x)-f^{\star})$.
-Finally, we complete this section with the following important consequences of (A6), see, e.g., [nesterov-lectures_on_cvxopt].
+Finally, we complete this section with the following important consequences of (A6), see, e.g., [55].
 Proposition 8.
@@
   1. 1.
-    To prove the first property, we follow steps similar to those in [li2020high, Lemma 1].
+    To prove the first property, we follow steps similar to those from Lemma 1 in [37].
 Let $y_{i}=\frac{z^{t}_{i}}{\sigma_{i}}$ and note that $\mathbb{E}\left[\exp\left(\|y_{i}\|^{2}\right)\>|\>\mathcal{F}_{t}\right]\leq\exp(1)$, from (A3).
@@
   3. 3.
-    Combining the previous result with Lemma 1 from [jin2019short], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
+    Combining the previous result with Lemma 1 from [27], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
     $\mathbb{P}\big(\|\overline{z}^{t}\|>\epsilon\big)\leq 2\exp\bigg(-\frac{n\epsilon^{2}}{24\sigma^{2}d}\bigg).$
-    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., [vershynin_2018, Proposition 2.5.2], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
+    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., Proposition 2.5.2 in [67], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
 We next restate Lemma 5.
@@
 Combining the definition of the network average model, the fact that the weight matrix is doubly stochastic and (14), it follows that $\overline{x}^{t+1}=\overline{x}^{t}-\alpha_{t}\overline{g}^{t}$.
-We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [nedic-subgrad, jakovetic-unification, ran-improved].
+We now introduce some notation useful for the analysis of decentralized methods, see, e.g., [53, 26, 75].
 Let ${\mathbf{x}}^{t}\coloneq col(x_{1}^{t},\ldots,x_{n}^{t})\in\mathbb{R}^{nd}$ denote the column vector stacking users’ local models.
@@
 [^10] This is equivalent to the client-server setup in terms of the update rule, as the server averages models from all clients in each iteration.
-The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [jakovetic-unification, ran-improved] and references therein.
+The interaction between matrices $W$ and $J$ represents an important part of any decentralized algorithm and we now list some known properties, see, e.g., [26, 75] and references therein.
 Proposition 9.
@@
   2. 2.
-    $\|W-J\|=\lambda$, where $\lambda\in[0,1)$ is the second largest singular value of $W$.
+    $\lambda\coloneqq\|W-J\|\in[0,1)$.
   3. 3.
@@
 Proof.
-The proof follows similar steps as in [ran-improved, Lemma 3]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
+The proof follows similar steps as in Lemma 3 from [75]. Starting from Proposition 7 and averaging across all costs $i\in[n]$, it readily follows that, for all $x,y\in\mathbb{R}^{d}$
 $f(x)\leq f(y)+\langle\nabla f(y),x-y\rangle+\frac{L}{2}\|x-y\|^{2}.$ (17)
@@
 We are now ready to prove Theorem 1, which we restate next, for convenience.
-Theorem 1.
+Theorem 3.4.
 Let (A1)-(A5) hold.
@@
     as the last term, which does not attain a linear speed-up, is no longer of higher order.
-We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [pmlr-v119-koloskova20a, ran-improved].
+We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [33, 75].
 Appendix F Proofs for Strongly Convex Costs
 In this section we prove Lemma 4 and Theorem 2.
-To do so, we follow a similar strategy to the one in, e.g., [conv-rates-dsgd-jakovetic, ran-improved], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense.
+To do so, we follow a similar strategy to the one in, e.g., [24, 75], where it is shown that the sequence of iterates generated by DSGD is bounded in the MSE sense.
 However, to establish guarantees in the HP sense, we instead work with the MGF of the iterates and have the following important result.
@@
 Let (A1)-(A4) and (A6) hold, let $a,t_{0},K>0$ and the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$, and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$.
-If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$
+If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\Big\{6,\frac{288\overline{\sigma}^{2}K}{\mu^{2}},\frac{3456\overline{\sigma}^{2}\lambda^{2}K}{\mu^{2}(1-\lambda)},\frac{12\lambda L\sqrt{10}}{\mu(1-\lambda)}\Big\}$, then for $K_{t+1}=(t+t_{0}+2)K$ and any $\nu\leq\min\left\{1,\frac{\mu^{2}}{144\sigma^{2}K}\right\}$, we have
 $\mathbb{E}\Big[\exp\Big(\nu K_{t+1}\sum_{i\in[n]}\|x^{t+1}_{i}-\overline{x}^{t+1}\|^{2}\Big)\Big]\leq\exp\bigg(\nu K_{t+1}\Big(\sum_{k\in[t]}\lambda^{t-k}S_{k}+\sum_{k\in[t]}\lambda^{t-k}D_{k}\Big)\bigg),$
-where $S_{k}=\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}=\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{5}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{6}}\Big)$.
+where $S_{k}\coloneqq\alpha_{k}^{2}\lambda^{2}\Big(n\sigma^{2}+\frac{5\|\nabla\mathbf{f}^{\star}\|^{2}}{1-\lambda}\Big)$ and $D_{k}\coloneqq\frac{5\alpha_{k}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{k}}{5}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(k+t_{0})^{6}}\Big)$.
 Proof.
@@
 We are now ready to prove Theorem 2. Prior to that, we restate it, for convenience.
-Theorem 2.
+Theorem 3.7.
 Let (A1)-(A4) and (A6) hold, let the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$.
@@
 Finally, it can be verified that the conditions on $a$, $t_{0}$ and $\nu$ in the statement of the theorem ensure that all the step-size conditions are satisfied, completing the proof. ∎
-We remark that, similarly to the proof of Theorem 1, one can analyze the strongly convex case with a fixed step-size, resulting in the dependence on some terms, e.g., optimality and iterate gaps $\Delta_{f}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}}^{\star}\|$, decaying exponentially fast, i.e., $\mathcal{O}\big((\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}+\Delta_{f})e^{-CT}\big)$, for some $C>0$, as shown in, e.g., [pmlr-v119-koloskova20a] for MSE guarantees in decentralized, or [liu2024revisiting] for HP guarantees in centralized settings.
+We remark that, similarly to the proof of Theorem 1, one can analyze the strongly convex case with a fixed step-size, resulting in the dependence on some terms, e.g., optimality and iterate gaps $\Delta_{f}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}}^{\star}\|$, decaying exponentially fast, i.e., $\mathcal{O}\big((\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}+\Delta_{f})e^{-CT}\big)$, for some $C>0$, as shown in, e.g., [33] for MSE guarantees in decentralized, or [41] for HP guarantees in centralized settings.
 For simplicity, we omit this analysis.
@@
 We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive semi-definite, making each $f_{i}$ strongly convex, with unbounded gradients.
-We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [XIAO200465].
+We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [73].
 When queried by user $i\in[n]$ in iteration $t\geq 1$, the $\mathcal{SFO}$ returns a noisy gradient of $f_{i}$ evaluated at $x^{t}_{i}$, i.e., $g^{t}_{i}=A_{i}x^{t}_{i}+b_{i}+z^{t}_{i}$, where $z^{t}_{i}\in\mathbb{R}^{d}$ is a zero-mean Gaussian random vector, making the noise consistent with assumption (A4).
@@
 To verify that linear speed-up in the number of users is achieved, we consider three networks, with $n=\{25,50,100\}$ users.
-In all three cases we use Erdős–Rényi communications graphs which satisfy $\lambda=\|W-J\|\approx 0.6$.
-To ensure that the effect of heterogeneity is consistent across different networks, we fix the matrices $A_{i}$, i.e., $A_{i}\equiv A=S+I_{d}$, where $S\in\mathbb{R}^{d\times d}$ is generated using sklearn’s make_sparse_spd_matrix, with $d=50$ and $\texttt{alpha}=0.9$, while vectors $b_{i}$ are given by $b_{i}=\beta_{i}\mathbf{1}_{d}$, where $\beta_{i}\in\{-2,-1,0,1,3\}$ are selected uniformly at random and in equal proportion across users (i.e., one fifth of users has $\beta_{i}=-2$, one fifth $\beta_{i}=-1$, etc).
+In all three cases, we use Erdős–Rényi communications graphs with Metropolis-Hastings weights that satisfy $\lambda=\|W-J\|\approx 0.6$.
+To ensure that the effect of heterogeneity is consistent across different networks, we fix the matrices $A_{i}$, i.e., $A_{i}\equiv A=S+I_{d}$, where $S\in\mathbb{R}^{d\times d}$ is generated using sklearn’s make_sparse_spd_matrix, with $d=50$ and $\texttt{alpha}=0.9$, while vectors $b_{i}$ are given by $b_{i}=\beta_{i}\mathbf{1}_{d}$, where $\beta_{i}\in\{-2,-1,0,1,3\}$ are selected uniformly at random and in equal proportion across users (i.e., one fifth of users has $\beta_{i}=-2$, one fifth has $\beta_{i}=-1$, etc).
 Generating the networks and costs in this manner ensures that both network connectivity and user heterogeneity are constant across all three network settings, allowing us to properly capture the effect of linear speed-up.
@@
 To compute the empirical probability, in each iteration we draw $S=600$ Monte-Carlo samples, with thresholds $\varepsilon=\big\{10^{-2},10^{-3},10^{-4}\big\}$.
-The results are presented in Figure 2, where the plots left to right and top to bottom show the MSE and tail probability behaviours for different values of $\varepsilon$, respectively.
+The results are presented in Figure 2, where the plots left to right and top to bottom respectively show the MSE and tail probability behaviours for different values of $\varepsilon$.
 We can again see that the empirical tail probability decays exponentially fast for all values of $\varepsilon$, with the decay being consistently faster for larger number of users, demonstrating linear speed-up in the HP sense.
@@
 In this section we provide details on the transient time resulting from our HP bounds.
-Similarly to [unified-refined], we focus on the dependence on network connectivity and the number of users, ignoring other problem related constants.
+Similarly to [1], we focus on the dependence on network connectivity and the number of users, ignoring other problem related constants.
 First, consider non-convex costs.
@@
 It is now evident that, for any $T\geq\max\Big\{\frac{n}{(1-\lambda)^{2}},\frac{n^{3}}{(1-\lambda)^{4}}\Big\}=\frac{n^{3}}{(1-\lambda)^{4}}$, the bound in (48) becomes $\mathcal{O}\Big(\frac{1}{\sqrt{nT}}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$.
-This matches the transient time of DSGD obtained from the MSE rate in [pmlr-v119-koloskova20a], see, e.g., [unified-refined, Table I].
+This matches the transient time of DSGD obtained from the MSE rate in [33], see, e.g., Table I in [1].
 Next, consider strongly convex costs. Recalling the expressions in (44) and (F), it follows that the full bound for strongly convex costs is of the form
@@
 where $t_{0}$ depends on the network connectivity via the expression $t_{0}=\Omega\Big(\max\Big\{\frac{\lambda}{1-\lambda},\frac{\lambda^{2}}{1-\lambda},\frac{1}{1-\lambda}\Big\}\Big)$.
-Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{\lambda}{1-\lambda}\Big)$, so that the bound in (H) becomes
+Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{1}{1-\lambda}\Big)$, so that the bound in (H) becomes
 $\quad\quad\mathcal{O}\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}$
@@
 It can be seen that, for any $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (H) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$.
-Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [pmlr-v119-koloskova20a], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., [unified-refined, Table II].
+Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [33], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., Table II in [1].
 As such, our HP analysis provides either matching or strictly sharper transient times for DSGD compared to the transient times stemming from the corresponding MSE results, further highlighting the tightness of our analysis.
@@
 where it is easy to see that choosing $\alpha=\Theta\Big(\frac{\sqrt{n}}{\sqrt{T}}\Big)$ leads to the desired linear speed-up.
-Works like [dsgd-high-prob, dsgd-online-noncooperative-games, dsgd-online-stochastic] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity.
+Works like [44, 45, 76] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity.
 As such, even if a fixed step-size was used, the authors would be unable to obtain a bound of the form (52) and achieve linear speed-up.
@@
 $\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}+\frac{n}{T}\bigg),$
-where the term not achieving linear speed-up is of higher order.
-This is consistent with MSE guarantees, e.g., [pmlr-v119-koloskova20a, ran-improved, unified-refined], where a fixed step-size is also needed to achieve linear speed-up.
+where the term not achieving linear speed-up is of higher order. This is consistent with MSE guarantees, e.g., [33, 75, 1], where a fixed step-size is also needed to achieve linear speed-up.
 Linear speed-up for strongly convex costs.
 For strongly convex costs, in addition to Lemma 2, a tighter control of the MGF of the process of interest is required, which is achieved in Lemma 5.
-To see why this is the case, consider the existing techniques in centralized settings, e.g., [pmlr-v99-harvey19a, pmlr-v206-bajovic23a], where the authors show via induction that the MGF of the process of interest is uniformly bounded, i.e., that there exists a $B>0$, such that, for all $t\geq 1$
-$\mathbb{E}[\exp(BF^{t})]\leq\exp(1).$ (51)
+To see why this is the case, consider the existing techniques in centralized settings, e.g., [21, 5], where the authors show via induction that the MGF of the process of interest is uniformly bounded, i.e., that there exists a $B>0$, such that, for all $t\geq 1$ and $\nu\in(0,B]$
+$\mathbb{E}[\exp(\nu F^{t})]\leq\exp\Big(\frac{\nu}{B}\Big).$ (53)
 Recalling that in our work $F^{t}=n(t+t_{0})\big(f(\overline{x}^{t})-f^{\star}\big)$, it then follows from (53) and Markov’s inequality that, for any $\epsilon>0$
-$\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}>\frac{\epsilon+1}{nB(t+t_{0})}\bigg)=\mathbb{P}\big(BF^{t}>\epsilon+1\big)\leq\exp(-\epsilon),$
-or equivalently, with probability at least $1-\delta$
+$\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}>\frac{\epsilon+1}{n\nu(t+t_{0})}\bigg)=\mathbb{P}\big(\nu F^{t}>\epsilon+1\big)\leq\exp(-\epsilon).$
+Choosing $\nu=B$, it follows that, for any $\delta\in(0,1)$, with probability at least $1-\delta$
 $f(\overline{x}^{t})-f^{\star}\leq\frac{\log(\nicefrac{{1}}{{\delta}})+1}{nB(t+t_{0})}.$ (54)
@@
 To see that this is the case, we go back to (53) and note that, in order for the identity to hold for $t=1$, we have
-$\mathbb{E}[\exp(BF^{1})]=\exp\Big(nB(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)\Big)\leq\exp(1),$
-where the equality follows from the definition of $F^{t}$ and the fact that the initial models are chosen deterministically, while the inequality follows if and only if $B\leq\frac{1}{n(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)}$.
+$\mathbb{E}[\exp(\nu F^{1})]=\exp\Big(\nu n(t_{0}+1)\big(f(\overline{x}^{1})-f^{\star}\big)\Big)\leq\exp\Big(\frac{\nu}{B}\Big),$
+where the equality follows from the definition of $F^{t}$ and the fact that the initial models are chosen deterministically, while the inequality follows if and only if $B\leq\frac{1}{n(t_{0}+1)(f(\overline{x}^{1})-f^{\star})}$.
 As such, in order for (53) to hold, the constant $B$ has to decay with $n$, canceling the linear speed-up in (54) and resulting in a HP bound of the form
@@
 for some problem related constants $G_{i}>0$, $i\in[4]$, providing a much tighter and more fine-grained control of the MGF compared to (53) and crucially allowing for a bound where the leading term achieves linear speed-up in the number of users (as $G_{1}$ is independent of $n$).
-On the other hand, using a different approach to [pmlr-v99-harvey19a, pmlr-v206-bajovic23a], the work [liu2024revisiting] establishes the following HP bound for the last iterate of SGD in centralized setting and smooth, strongly convex costs
+On the other hand, using a different approach to [21, 5], the work [41] establishes the following HP bound for the last iterate of SGD in centralized setting and smooth, strongly convex costs
 $f(x^{t})-f^{\star}=\mathcal{O}\bigg(\frac{\log(t)\log(\nicefrac{{1}}{{\delta}})}{t+t_{0}}+\frac{1}{(t+t_{0})^{2}}\bigg).$
-While such result could potentially be leveraged in decentralized settings to maintain linear speed-up (provided that the constant growing linearly with $n$ affects only the higher-order term), it still accrues a factor of $\log(t)$ in the leading term, whereas in our bound it appears only in the higher-order term, decaying at rate $\mathcal{O}\Big(\frac{\log(t+t_{0})}{(t+t_{0})^{3}}\Big)$.
+While such result could potentially be leveraged in decentralized settings to maintain linear speed-up (provided that the constant growing linearly with $n$ affects only the higher-order term), it still accrues a factor of $\log(t)$ in the leading term, whereas in our bound the $\log$ term only affects a higher-order term, decaying at rate $\mathcal{O}\Big(\frac{\log(t+t_{0})}{(t+t_{0})^{3}}\Big)$.
 As such, Lemma 5 provides the most fine-grained control of the MGF and is of independent interest when studying “almost decreasing” type processes, even in centralized settings.
@@
 As discussed in the main body, to model the communication network, we use assumption (A1), which subsumes fixed undirected networks, as well as a class of fixed directed networks.
-The authors in [dsgd-high-prob] consider a more general case of directed, time-varying networks, facilitated by the following condition.
+The authors in [44] consider a more general case of directed, time-varying networks, facilitated by the following condition.
 […]
@@
 To see why this is the case, note that in our proofs we only use the fact that under (A1), we have $\lambda=\|W-J\|\in[0,1)$.
-The authors in [dsgd-high-prob, Lemma 1] use a similar result for time-varying networks, namely that under (A1), for any time indices $t\geq s$
+The authors in [44] use a similar result for time-varying in networks (see Lemma 1 [44]), namely that under (A1), for any time indices $t\geq s$
 $\Big|[\Phi(t,s)]_{ij}-\frac{1}{n}\Big|\leq M\rho^{t-s},$ (55)
 where $\Phi(t,s)=W^{t-1}\times\ldots\times W^{s}$ is the transition matrix from time $s$ to $t$ (with $\Phi(t,t)=I_{n}$), while $M>0$ and $\rho\in(0,1)$ are graph related constants.
-As discussed in [dsgd-high-prob], the constant $\rho$ plays the same role of the network connectivity parameter, as $\lambda$ does in our work.
+As discussed in [44], the constant $\rho$ plays the same role of the network connectivity parameter, as $\lambda$ does in our work.
 It can be readily verified that replacing (A1) with (A1) and using (55) to control the consensus gap, our results readily go through, with $\lambda$ being replaced by $\rho$.
@@
 $\mathbb{P}\big(\|x\|>\epsilon\big)\leq 2\exp\bigg(-\frac{\epsilon^{2}}{2\sigma^{2}d}\bigg),$
-see [jin2019short, Lemma 1].
+see Lemma 1 in [27].
 This argument allows us to show the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian (in the sense of (A4)), at the cost of introducing a $\sqrt{d}$ dependence.
 One way to mitigate this is to directly assume that the noise $z^{t}_{i}$ at each user is zero-mean and $\sigma_{i}$-norm-sub-Gaussian.
-Using a similar argument to the one in Lemma 6 and Corollary 7 in [jin2019short], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$.
+Using a similar argument to the one in Lemma 6 and Corollary 7 in [27], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$.
 This, however, comes at the cost of imposing a slightly stronger noise condition, as any $\sigma$-norm-sub-Gaussian random vector is also $\sigma$-sub-Gaussian (in the inner product sense), while the opposite implication inevitably introduces a $\sqrt{d}$ factor, as is the case in Lemma 2.

### v5

@@
 Another line of work studies HP convergence under heavy-tailed noise, [^2] Since heavy-tailed noise is not the focus of our work, we point the reader to [53, 60, 4] for various conditions used to study heavy-tailed noise.
-e.g., [61, 56, 40, 58, 23, 35, 3, 2, 4], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1).
+e.g., [21, 68, 60, 43, 63, 25, 37, 3, 2, 4], where it is necessary to introduce algorithmic modifications, e.g., clipping, normalization, or sign, to ensure concentration of the form in (1).
 Crucially, convergence in the HP sense is achieved under the same conditions on the cost function as in the MSE sense, for both light-tailed and heavy-tailed noise.
@@
 [64, 84] study HP convergence of decentralized algorithms with clipping, for convex costs under heavy-tailed noise, while [85] also consider general non-convex costs.
+Finally, [22] study clipped SGD under heavy-tailed noise, in the distributed client-server (i.e., fully connected network) setting.
 It is worth mentioning works like [7, 6, 50, 51, 8], where the authors study asymptotic large deviation guarantees for decentralized problems such as detection and inference.
@@
 Assumption (A1) is satisfied by connected undirected graphs, as well as a class of strongly-connected directed graphs with doubly stochastic weights, e.g., [81].
-[^5] Our work can be readily extended to the setting of time-varying networks, see Appendix J for a detailed discussion.
+[^5] Our work can be readily extended to the setting of time-varying networks, see Appendix K for a detailed discussion.
 Let $\lambda\coloneqq\|W-J\|$ denote the network connectivity parameter, where $J\coloneqq\frac{1}{n}\mathbf{1}_{n}\mathbf{1}_{n}^{\top}\in\mathbb{R}^{n\times n}$ is the ideal communication matrix.
@@
 Assumptions (A2)-(A3) are standard in smooth non-convex optimization, e.g., [20, 82].
+While some works, e.g., [38], assume gradients of loss $\ell$ are $L$-Lipschitz continuous, we note that while similar, such assumption and (A3) aim to exploit different structural properties.
+However, it is known that under certain regularity conditions, $L$-Lipschitz gradients of $\ell$ imply $L$-Lipschitz gradients of $f_{i}$, making such condition slightly stronger than (A3).
 The stochastic quantities satisfy the following:
@@
   3. 3.
-    The noise at user $i\in[n]$ is $\sigma_{i}$-sub-Gaussian (light-tailed), i.e., for all $t\geq 1$ and $\mathcal{F}_{t}$-measurable $x\in\mathbb{R}^{d}$
+    The gradient noise $z^{t}_{i}$ at user $i\in[n]$ and time $t\geq 1$ is $\sigma_{i}$-sub-Gaussian, i.e., we have
     $\mathbb{E}\left[\exp\bigg(\frac{\|z_{i}^{t}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1).$
-The first condition in (A4) is standard in decentralized stochastic optimization [75, 79], while the second and third require noise to be unbiased and light-tailed (i.e., sub-Gaussian).
+The first condition in (A4) is standard in decentralized stochastic optimization [82, 86], while the second and third require noise to be unbiased and sub-Gaussian (i.e., light-tailed).
 Light tails are necessary to achieve exponentially decaying tail bounds of the form in (1) if the vanilla stochastic gradient estimator is employed without any modifications (e.g., using clipping operator or estimators like median-of-means) and are widely used in centralized settings, e.g., [58, 20, 39, 40, 23, 42, 5].
 [^6] While it is possible to achieve concentration of the form in (1) under, e.g., sub-Weibull noise, see [49] and references therein, the focus of this work is on sub-Gaussian noise.
-The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
+The gradients of users have bounded heterogeneity, i.e., for all $x\in\mathbb{R}^{d}$ and some $A,B\geq 0$, [^7] With at least one of $A,B$ being strictly positive.
+we have $\max_{i\in[n]}\|\nabla f_{i}(x)\|^{2}\leq A^{2}+B^{2}\|\nabla f(x)\|^{2}$.
 A heterogeneity bound of the type in (A5) is in fact required to ensure that DSGD converges for non-convex costs [41, 14, 35, 79].
@@
   3. 3.
-    The average noise $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian, i.e., $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{120d\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1)$.
-Lemma 2 provides some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian, establishing the variance reduction benefit of decentralized optimization in the HP sense.
+    The average noise $\overline{z}^{t}$ satisfies $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{15\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp\big(1+\log(2d)\big)$.
+Lemma 2 provides some properties of noise, importantly showing that the average noise is $\mathcal{O}\Big(\frac{\sigma}{\sqrt{n}}\Big)$-sub-Gaussian, establishing the variance reduction benefit of decentralized optimization in the HP sense.
 This result is crucial toward showing that DSGD achieves a linear speed-up in the number of users, which we do in the following subsections.
-Note that the dependence on problem dimension $d$ is unavoidable due to the definition of sub-Gaussianity used in (A4), with further discussion provided in Section 4 and Appendix K.
+We note that the third property introduces a mild logarithmic dependence on the problem dimension $d$, with further discussion provided in Section 4 and Appendix L.
 3.2 Non-Convex Costs
@@
 Let (A1)-(A5) hold.
-If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
-$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$
+If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{15LT}}\right\}$, where $C>0$ is a problem related constant satisfying
+$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$
 then for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
-Theorem 1 establishes HP convergence of DSGD for general smooth non-convex costs and a fixed step-size. We now discuss the bound from a few different perspectives.
+$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{L}\big(\Delta_{f}+\log(\nicefrac{{2d}}{{\delta}})\big)}{\sqrt{nT}}+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}(1-\lambda)^{2}T}\Bigg).$
+Theorem 1 establishes HP convergence of DSGD to a stationary point (or a set of stationary points), for general smooth non-convex costs and a fixed step-size.
+We now discuss the bound from a few different perspectives.
 Centralized SGD and linear speed-up.
@@
 Moreover, the rate in Theorem 1 implies that the transient time of DSGD in the HP sense is of order $\mathcal{O}\Big(\frac{n^{3}}{(1-\lambda)^{4}}\Big)$, matching the transient time of DSGD implied by MSE rates, see, e.g., Table I in [1].
-For details on the derivation of the transient time, see Appendix H.
+For details on the derivation of the transient time, see Appendix I.
 Effects of noise and heterogeneity.
@@
 Next, [47] are unable to establish a linear speed-up, which is a byproduct of both their step-size choice, as well as their treatment of the noise.
-In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, see Appendix E for a detailed discussion.
+In particular, choosing a time-varying step-size for non-convex costs results in a loss of linear speed-up, see Appendix F for a detailed discussion.
 However, it would be impossible for [47] to obtain a linear speed-up even under a fixed step-size, as they show no variance reduction benefit of decentralized learning in the HP sense, instead only using the definition of sub-Gaussianity.
-On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used, see Appendix I for detailed discussion on the results needed to achieve linear speed-up.
+On the other hand, the variance reduction benefits of decentralized learning established in Lemma 2 allow us to achieve linear speed-up in the number of users when a fixed step-size is used, see Appendix J for detailed discussion on the results needed to achieve linear speed-up.
 [^9] This is consistent with MSE results, e.g., [35, 82, 1], where linear speed-up in non-convex case is achieved only via a fixed step-size.
-Finally, while we consider static networks, our results can be extended to the time-varying setting in [44], see Appendix J for details.
+Finally, while we consider static networks, our results can be extended to the time-varying setting in [47], see Appendix K for details.
 3.3 Strongly Convex Costs
@@
 where $S_{k},D_{k}>0$ are problem related constants.
-Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5), with $S_{k},D_{k}$ depending on the step-size and other problem parameters, see Appendix F for explicit forms of $S_{k},D_{k}$.
+Lemma 4 gives a tight bound on the MGF of the consensus gap, without requiring the bounded heterogeneity condition (A5), with $S_{k},D_{k}$ depending on the step-size and other problem parameters, see Appendix G for explicit forms of $S_{k},D_{k}$.
 We note that the requirement of same model initialization across users can be removed, at the cost of an additional, geometrically decaying term in the exponent on the RHS, however, we use same initialization, for ease of exposition.
@@
 Lemma 5 provides a tight bound on the MGF of an “almost decreasing” process, with the optimality gap of strongly convex costs being an instance of such a process.
-Compared to bounds used in the centralized setting, e.g., [21, 5, 41], Lemma 5 is significantly sharper and incorporates higher-order terms, which are crucial for ensuring linear speed-up, see Section 4 and Appendix I for detailed discussions.
+Compared to bounds used in the centralized setting, e.g., [23, 5, 44], Lemma 5 is significantly sharper and incorporates higher-order terms, which are crucial for ensuring linear speed-up, see Section 4 and Appendix J for detailed discussions.
 We are now ready to state the main result for strongly convex costs.
@@
 Let (A1)-(A4) and (A6) hold, let the step-size schedule be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$.
-If constants $a$ and $t_{0}$ are chosen as $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
-$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+d\sigma^{2}\kappa/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(T+t_{0})^{2}}\bigg),$
+If constants $a$ and $t_{0}$ are chosen as $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{1960\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
+$\frac{1}{n}\sum_{i\in[n]}$ $\big(f(x_{i}^{T})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}\kappa\big(1+\log(2d)\big)/\mu}{n(T+t_{0})}+\frac{\lambda^{2}L(1+L)(n\sigma^{2}+\nicefrac{{\|\nabla\mathbf{f}^{\star}\|^{2}(1+\kappa^{2})}}{{(1-\lambda))}}}{(1-\lambda)n(T+t_{0})^{2}}\bigg),$
 where $\mathcal{O}(\cdot)$ hides some higher-order terms.
-Theorem 2 establishes HP convergence for smooth strongly convex costs and time-varying step-size.
-In Appendix F we provide the full bound, with additional higher-order terms.
+Theorem 2 establishes HP convergence of DSGD to the global minumum, for smooth strongly convex costs and time-varying step-size.
+In Appendix G we provide the full bound, with additional higher-order terms.
 Next, we discuss multiple aspects of our results.
@@
 We can again see that network connectivity only affects higher-order terms.
-Morevoer, the rate in Theorem 2 implies a transient time of DSGD in the HP sense, of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$.
+Moreover, the rate in Theorem 2 implies a transient time of DSGD in the HP sense, of order $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$.
 This is strictly sharper than the transient time of DSGD implied by MSE rates, of order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see Table II in [1], further highlighting the tightness of our results and improved analysis.
 Effects of noise and heterogeneity.
-While heterogeneity only affects higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see Appendix F for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule.
+While heterogeneity only affects higher-order terms (through $\|\nabla\mathbf{f}^{\star}\|^{2}$ and $\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}$, see Appendix G for the full bound), we can see that the noise affects both the leading and higher-order terms, stemming from the time-varying step-size schedule.
 Similarly to the discussion after Theorem 1, the effect of noise on the leading term can be mitigated by employing a carefully selected fixed step-size.
 Due to time-varying step-sizes being sufficient for achieving linear speed-up and for ease of exposition, the results using a fixed step-size are omitted.
+Extending to slower decaying step-size and Polyak-Ruppert average.
+While Theorem 2 provides the optimal convergence rate for strongly convex costs, it is also of interest to consider step-sizes that decay at a slightly slower rate, e.g., $\alpha_{t}\propto(t+t_{0})^{-\eta}$, for $\eta\in[1/2,1)$.
+The techniques used in our proof can readily handle such step-sizes, but would inevitably lead to sub-optimal rates of order $\mathcal{O}\Big(\frac{1}{nt^{\eta}}\Big)$, while noting that the step-size schedule $\alpha_{t}\propto(t+t_{0})^{-1}$ used in Theorem 2 is known to be optimal for strongly convex costs, e.g., [66, 44].
+Additionally, for (strongly) convex costs, it is often of interest to provide convergence guarantees in terms of the Polyak-Ruppert average model [67, 61].
+Such guarantees readily follow, e.g., from Theorem 1 and using (A6).
+Indeed, defining the Polyak-Ruppert average as $\widetilde{x}^{t}=\frac{1}{nT}\sum_{i\in[n],t\in[T]}x^{t}_{i}$, using the results of Theorem 1 with step size $\alpha\propto\frac{1}{\sqrt{nT}}$ and (A6), we get
+$2\mu\big(f(\widetilde{x}^{t})-f^{\star})\leq\frac{1}{nT}\sum_{i\in[n],t\in[T]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}\bigg).$
+However, we note that the guarantees of Theorem 2 are stronger in multiple facets: (i) the step-size and rate in Theorem 2 are optimal for strongly convex costs; (ii) the last iterate is practically more viable than the Polyak-Ruppert average and does not require the users to maintain an extra variable (i.e., the running average); (iii) the analysis for the last iterate is significantly more involved, as demonstrated in the appendix, whereas the result for Polyak-Ruppert average follow directly from Theorem 1 and the properties of strong convexity.
 Comparison with [47].
@@
 On the other hand, we show optimal convergence rate with linear speed-up using standard assumptions, with sharper transient times compared to MSE results.
-These improvements stem not only from a tighter analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity condition (A5) and ensure linear speed-up is achieved, see Appendix I for a further discussion.
+These improvements stem not only from a tighter analysis and Lemma 2, but also Lemmas 4 and 5, which allow us to remove the bounded heterogeneity condition (A5) and ensure linear speed-up is achieved, see Appendix J for a further discussion.
 Numerical experiments.
-In Appendix G we provide numerical experiments which validate the main findings of the paper, notably showing that (i) DSGD converges in HP under the same conditions on the cost as in the MSE sense and (ii) DSGD achieves linear speed-up in the HP sense.
+In Appendix H we provide numerical experiments which validate the main findings of the paper, notably showing that (i) DSGD converges in HP under the same conditions on the cost as in the MSE sense and (ii) DSGD achieves linear speed-up in the HP sense.
 The reader is kindly referred to the Appendix for details.
@@
 We use Lemma 3 to control the last term on the RHS, while to deal with $\sum_{t\in[T]}\alpha_{t}\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle$ and remove the need for bounded gradients, we use Lemma 2 and the “offset trick”, e.g., [40, 42, 2], i.e., we subtract $\sum_{t\in[T]}\frac{9\sigma^{2}\alpha_{t}^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{4n}$ from both sides of (4), ensuring that the effects of the inner product is absorbed by the left-hand side.
-The rest of the proof relies on Lemma 2, some technical results and a careful selection of the step-size, see Appendix E for details.
+The rest of the proof relies on Lemma 2, some technical results and a careful selection of the step-size, see Appendix F for details.
 Proof sketch of Theorem 2.
@@
 We use (4), Lemma 4, and a careful analysis to show that the MGF of $F^{t}$ satisfies the conditions of Lemma 5, which we can then apply.
-The proof is completed by using these results to show that both optimality and consensus gaps are small with high probability, see Appendix F for details.
+The proof is completed by using these results to show that both optimality and consensus gaps are small with high probability, see Appendix G for details.
 Comparison with centralized works.
@@
 $\mathbb{P}\bigg(f(\overline{x}^{t})-f^{\star}\leq\frac{\log(\nicefrac{{1}}{{\delta}})+1}{nB(t+t_{0})}\bigg)\geq 1-\delta,$ (7)
-losing the linear speed-up, as $B=\mathcal{O}(1/n)$, which we show in Appendix I.
+losing the linear speed-up, as $B=\mathcal{O}(1/n)$, which we show in Appendix J.
 On the other hand, Lemma 5 provides a much sharper bound on the MGF, ensuring that linear speed-up is achieved.
-Similar observations hold with respect to [41], see Appendix I for details.
+Similar observations hold with respect to [44], see Appendix J for details.
 On the dimension dependence.
-Our bounds contain a factor of $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, which is not the case for MSE or HP bounds in [33, 44].
-This stems from Lemma 2, where $d$ appears in the sub-Gaussianity constant and is unavoidable, due to the definition of sub-Gaussianity used in (A4).
-One way to relax this dependence to $\log(d)$ is to assume that the noise is norm-sub-Gaussian, which comes with the cost of a slightly stronger noise condition, see, e.g., Corollary 7 in [27]; see also Appendix K for a further discussion on the dimension dependence.
+Our bounds contain a factor of $\log(d)$, introducing a mild dependence on the problem dimension, which is not the case for either MSE results [35], or HP bounds in [47].
+This dependence stems from Lemma 2, where we show that variance-reduction is achieved in the HP sense, with $\log(d)$ appearing in the exponent bounding the MGF of $\overline{z}^{t}$.
+While mild, it is not clear if this dependence on the problem dimension can be completely removed for sub-Gaussian vectors, as noted also in [29].
+For a further discussion on sub-Gaussian vectors and dimension dependence, see Appendix L.
 5 Conclusion
@@
 Appendix A Introduction
-The Appendix contains results omitted from the main body.
-Section B collects some important facts used in our proofs, Section C provides some technical results, Section D defines some notions used in the analysis, Sections E and F provide proofs for non-convex and strongly convex costs, respectively, Section G provides numerical results, Section H derives the transient times, Section I provides a discussion on the novel results needed for achieving linear speed-up, Section J discusses the extension to time-varying networks, while Section K contains further discussion on dimension dependence in our bounds.
-Appendix B Useful Inequalities
+Appendix contains results omitted from the main body.
+Section B discusses the differences between MSE and HP guarantees, Section C collects some important facts used in our proofs, Section D provides some technical results, Section E defines some notions used in the analysis, Sections F and G provide proofs for non-convex and strongly convex costs, respectively, Section H provides numerical results, Section I derives the transient times, Section J provides a discussion on the novel results needed for achieving linear speed-up, Section K discusses the extension to time-varying networks, while Section L contains further discussion on dimension dependence in our bounds and the notion of sub-Gaussianity used in our work.
+Appendix B On MSE and HP Results
+Although both HP and MSE results establish convergence, the nature of these guarantees is very different.
+In particular, MSE results quantify the average behaviour across many runs of an algorithm, while HP results quantify the behaviour of an individual run.
+As mentioned briefly in the introduction, this distinction is very important in huge-scale applications like LLM training, where it is often impossible to perform more than a single training run, both resource- and time-wise.
+For example, [54] estimate that the GPT-3 model with 175 billion parameters was trained for 34 days, while the variant with 1 trillion parameters can be trained for approximately 84 days.
+This is further exacerbated by phenomena like heavy-tailed noise, frequently observed during training of deep learning models and transformers, e.g., [89, 72], which can cause the performance of an individual run to significantly deviate from the average performance.
+Hence, strong guarantees with respect to an individual run are very important in such applications.
+While MSE results can be used to provide guarantees with respect to an individual run, they can be quite loose compared to HP bounds.
+In particular, if we have a MSE bound of the form $\mathbb{E}[X_{t}^{2}]=1/t$, Chebyshev’s inequality implies $\mathbb{P}\big(X_{t}^{2}>\nicefrac{{1}}{{\delta t}}\big)\leq\delta$, which shows an inversely proportional dependence on the confidence level $\delta\in(0,1)$.
+On the other hand, HP results establish guarantees of the type $\mathbb{P}(X_{t}^{2}>\log(\nicefrac{{1}}{{\delta}})/t)\leq\delta$, having a much milder, logarithmic dependence on $1/\delta$.
+Equivalently, for any threshold $\epsilon>0$, MSE + Chebyshev implies $\mathbb{P}(X_{t}^{2}>\epsilon)\leq\nicefrac{{1}}{{t\epsilon}}$, while HP gives $\mathbb{P}(X_{t}^{2}>\epsilon)\leq\exp(-t\epsilon)$, providing a much sharper bound on the tail probability.
+Since our goal is to show that $X_{t}^{2}\leq\epsilon$ (e.g., if $X_{t}=\|x^{t}-x^{\star}\|$, or if $X_{t}=\min_{k\in[t]}\|\nabla f(x^{k})\|$), then $\mathbb{P}(X_{t}^{2}>\epsilon)$ represents the probability of failure, with MSE + Chebyshev implying that the probability of failure decays polynomially in $t$, while HP bound implies that the probability of failure decays exponentially fast in $t$, giving much stronger guarantees on the convergence of an individual run.
+Appendix C Useful Inequalities
 In this section we outline some well-known inequalities and results used in our proofs. We start with Jensen’s inequality for convex/concave functions.
@@
     $\|\nabla f(x)\|^{2}\leq 2L(f(x)-f^{\star})$.
-Finally, we complete this section with the following important consequences of (A6), see, e.g., [55].
+Next, we state some important consequences of (A6), see, e.g., [59].
 Proposition 8.
@@
     $f(x)-f^{\star}\geq\frac{\mu}{2}\|x-x^{\star}\|^{2}$.
-Appendix C Technical Results
+Finally, we provide a useful martingale concentration result from [29].
+Proposition 9.
+Let $X_{1},\ldots,X_{n}\in\mathbb{R}^{d}$ be random vectors and denote by $\mathcal{F}_{n}=\sigma\big(\{X_{1},\ldots,X_{n}\}\big)$ the corresponding filtration.
+If for all $i\in[n]$, the vector $X_{i}$ satisfies $\mathbb{E}[X_{i}\>|\>\mathcal{F}_{i-1}]=0$ and
+$\mathbb{P}(\|X_{i}\|>\epsilon\>|\>\mathcal{F}_{i-1})\leq 2\exp\bigg(-\frac{t^{2}}{2\sigma_{i}^{2}}\bigg),$
+for some $\mathcal{F}_{i-1}$-measurable $\sigma_{i}>0$ and all $\epsilon>0$, then for any $\delta\in(0,1)$, we have
+$\mathbb{P}\Big(\|\sum_{i\in[n]}X_{i}\|^{2}>2\sqrt{2}\sum_{i\in[n]}\sigma_{i}^{2}\log(\nicefrac{{2d}}{{\delta}})\Big)\leq\delta.$
+Appendix D Technical Results
 In this section we prove Lemmas 2 and 5. We also provide another technical result used in our proofs. For the reader’s convenience, we restate Lemmas 2 and 5 below.
@@
   3. 3.
-    The average noise $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian, i.e., $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{120d\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq\exp(1)$.
+    The average noise $\overline{z}^{t}$ satisfies $\mathbb{E}\left[\exp\bigg(\frac{n\|\overline{z}^{t}\|^{2}}{15\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\right]\leq 2d\exp(1)$.
 Proof.
@@
   3. 3.
-    Combining the previous result with Lemma 1 from [27], it can be readily seen that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{3d}}{\sqrt{n}}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$, we have
-    $\mathbb{P}\big(\|\overline{z}^{t}\|>\epsilon\big)\leq 2\exp\bigg(-\frac{n\epsilon^{2}}{24\sigma^{2}d}\bigg).$
-    Using the equivalence between different conditions for scalar sub-Gaussian random variables, see, e.g., Proposition 2.5.2 in [67], with $\|\overline{z}^{t}\|$ being the variable of interest, it can be readily verified that $\mathbb{E}\Big[\exp\Big(\frac{\|\overline{z}^{t}\|^{2}}{K^{2}}\Big)\Big]\leq\exp(1)$, where $K\leq\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$, implying that $\overline{z}^{t}$ is $\frac{2\sigma\sqrt{30d}}{\sqrt{n}}$-sub-Gaussian.
+    From (A4) we know that conditioned on $\mathcal{F}_{t}$, the vectors $z_{1}^{t},\ldots,z_{n}^{t}$ are independent and zero-mean. Moreover, for each $i\in[n]$ and any $\epsilon>0$, we have
+    $\mathbb{P}\big(\|z^{t}_{i}\|>$ $\epsilon\>|\>\mathcal{F}_{t}\big)=\mathbb{P}\bigg(\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{2\sigma_{i}^{2}}\bigg)>\exp\bigg(\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\>\Big|\>\mathcal{F}_{t}\bigg)\stackrel{{\scriptstyle(a)}}{{\leq}}\exp\bigg(-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{2\sigma_{i}^{2}}\bigg)\>\Big|\>\mathcal{F}_{t}\bigg]$
+    $\stackrel{{\scriptstyle(b)}}{{\leq}}\exp\bigg(-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\Big|\>\mathcal{F}_{t}\bigg]\Bigg)^{1/2}\stackrel{{\scriptstyle(c)}}{{\leq}}\exp\bigg(\frac{1}{2}-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg)\stackrel{{\scriptstyle(d)}}{{\leq}}2\exp\bigg(-\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg),$
+    where $(a)$ follows from Markov’s inequality, $(b)$ follows from Proposition 1, in $(c)$ we use (A4), while $(d)$ follows form the fact that $\exp(1/2)\leq 2$.
+Therefore, vectors $z_{1}^{t},\ldots,z_{n}^{t}$ satisfy the conditions of Proposition 9 with respect to $\mathcal{F}_{t}$, and it follows that, for any $\delta\in(0,1)$
+    $\mathbb{P}\bigg(\|\overline{z}^{t}\|^{2}>\frac{2\sqrt{2}\sigma^{2}\log(\nicefrac{{2d}}{{\delta}})}{n}\bigg)\leq\delta,$
+    or equivalently, for any $\epsilon>0$, we have
+    $\mathbb{P}(\|\overline{z}^{t}\|^{2}>\epsilon)\leq 2d\exp\bigg(-\frac{n\epsilon}{2\sigma^{2}\sqrt{2}}\bigg)\leq 2d\exp\bigg(-\frac{n\epsilon}{3\sigma^{2}}\bigg).$ (8)
+    Next, define $Y=\frac{\sqrt{n}\|\overline{z}^{t}\|}{\sigma\sqrt{3}}$. Using the layer-cake formula for the expectation of a non-negative random variable, we then have, for any integer $p\geq 1$
+    $\mathbb{E}\big[\|Y\|^{2p}\big]$ $=\int_{0}^{\infty}\mathbb{P}\big(\|Y\|^{2p}>y\big)dy\stackrel{{\scriptstyle(a)}}{{=}}p\int_{0}^{\infty}s^{p-1}\mathbb{P}\big(\|Y\|^{2}>s\big)ds$
+    $=p\int_{0}^{\infty}s^{p-1}\mathbb{P}\bigg(\|\overline{z}^{t}\|^{2}>\frac{3\sigma^{2}s}{n}\bigg)ds\stackrel{{\scriptstyle(b)}}{{\leq}}2dp\int_{0}^{\infty}s^{p-1}\exp(-s)ds\stackrel{{\scriptstyle(c)}}{{\leq}}2dp^{p+1},$ (9)
+    where $(a)$ follows by introducing the substitution $y=s^{p}$, $(b)$ follows from (8), while $(c)$ follows from the definition of $\Gamma$ function and the fact that $\Gamma(p)\leq p^{p}$.
+For any $c>0$, we then have
+    $\mathbb{E}\bigg[\exp\bigg(\frac{cn\|\overline{z}^{t}\|^{2}}{3\sigma^{2}}\bigg)\bigg]$ $=1+\sum_{k=1}^{\infty}\frac{c^{k}\mathbb{E}\big[\|Y\|^{2k}\big]}{k!}\stackrel{{\scriptstyle(i)}}{{\leq}}1+\sum_{k=1}^{\infty}\frac{2dk(ck)^{k}}{k!}\stackrel{{\scriptstyle(ii)}}{{\leq}}2d\bigg(1+\frac{1}{e}\sum_{k=1}^{\infty}k(ec)^{k}\bigg),$
+    where $(i)$ follows from (3), while $(ii)$ follows from $k!\geq e(k/e)^{k}$.
+Next, note that the sequence $\sum_{k=1}^{\infty}kp^{k}$ converges if $p<1$, with $\sum_{k=1}^{\infty}kp^{k}=\frac{p}{(1-p)^{2}}$, therefore, choosing $c<\frac{1}{e}$ in the relation above, we get
+    $\mathbb{E}\bigg[\exp\bigg(\frac{cn\|\overline{z}^{t}\|^{2}}{3\sigma^{2}}\bigg)\bigg]\leq 2d\bigg(1+\frac{c}{(1-ec)^{2}}\bigg).$ (10)
+    It can be readily verified that $\frac{c}{(1-ec)^{2}}\leq e-1$, for any $c\leq\frac{e-1}{e^{2}}$. Choosing $c=\frac{1}{5}$ and plugging into (10) completes the proof.
 We next restate Lemma 5.
@@
 $=b_{t}Y^{t}+\sum_{i\in[M]}\frac{C_{i}}{(t+t_{0})^{i}},$ (12)
-where the second inequality follows from the fact that $b_{t}\in(0,1)$ and Proposition 1. Unrolling the recursion (C) and noting that $Y_{1}=X_{1}$, since $X_{1}>0$ is deterministic, we get
+where the second inequality follows from the fact that $b_{t}\in(0,1)$ and Proposition 1. Unrolling the recursion (D) and noting that $Y_{1}=X_{1}$, since $X_{1}>0$ is deterministic, we get
 $Y^{t+1}$ $\leq X^{1}\prod_{k\in[t]}b_{k}+\sum_{i\in[M]}C_{i}\sum_{k\in[t]}\frac{1}{(k+t_{0})^{i}}\prod_{s=k+1}^{t}b_{s}$
@@
 $\sum_{k\in[t]}(k+t_{0})^{a-i}\leq\begin{cases}\frac{(t+1+t_{0})^{a}}{a},&i=1\\ \frac{(t+1+t_{0})^{a-1}}{a-1},&i=2\\ \ln(t+t_{0}+1),&i=3\\ \frac{t_{0}^{3-i}}{i-3},&i\geq 4\end{cases}.$ (14)
-Plugging (11) into (C), we get
+Plugging (14) into (D), we get
 $Y^{t+1}$ $\leq\frac{(t_{0}+1)^{a}X^{1}}{(t+1+t_{0})^{a}}+\frac{2^{a}C_{1}}{a}+\frac{2^{a}C_{2}}{(a-1)(t+t_{0}+1)}$
@@
 (16)
-Plugging (13) into (C) completes the proof. ∎
-Appendix D Analysis Setup
+Plugging (16) into (D) completes the proof. ∎
+Appendix E Analysis Setup
 In this section we define some notation useful for the analysis.
@@
 Note that $\|\nabla\mathbf{f}^{\star}\|^{2}=\sum_{i\in[n]}\|\nabla f_{i}(x^{\star})\|^{2}$ is a useful measure of heterogeneity.
-Appendix E Proofs for Non-convex Costs
+Appendix F Proofs for Non-convex Costs
 In this section we prove Lemmas 1 and 3, as well as Theorem 1. For the reader’s convenience, we restate the results below, starting with Lemma 1.
@@
 $\|\overline{\nabla}f_{t}-\nabla f(\overline{x}^{t})\|^{2}=\Big\|\frac{1}{n}\sum_{i\in[n]}\big[\nabla f_{i}(x^{t}_{i})-\nabla f_{i}(\overline{x}^{t})\big]\Big\|^{2}\leq\frac{L^{2}}{n}\sum_{i\in[n]}\|x^{t}_{i}-\overline{x}^{t}\|^{2},$ (23)
-where we used Proposition 1 and (A3) in the last inequality. Plugging (20) in (E) gives the desired result. ∎
+where we used Proposition 1 and (A3) in the last inequality. Plugging (23) in (F) gives the desired result. ∎
 Next, we restate Lemma 3.
@@
 We are now ready to prove Theorem 1, which we restate next, for convenience.
-Theorem 3.4.
+Theorem 1.
 Let (A1)-(A5) hold.
-If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$, where $C>0$ is a problem related constant satisfying
-$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
+If for any $T\geq 1$, the step-size is chosen such that $\alpha_{T}\equiv\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{15LT}}\right\}$, where $C>0$ is a problem related constant satisfying
+$C\leq\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\},$
 we then have, for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\sqrt{nT}}$
-$+\frac{\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})}{CT}+\frac{L^{2}\Delta_{x}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(A^{2}+\sigma^{2})}{\sigma^{2}d(1-\lambda)^{2}T}\Bigg).$
+$\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{L}\big(\Delta_{f}+\log(\nicefrac{{2d}}{{\delta}})+1\big)}{\sqrt{nT}}$
+$+\frac{\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}T}\Bigg).$
 Proof.
@@
 $\leq\frac{1}{1-\lambda}\sum_{t\in[T]}\alpha_{t}^{3}\Big[\|z^{t}_{i}\|^{2}+A^{2}+B^{2}\|\nabla f(x^{t}_{i})\|^{2}\Big].$ (29)
-Plugging (E) into (25), we get
+Plugging (F) into (28), we get
 $\frac{1}{8n}$ $\sum_{t\in[T]}\sum_{i\in[n]}\alpha_{t}\|\nabla f(x^{t}_{i})\|^{2}\leq\Delta_{f}-\sum_{t\in[T]}\alpha_{t}\Big(\langle\nabla f(\overline{x}^{t}),\overline{z}^{t}\rangle+\frac{9\alpha_{t}\sigma^{2}}{4n}\|\nabla f(\overline{x}^{t})\|^{2}\Big)+L\sum_{t\in[T]}\alpha_{t}^{2}\|\overline{z}^{t}\|^{2}$
@@
 $\mathbb{E}[\exp(B_{2,T})]$ $=\mathbb{E}\left[\exp(B_{2,T-1})\mathbb{E}\left[\exp\left(3\alpha_{T}^{2}L\|\overline{z}^{T}\|^{2}\right)\>|\>\mathcal{F}_{T}\right]\right]$
-$=\mathbb{E}\left[\exp(B_{2,T-1})\mathbb{E}\left[\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\frac{n\|\overline{z}^{T}\|^{2}}{120\sigma^{2}d}\right)\>\Big|\>\mathcal{F}_{T}\right]\right]$
-$\leq\mathbb{E}\left[\exp(B_{2,T-1})\left(\mathbb{E}\left[\exp\left(\frac{n\|\overline{z}^{T}\|^{2}}{120\sigma^{2}d}\right)\>\Big|\>\mathcal{F}_{T}\right]\right)^{\nicefrac{{360\alpha_{T}^{2}\sigma^{2}dL}}{{n}}}\right],$
-where the last inequality follows by choosing $\alpha_{t}\leq\frac{\sqrt{n}}{6\sigma\sqrt{10dL}}$ and using Proposition 1. Using Lemma 2, we get
-$\mathbb{E}[\exp(B_{2,T})]\leq\mathbb{E}\left[\exp(B_{2,T-1})\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\right)\right]=\exp\left(\frac{360\alpha_{T}^{2}\sigma^{2}dL}{n}\right)\mathbb{E}[\exp(B_{2,T-1})].$
-Unrolling the recursion, it follows that $\mathbb{E}[\exp(B_{2,T})]\leq\exp\left(\frac{3\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}\right)$.
+$=\mathbb{E}\left[\exp(B_{2,T-1})\mathbb{E}\left[\exp\left(\frac{45\alpha_{T}^{2}\sigma^{2}L}{n}\times\frac{n\|\overline{z}^{T}\|^{2}}{15\sigma^{2}}\right)\>\Big|\>\mathcal{F}_{T}\right]\right]$
+$\leq\mathbb{E}\left[\exp(B_{2,T-1})\left(\mathbb{E}\left[\exp\left(\frac{n\|\overline{z}^{T}\|^{2}}{15\sigma^{2}}\right)\>\Big|\>\mathcal{F}_{T}\right]\right)^{\nicefrac{{45\alpha_{T}^{2}\sigma^{2}L}}{{n}}}\right],$
+where the last inequality follows by choosing $\alpha_{t}\leq\frac{\sqrt{n}}{3\sigma\sqrt{5L}}$ and using Proposition 1. Using the third property in Lemma 2, we get
+$\mathbb{E}[\exp(B_{2,T})]\leq\exp\left(\frac{45\alpha_{T}^{2}\sigma^{2}L}{n}\big(1+\log(2d)\big)\right)\mathbb{E}\left[\exp(B_{2,T-1})\right].$
+Unrolling the recursion, it follows that $\mathbb{E}[\exp(B_{2,T})]\leq\exp\left(\frac{45\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}\right)$.
 Finally, to bound $\mathbb{E}[\exp(B_{3,T})]$, we can proceed in the same way, using the conditional independence of noise across agents (recall (A4)) to note that, if $\alpha_{t}\leq\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2/3}L^{2/3}\sqrt[3]{9}}$, then $\mathbb{E}[\exp(B_{3,T})]\leq\exp\left(\frac{9\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\right)$.
 Combining everything, we get
-$\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg).$ (29)
+$\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg).$ (32)
 Using Markov’s inequality and (32), we get, for any $\epsilon>0$
-$\mathbb{P}\left(M_{T}>\epsilon\right)\leq\exp(-\epsilon)\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(-\epsilon+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg),$
+$\mathbb{P}\left(M_{T}>\epsilon\right)\leq\exp(-\epsilon)\mathbb{E}[\exp(M_{T})]\leq\exp\bigg(-\epsilon+\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}\bigg),$
 or equivalently, for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$M_{T}\leq\log(\nicefrac{{1}}{{\delta}})+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
-Recalling the definition of $M_{T}$ and using the fact that the sequence of step-sizes in non-increasing we then have, with probability at least $1-\delta$
+$M_{T}\leq\log(\nicefrac{{1}}{{\delta}})+\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
+Recalling the definition of $M_{T}$ and using the fact that the sequence of step-sizes is non-increasing we then have, with probability at least $1-\delta$
 $\frac{\alpha_{T}}{16n}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}$ $\leq\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+\frac{3\Delta_{x}L^{2}}{2}\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}$
-$+\frac{120\sigma^{2}dL}{n}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
+$+\frac{15\sigma^{2}L}{n}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{3\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
 Dividing both sides by $\frac{\alpha_{T}T}{16}$, it follows that, with probability at least $1-\delta$
 $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}$ $\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha_{T}T}+\frac{24\Delta_{x}L^{2}}{\alpha_{T}T}\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}$
-$+\frac{1920\sigma^{2}dL}{n\alpha_{T}T}\sum_{t\in[T]}\alpha_{t}^{2}+\frac{48\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\alpha_{T}T(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
+$+\frac{240\sigma^{2}L}{n\alpha_{T}T}\big(1+\log(2d)\big)\sum_{t\in[T]}\alpha_{t}^{2}+\frac{48\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{\alpha_{T}T(1-\lambda)^{2}}\sum_{t\in[T]}\alpha_{t}^{3}.$
 Next, consider two regimes with respect to the time horizon $T$.
@@
 Noticing that $\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}\leq\frac{\alpha}{1-\lambda^{2}}$, we then have, with probability at least $1-\delta$
-    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\alpha\sigma^{2}dL}{n}+\frac{48\alpha^{2}\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}.$
-    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{dLT}}$, then with probability at least $1-\delta$
-    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{1920\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
-    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{dLT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{dLT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{dLT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma\sqrt{dL}}{\sqrt{nT}}+\frac{1}{CT}$.
+    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}$ $\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}$
+    $+\frac{240\alpha\sigma^{2}L\big(1+\log(2d)\big)}{n}+\frac{48\alpha^{2}\lambda^{2}L^{2}(\sigma^{2}+A^{2})}{(1-\lambda)^{2}}.$
+    If the step-size satisfies $\alpha\leq\frac{\sqrt{n}}{\sigma\sqrt{15LT}}$, then with probability at least $1-\delta$
+    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}$ $\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{\alpha T}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}$
+    $+\frac{16\sigma\sqrt{15L}\big(1+\log(2d)\big)}{\sqrt{nT}}+\frac{4n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}T}.$
+    Setting $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$ and $\alpha=\min\left\{C,\frac{\sqrt{n}}{\sigma\sqrt{15LT}}\right\}$ guarantees that all the step-size conditions are satisfied and it readily follows that $\frac{1}{\alpha}=\max\left\{\frac{1}{C},\frac{\sigma\sqrt{15LT}}{\sqrt{n}}\right\}\leq\frac{1}{C}+\frac{\sigma\sqrt{15LT}}{\sqrt{n}}$, implying that $\frac{1}{\alpha T}\leq\frac{\sigma\sqrt{15L}}{\sqrt{nT}}+\frac{1}{CT}$.
 Therefore, for any $\delta\in(0,1)$, with probability at least $1-\delta$, we finally have
-    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})+120\big)}{\sqrt{nT}}$
-    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{48n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}T}.$
+    $\frac{1}{nT}\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\>\leq\>\frac{16\sigma\sqrt{15L}\big(\Delta_{f}+\log(\nicefrac{{2d}}{{\delta}})+1\big)}{\sqrt{nT}}$
+    $+\frac{16\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{CT}+\frac{24\Delta_{x}L^{2}}{(1-\lambda^{2})T}+\frac{4n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}T}.$
   2. 2.
     Unknown time horizon.
-In this case, we choose a time-varying step-size $\alpha_{t}=\frac{C^{\prime}}{\sqrt{t+1}}$, for all $t\geq 1$ and $C^{\prime}=\sqrt{2}C$, where $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$, again guaranteeing that all the step-size conditions are satisfied.
+In this case, we choose a time-varying step-size $\alpha_{t}=\frac{C^{\prime}}{\sqrt{t+1}}$, for all $t\geq 1$ and $C^{\prime}=\sqrt{2}C$, where $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$, again guaranteeing that all the step-size conditions are satisfied.
 Noting that $\alpha_{t}\leq C$, we get $\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}\leq\frac{C}{(1-\lambda^{2})}$, hence, with probability at least $1-\delta$
     $\frac{1}{nT}\sum_{t\in[T]}$ $\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}\leq\frac{16\sqrt{2}\big(\Delta_{f}+\log(\nicefrac{{1}}{{\delta}})\big)}{C\sqrt{T+1}}+\frac{24\sqrt{2}\Delta_{x}L^{2}}{C\sqrt{T+1}(1-\lambda^{2})}$
-    $+\frac{3840\sqrt{2}\sigma^{2}dLC\log(T+1)}{n\sqrt{T+1}}+\frac{384\lambda^{2}L^{2}(\sigma^{2}+A^{2})C^{2}}{(1-\lambda)^{2}\sqrt{T+1}}.$
-    Note that in the unknown time horizon case we lose the linear speed-up in the number of users, as, even in the case $C=\frac{\sqrt{n}}{6\sigma\sqrt{10dL}}$, we have
-    $\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{T+1}}{{\delta}})\big)}{\sqrt{n(T+1)}}$
-    $+\frac{\sigma\sqrt{d}\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
+    $+\frac{240\sqrt{2}\sigma^{2}LC\log(T+1)\big(1+\log(2d)\big)}{n\sqrt{T+1}}+\frac{384\lambda^{2}L^{2}(\sigma^{2}+A^{2})C^{2}}{(1-\lambda)^{2}\sqrt{T+1}}.$
+    Note that in the unknown time horizon case we lose the linear speed-up in the number of users, as, even in the case $C=\frac{\sqrt{n}}{3\sigma\sqrt{5L}}$, we have
+    $\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{L}\big(\Delta_{f}+\log(\nicefrac{{2dT}}{{\delta}})\big)}{\sqrt{n(T+1)}}$
+    $+\frac{\sigma\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
     as the last term, which does not attain a linear speed-up, is no longer of higher order.
 We note that linear speed-up in the MSE sense is also achieved under a known time horizon and fixed step-size, e.g., [35, 82].
-Appendix F Proofs for Strongly Convex Costs
+Appendix G Proofs for Strongly Convex Costs
 In this section we prove Lemma 4 and Theorem 2.
@@
 where in $(i)$ we used the fact that $(1+\frac{a}{2})(1-a)\leq(1-\frac{a}{2})$ for any $a>0$, while $(ii)$ follows by setting $\nu\leq\frac{\mu}{24a\overline{\sigma}^{2}K}$ and from the step-size choice $\alpha_{t}\leq\frac{1}{\mu}$.
-Plugging (F) in (F), using the shorthand $E_{t}=\alpha_{t}^{2}n\sigma^{2}+9\alpha_{t}\|\nabla\mathbf{f}^{\star}\|^{2}/2\mu$ and taking the full expectation, we get
+Plugging (G) in (G), using the shorthand $E_{t}=\alpha_{t}^{2}n\sigma^{2}+9\alpha_{t}\|\nabla\mathbf{f}^{\star}\|^{2}/2\mu$ and taking the full expectation, we get
 $\mathbb{E}[\exp(\nu$ $K_{t+1}\|{\mathbf{x}^{t+1}}-{\mathbf{x}^{\star}}\|^{2})]\leq\exp(\nu K_{t+1}E_{t})\mathbb{E}[\exp((1-\alpha_{t}\mu)\nu K_{t+1}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|^{2})]$
@@
 where $(i)$ follows from Proposition 3, in $(ii)$ we set $\theta=\frac{\widetilde{\lambda}}{2}$, while $(iii)$ follows from the fact that $\|\nabla\mathbf{f}^{t}\|^{2}\leq 2L^{2}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|+2\|\nabla\mathbf{f}^{\star}\|^{2}$ (recall Proposition 7).
-Choosing $\alpha_{t}\leq\frac{\sqrt{1-\lambda}}{2\overline{\sigma}\lambda\sqrt{6(t+t_{0}+2)K}}$ and plugging (F) in (F), we get
+Choosing $\alpha_{t}\leq\frac{\sqrt{1-\lambda}}{2\overline{\sigma}\lambda\sqrt{6(t+t_{0}+2)K}}$ and plugging (G) in (G), we get
 $\mathbb{E}_{t}[\exp(\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2})]\leq\exp\bigg(\nu K_{t+1}\Big(1+\frac{\widetilde{\lambda}}{4}\Big)\Big[\lambda\Big(1-\frac{\widetilde{\lambda}}{2}\Big)\|\widetilde{{\mathbf{x}}}_{t}\|^{2}$
@@
 where we used Proposition 1 in the first and Lemma 11 in the second inequality, with $D_{t}=\frac{5\alpha_{t}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{t}}{a\mu-1}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{a\mu}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(t+t_{0})^{a\mu}}\Big)$.
-Similarly, we use (F), to get
+Similarly, we use (G), to get
 $\bigg(\mathbb{E}\big[\exp(\nu K_{t}\|\widetilde{{\mathbf{x}}}_{t}\|^{2})\big]\bigg)^{\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})\frac{t+t_{0}+2}{t+t_{0}+1}}$ $\leq\exp(\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})S_{t-1}\nu K_{t+1})\bigg(\mathbb{E}\big[\exp(\nu K_{t-1}\|\widetilde{{\mathbf{x}}}_{t-1}\|^{2})\big]\bigg)^{\lambda^{2}(1-\nicefrac{{\widetilde{\lambda}}}{{4}})^{2}\frac{t+t_{0}+2}{t+t_{0}}}$ (41)
 $\times\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{20\alpha_{t}^{2}\lambda^{2}L^{2}\nu K_{t}}{(1-\lambda)^{2}}\|{\mathbf{x}}_{t-1}-{\mathbf{x}^{\star}}\|^{2}\bigg)\bigg]\Bigg)^{\nicefrac{{\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})\widetilde{\lambda}}}{{4}}}.$
-Plugging (F) and (38) into (F), we get
+Plugging (G) and (41) into (G), we get
 $\mathbb{E}[\exp(\nu K_{t+1}\|\widetilde{{\mathbf{x}}}^{t+1}\|^{2})]$ $\leq\exp\Big(\nu K_{t+1}\big(S_{t}+S_{t-1}\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})\big)+D_{t}+\lambda(1-\nicefrac{{\widetilde{\lambda}}}{{4}})D_{t-1}\big)\Big)$
@@
 We are now ready to prove Theorem 2. Prior to that, we restate it, for convenience.
-Theorem 3.7.
+Theorem 2.
 Let (A1)-(A4) and (A6) hold, let the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$.
-If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
-$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
-$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
-$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
+If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{1960\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
+$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}\kappa\big(1+\log(2d)\big)/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
+$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
 Proof.
@@
 $\stackrel{{\scriptstyle(c)}}{{\leq}}\exp\bigg((1-\alpha_{t}\mu)\frac{t+t_{0}+1}{t+t_{0}}\nu F_{t}+\frac{\alpha_{t}\nu(t+t_{0}+1)L^{2}}{2}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\bigg)$
-$\times\exp\bigg(\frac{3\alpha_{t}^{2}\nu^{2}n(t+t_{0}+1)^{2}\sigma^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{2}+120\alpha_{t}^{2}\nu\sigma^{2}d(t+t_{0}+1)L\bigg)$
+$\times\exp\bigg(\frac{3\alpha_{t}^{2}\nu^{2}n(t+t_{0}+1)^{2}\sigma^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{2}+15\alpha_{t}^{2}\nu\sigma^{2}(t+t_{0}+1)L\big(1+\log(2d)\big)\bigg)$
 $\stackrel{{\scriptstyle(d)}}{{\leq}}\exp\bigg(\nu\Big(b_{t}F_{t}+c_{t}\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}+d_{t}\Big)\bigg),$
-where in $(a)$ we used the fact that $F_{t}$ and $\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}$ are $\mathcal{F}_{t}$-measurable, $(b)$ follows from Proposition 4, in $(c)$ we use Lemma 2, Proposition 1 and impose the condition $\alpha_{t}\leq\frac{1}{4\sigma\sqrt{15(t+t_{0}+1)dL}}$, while $(d)$ follows from Proposition 7 and the definition of $F_{t}$, with $b_{t}=\Big(1-\alpha_{t}\mu+3\alpha_{t}^{2}\nu(t+t_{0}+1)L\Big)\frac{t+t_{0}+1}{t+t_{0}}$, $c_{t}=\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}$ and $d_{t}=120\alpha_{t}^{2}\sigma^{2}d(t+t_{0}+1)L$.
+where in $(a)$ we used the fact that $F_{t}$ and $\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}$ are $\mathcal{F}_{t}$-measurable, $(b)$ follows from Proposition 4, in $(c)$ we use Lemma 2, Proposition 1 and impose the condition $\alpha_{t}\leq\frac{1}{\sigma\sqrt{30(t+t_{0}+1)L}}$, while $(d)$ follows from Proposition 7 and the definition of $F_{t}$, with $b_{t}=\Big(1-\alpha_{t}\mu+3\alpha_{t}^{2}\nu(t+t_{0}+1)L\Big)\frac{t+t_{0}+1}{t+t_{0}}$, $c_{t}=\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}$ and $d_{t}=15\alpha_{t}^{2}\sigma^{2}(t+t_{0}+1)L\big(1+\log(2d)\big)$.
 Taking the full expectation and applying Proposition 4, we get
@@
 $qc_{t}=\bigg(1+\frac{4}{\alpha_{t}\mu}\bigg)\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}=\bigg(\frac{\alpha_{t}}{2}+\frac{2}{\mu}\bigg)(t+t_{0}+1)L^{2}\leq\frac{3L^{2}}{\mu}(t+t_{0}+1),$ (44)
-where the first inequality follows from $\alpha_{t}\leq\frac{1}{\mu}$. Using (40) and (41) in (F), it follows that
+where the first inequality follows from $\alpha_{t}\leq\frac{1}{\mu}$. Using (43) and (44) in (G), it follows that
 $\mathbb{E}\big[\exp(\nu F_{t+1})\big]$ $\leq\exp\big(d_{t}\nu\big)\sqrt[p]{\big(\mathbb{E}\big[\exp(\nu F_{t})\big]\big)^{pb_{t}}}\sqrt[q]{\mathbb{E}\big[\exp\big(\nu(t+t_{0}+1)3\kappa L\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}\big)\big]}$
@@
 where $N_{t}\coloneqq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}$.
-Define $G_{1}\coloneqq\frac{8640\kappa\sigma^{2}d}{\mu}$, $G_{2}\coloneqq\frac{1296\kappa^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)}$, $G_{4}\coloneqq\frac{6480\kappa^{3}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and $G_{3}\coloneqq\frac{373248n\kappa^{4}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}}$, and plug into (F), to get
+Define $G_{1}\coloneqq\frac{1080\kappa\sigma^{2}(1+\log(2d))}{\mu}$, $G_{2}\coloneqq\frac{1296\kappa^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)}$, $G_{4}\coloneqq\frac{6480\kappa^{3}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and $G_{3}\coloneqq\frac{373248n\kappa^{4}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}}$, and plug into (G), to get
 $\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]\leq\Big(\mathbb{E}\big[\exp\big(\nu F_{t}\big)\big]\Big)^{b_{t}}\exp\bigg(\sum_{i\in[3]}\frac{\nu G_{i}}{(t+t_{0})^{i}}+\frac{\nu G_{4}}{(t+t_{0})^{8}}\bigg).$
 Recalling the definition of $b_{t}$ and (43), it follows that $b_{t}\leq 1-\frac{a\mu/2-1}{t+t_{0}}=1-\frac{2}{t+t_{0}}$, therefore we can bound the MGF of $\nu F_{t+1}$ using Lemma 5 with $a=2$, $M=8$, $C_{i}=G_{i}$ for $i\in[3]$, $C_{8}=G_{4}$ and $C_{j}=0$, for $j\in\{4,\ldots,7\}$, to finally get
-$\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]$ $\leq\exp\bigg(\frac{(t_{0}+2)^{3}\nu\Delta_{f}}{(t+1+t_{0})^{2}}+4\nu G_{1}+\frac{4\nu G_{2}}{t+1+t_{0}}\bigg)$
+$\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]$ $\leq\exp\bigg(\frac{n(t_{0}+1)^{3}\nu\Delta_{f}}{(t+1+t_{0})^{2}}+4\nu G_{1}+\frac{4\nu G_{2}}{t+1+t_{0}}\bigg)$
 $\times\exp\bigg(\frac{4\nu G_{3}\log(t+t_{0}+1)}{(t+t_{0}+1)^{2}}+\frac{4\nu G_{4}}{5(t_{0}+1)^{5}(t+1+t_{0})^{2}}\bigg).$ (46)
@@
 $\mathbb{P}\left(f(\overline{x}^{t+1})-f^{\star}>\epsilon\right)=\mathbb{P}\big(\nu F_{t+1}>\nu n(t+1+t_{0})\epsilon\big)\leq\exp\left(-\nu n(t+1+t_{0})\epsilon\right)\mathbb{E}\big[\big(\nu F_{t+1}\big)\big].$
-Using (F), it can be readily verified that choosing
-$\epsilon_{t}^{1}$ $=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})+4G_{1}}{n(t+t_{0})}+\frac{4G_{2}}{n(t+t_{0})^{2}}+\frac{(t_{0}+2)^{3}\Delta_{f}+4G_{3}\log(t+t_{0})+4G_{4}/5(t_{0}+1)^{5}}{n(t+t_{0})^{3}},$ (44)
+Using (G), it can be readily verified that choosing
+$\epsilon_{t}^{1}$ $=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})+4G_{1}}{n(t+t_{0})}+\frac{4G_{2}}{n(t+t_{0})^{2}}+\frac{(t_{0}+1)^{3}\Delta_{f}}{(t+t_{0})^{3}}+\frac{4G_{3}\log(t+t_{0})+4G_{4}/5(t_{0}+1)^{5}}{n(t+t_{0})^{3}},$ (47)
 for any $\delta\in(0,1)$, results in $\mathbb{P}\left(f(\overline{x}^{t})-f^{\star}>\epsilon_{t}^{1}\right)\leq\delta$. Next, using Proposition 7 with $x=x^{t}_{i}$ and $y=\overline{x}^{t}$, we get
@@
 Therefore, we get, for any $\delta\in(0,1)$, with probability at least $1-\delta$
-$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
-$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
-$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
+$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}\kappa\big(1+\log(2d)\big)/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
+$+\frac{(1+t_{0})^{3}\Delta_{f}}{(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
 Finally, it can be verified that the conditions on $a$, $t_{0}$ and $\nu$ in the statement of the theorem ensure that all the step-size conditions are satisfied, completing the proof. ∎
@@
 For simplicity, we omit this analysis.
-Appendix G Numerical Experiments
-In this section we provide some numerical results, which serve to verify our theory.
+Appendix H Numerical Experiments
+In this section we perform numerical experiments, with the aim of validating our theory.
 Methodology.
-We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive semi-definite, making each $f_{i}$ strongly convex, with unbounded gradients.
+We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive definite, making each $f_{i}$ strongly convex, with unbounded gradients.
 We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [80].
@@
 We can see that DSGD consistently achieves faster exponential tail decay for larger networks, across all values of threshold $\varepsilon$, illustrating the effect of the linear speed-up in the HP sense.
-Appendix H Deriving the Transient Time
+Appendix I Deriving the Transient Time
 In this section we provide details on the transient time resulting from our HP bounds.
@@
 This matches the transient time of DSGD obtained from the MSE rate in [35], see, e.g., Table I in [1].
-Next, consider strongly convex costs. Recalling the expressions in (44) and (F), it follows that the full bound for strongly convex costs is of the form
-$\mathcal{O}$ $\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}+\frac{t_{0}^{3}}{n(T+t_{0})^{3}}$
+Next, consider strongly convex costs. Recalling the expressions in (47) and (G), it follows that the full bound for strongly convex costs is of the form
+$\mathcal{O}$ $\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}+\frac{t_{0}^{3}}{(T+t_{0})^{3}}$
 $+\frac{\log(T+t_{0})}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{t_{0}}{n(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{t_{0}^{6}}{n(1-\lambda)^{2}(T+t_{0})^{8}}\bigg),$ (52)
 where $t_{0}$ depends on the network connectivity via the expression $t_{0}=\Omega\Big(\max\Big\{\frac{\lambda}{1-\lambda},\frac{\lambda^{2}}{1-\lambda},\frac{1}{1-\lambda}\Big\}\Big)$.
-Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{1}{1-\lambda}\Big)$, so that the bound in (H) becomes
-$\quad\quad\mathcal{O}\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}$
-$+\frac{\log(T+t_{0})}{(1-\lambda)^{2}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{3}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{8}(T+t_{0})^{8}}\bigg).$ (50)
-It can be seen that, for any $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (H) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$.
+Since we are again interested in the worst-case transient time, it follows that $t_{0}=\Omega\Big(\frac{1}{1-\lambda}\Big)$, so that the bound in (I) becomes
+$\quad\quad\mathcal{O}\bigg(\frac{1}{n(T+t_{0})}+\frac{1}{(1-\lambda)(T+t_{0})^{2}}+\frac{1}{n(1-\lambda)^{2}(T+t_{0})^{2}}+\frac{\log(T+t_{0})}{(1-\lambda)^{2}(T+t_{0})^{3}}$
+$\quad\quad+\frac{1}{(1-\lambda)^{3}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{3}(T+t_{0})^{3}}+\frac{1}{n(1-\lambda)^{8}(T+t_{0})^{8}}\bigg).$ (53)
+It can be seen that for $T\geq\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}},\frac{\sqrt{n}}{1-\lambda},\frac{1}{(1-\lambda)^{\frac{3}{2}}},\frac{\sqrt{n}}{(1-\lambda)^{\frac{3}{2}}},\frac{1}{(1-\lambda)^{\frac{8}{7}}}\Big\}=\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}$, the bound in (I) becomes $\mathcal{O}\Big(\frac{\log(T+t_{0})}{n(T+t_{0})}\Big)$, implying a transient time of at most $\mathcal{O}\Big(\max\Big\{\frac{n}{1-\lambda},\frac{1}{(1-\lambda)^{2}}\Big\}\Big)$.
 Interestingly, this is strictly sharper than the transient time of DSGD obtained from the MSE rate in [35], which is of the order $\mathcal{O}\Big(\frac{n}{(1-\lambda)^{2}}\Big)$, see, e.g., Table II in [1].
 As such, our HP analysis provides either matching or strictly sharper transient times for DSGD compared to the transient times stemming from the corresponding MSE results, further highlighting the tightness of our analysis.
-Appendix I Achieving Linear Speed-up
+Appendix J Achieving Linear Speed-up
 In this section we provide further discussions on the main results and techniques needed to obtain linear speed-up, for both non-convex and strongly convex costs.
@@
 Works like [47, 48, 83] provide no results on the variance reduction benefit of decentralized learning in the HP sense, instead relying only on the definition of sub-Gaussianity.
+[^12] We note that [48] consider the problem of online noncooperative games and show their HP convergence guarantees in terms of dynamic regret $R_{i}(t)$, for each individual $i\in[n]$.
+While this is different from the metrics considered in our work (i.e., user-averaged gradient norm/optimality gap), the analysis in [48] suffers from the same shortcomings as the decentralized works, i.e., no result showing variance reduction in the HP sense is provided, making linear speed-up impossible to achieve even if an average of $R_{i}(t)$’s is considered.
 As such, even if a fixed step-size was used, the authors would be unable to obtain a bound of the form (55) and achieve linear speed-up.
-As discussed in Appendix E, using a time-varying step-size for non-convex costs inevitably leads to a loss of linear speed-up, as all terms decay at the same rate, leading to a bound of the form
+As discussed in Appendix F, using a time-varying step-size for non-convex costs inevitably leads to a loss of linear speed-up, as all terms decay at the same rate, leading to a bound of the form
 $\mathcal{O}\bigg(\frac{1}{\sqrt{nT}}+\frac{n}{\sqrt{T}}\bigg).$
@@
 As such, Lemma 5 provides the most fine-grained control of the MGF and is of independent interest when studying “almost decreasing” type processes, even in centralized settings.
-Appendix J Extending to Time-varying Networks
+Appendix K Extending to Time-varying Networks
 As discussed in the main body, to model the communication network, we use assumption (A1), which subsumes fixed undirected networks, as well as a class of fixed directed networks.
@@
 As such, our results can be easily extended to the (directed) time-varying network setting.
-Appendix K On Dimension Dependence
-As mentioned in Section 4 in the main body, our rates in Theorems 1 and 2 exhibit a dependence on the problem dimension, of order $\sqrt{d}$ for non-convex and $d$ for strongly convex costs, with the dependence stemming from Lemma 2, where it is shown that $\overline{z}^{t}$ is $\mathcal{O}\big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\big)$-sub-Gaussian.
-This is a consequence of working with random vectors, where we simultaneoulsy need to show a bound on the MGF of the inner product $\langle\overline{z}^{t},v\rangle$ for any $\mathcal{F}_{t}$-measurable vector $v\in\mathbb{R}^{d}$ (shown in point 2 of Lemma 2), as well as on the MGF of the norm-squared $\|\overline{z}^{t}\|^{2}$.
+Appendix L On Sub-Gaussian Vectors and Dimension Dependence
+As mentioned in Section 4 in the main body, our rates in Theorems 1 and 2 exhibit a mild dependence on the problem dimension, of order $\log(d)$, with the dependence stemming from Lemma 2.
+This is a consequence of working with random vectors, where we simultaneoulsy need to show a bound on the MGF of the inner product $\langle\overline{z}^{t},v\rangle$ for any $\mathcal{F}_{t}$-measurable vector $v\in\mathbb{R}^{d}$ (shown in point 2 of Lemma 2), as well as on the MGF of the squared norm $\|\overline{z}^{t}\|^{2}$ (shown in point 3 of Lemma 2).
 While the inner product maintains some desirable properties, such as being zero-mean and linear (in the sense that $\langle\overline{z}^{t},v\rangle=\frac{1}{n}\sum_{i\in[n]}\langle z^{t}_{i},v\rangle$), this is not the case with the squared norm, which is neither zero-mean, nor linear.
@@
 $\mathbb{E}\bigg[\exp\bigg(\frac{\|\overline{z}^{t}\|^{2}}{\sigma^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]$ $\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{E}\Bigg[\exp\Bigg(\frac{\big(\sum_{i\in[n]}\|z^{t}_{i}\|\big)^{2}}{n\sum_{i\in[n]}\sigma_{i}^{2}}\Bigg)\>\Big|\>\mathcal{F}_{t}\Bigg]\stackrel{{\scriptstyle(ii)}}{{\leq}}\mathbb{E}\bigg[\exp\bigg(\frac{1}{n}\sum_{i\in[n]}\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]$
-$\stackrel{{\scriptstyle(iii)}}{{\leq}}\prod_{i\in[n]}\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]\Bigg)^{1/n}\stackrel{{\scriptstyle(iv)}}{{\leq}}\prod_{i\in[n]}\exp\bigg(\frac{1}{n}\bigg)=\exp(1),$
+$\stackrel{{\scriptstyle(iii)}}{{\leq}}\prod_{i\in[n]}\Bigg(\mathbb{E}\bigg[\exp\bigg(\frac{\|z^{t}_{i}\|^{2}}{\sigma_{i}^{2}}\bigg)\>\big|\>\mathcal{F}_{t}\bigg]\Bigg)^{1/n}\stackrel{{\scriptstyle(iv)}}{{\leq}}\prod_{i\in[n]}\exp\bigg(\frac{1}{n}\bigg)=\exp(1),$ (59)
 where $(i)$ follows from Proposition 1, $(ii)$ follows from Sedrakyan’s inequality, namely that
@@
 Therefore, a direct approach yields that $\overline{z}^{t}$ is $\sigma$-sub-Gaussian, failing to show the variance reduction benefit.
-To circumvent this issue, we use a different argument, namely that a zero-mean random vector $x\in\mathbb{R}^{d}$, which is $\sigma$-sub-Gaussian in the inner product sense, i.e., for any $v\in\mathbb{R}^{d}$
-$\mathbb{E}\big[\exp\big(\langle x,v\rangle\big)\big]\leq\exp\bigg(\frac{\sigma^{2}\|v\|^{2}}{2}\bigg),$
-is also $\sigma\sqrt{d}$-norm-sub-Gaussian, i.e., for any $\epsilon>0$
-$\mathbb{P}\big(\|x\|>\epsilon\big)\leq 2\exp\bigg(-\frac{\epsilon^{2}}{2\sigma^{2}d}\bigg),$
-see Lemma 1 in [27].
-This argument allows us to show the variance reduction benefit of decentralized learning, i.e., that $\overline{z}^{t}$ is $\mathcal{O}\Big(\frac{\sigma\sqrt{d}}{\sqrt{n}}\Big)$-sub-Gaussian (in the sense of (A4)), at the cost of introducing a $\sqrt{d}$ dependence.
-One way to mitigate this is to directly assume that the noise $z^{t}_{i}$ at each user is zero-mean and $\sigma_{i}$-norm-sub-Gaussian.
-Using a similar argument to the one in Lemma 6 and Corollary 7 in [27], we could then show that $\overline{z}^{t}$ is $\frac{\sigma}{n}$-sub-Gaussian, while reducing the dependence on problem dimension to $\log(d)$.
-This, however, comes at the cost of imposing a slightly stronger noise condition, as any $\sigma$-norm-sub-Gaussian random vector is also $\sigma$-sub-Gaussian (in the inner product sense), while the opposite implication inevitably introduces a $\sqrt{d}$ factor, as is the case in Lemma 2.
-Finally, we note that, in the regime $n\ll d$, we can simply use the above argument following Sedrakyan’s inequality, to conclude that $\overline{z}^{t}$ is $\sigma$-sub Gaussian, reducing the dependence on problem dimension $d$, at the cost of losing linear speed-up in the number of users.
+To circumvent this issue, we use a different argument, namely that condition (A4) is equivalent to norm-sub-Gausiannity, i.e., that for any $\epsilon>0$, we have
+$\mathbb{P}\big(\|z^{t}_{i}\|^{2}>\epsilon\>|\>\mathcal{F}_{t}\big)\leq 2\exp\bigg(\frac{\epsilon^{2}}{2\sigma_{i}^{2}}\bigg),$
+which, combined with Proposition 9 (i.e., Corollary 7 in [29]), implies that
+$\mathbb{P}\big(\|\overline{z}^{t}\|^{2}>\epsilon\big)\leq 2d\exp\bigg(\frac{n\epsilon}{3\sigma^{2}}\bigg).$
+This result can be further leveraged to show the desired bound on the MGF of $\overline{z}^{t}$, while introducing the mild $\log(d)$ dependence, see the proof of Lemma 2 for details.
+It can be noted that, when there is no noise, i.e., $\sigma=0$, the dimension dependence vanishes, highlighting that this dependence stems from the noise.
+As mentioned in [29], while mild, it is not clear whether this dependence on $d$ can be completely removed.
+We now highlight three important facts.
+First, if a different definition of sub-Gaussian vectors is used, namely the one via inner products, which states that, for any $\mathcal{F}_{t}$-measurable vector $v$, we have
+$\mathbb{E}\big[\exp\big(\langle v,z^{t}_{i}\rangle\big)\>|\>\mathcal{F}_{t}\big]\leq\exp\bigg(\frac{\sigma_{i}^{2}\|v\|^{2}}{2}\bigg),$
+this would inevitably lead to a worse dependence on the problem dimension, as in order to bound the MGF of the squared norm of the average noise, i.e., $\|\overline{z}^{t}\|^{2}$, one would need to use a similar argument as the one for Lemma 1 in [29], using covers of the unit sphere and picking up a factor of $d$ in the final rate, which is much worse than our $\log(d)$ dependence.
+Second, while the notion of sub-Gaussian random vectors used in (A4) is equivalent to the notion of norm-sub-Gaussian vectors from [29] and stronger than the notion of sub-Gaussian vectors defined via inner products above, it is in fact the standard notion of sub-Gaussianity used in almost all HP results, in both centralized and decentralized settings, see [58, 20, 40, 42, 44, 47, 48, 83] and references therein.
+Finally, we note that in the extreme regime $n\ll\log(d)$, we can simply use the argument in (L), to conclude that $\overline{z}^{t}$ is $\sigma$-sub-Gaussian, completely removing the dependence on dimension $d$, at the cost of losing linear speed-up in the number of users.